### v1

# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian
Satadal Ganguly, Rachita Guria Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, 203 Barrackpore Trunk Road, Kolkata-700108, India. sgisical@gmail.com guriarachita2011@gmail.com
Key words and phrases:
Lattice points, Fourier coefficients of cusp forms, Kloosterman sum, Large sieve inequality
2020 Mathematics Subject Classification:
Primary 11P21, 11N45, 11F30, 11F72; Secondary 11N36, 11L07
1. Introduction
The famous Gauss circle problem asks for a precise count for the number $N(r)$ of lattice points, i.e., points with integer coordinates, in the Euclidean plane inside a circle of radius $r$ as $r$ increases to infinity. Obtaining the asymptotic relation $N(r)\sim\pi r^{2}$ (as $r\longrightarrow\infty$) is an elementary exercise and the real task is to estimate the growth of the error term $E(r)=N(r)-\pi r^{2}$. The current record is
$E(r)=O(r^{131/208+\varepsilon}),\ \text{for any }\varepsilon>0,$
due to Huxley [Hux96]  [^1] Recently, Xiaochun Li and Xuerui Yang have posted an article (see https://arxiv.org/abs/2308.14859) that claims to improve this bound of Huxley., while the conjectured bound $O(r^{1/2+\varepsilon})$ remains hopelessly out of reach of the current technology. If we replace the circle by a hyperbola $xy=N$, with $N\longrightarrow\infty$, the problem of counting lattice points with positive coordinates under the hyperbola is known as Dirichlet’s divisor problem and the situation here is completely analogous to the Gauss circle problem (see [IK04, Chap. 1]). An interesting lattice point counting problem concerns counting elements $\gamma\in SL(2,{\mathbb{Z}})$; i.e., integer matrices $\gamma=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ with determinant $ad-bc=1$, such that $||\gamma||\leq x$, where we take the norm to be the standard Euclidean norm on ${{\mathbb{R}}}^{4}$ after identifying the matrix $\gamma$ with the point $(a,b,c,d)$; i.e.,
$||\gamma||^{2}=a^{2}+b^{2}+c^{2}+d^{2}.$
This problem can be interpreted as the analogue of the Gauss circle problem in the hyperbolic plane $\mathbb{H}$. To see why, note that if $d(z,w)$ denotes the standard hyperbolic distance $\mathbb{H}$ defined by
$d(z,w)=\frac{4|z-w|^{2}}{\Im z\Im w}+2,$
then $d(\gamma.i,i)=a^{2}+b^{2}+c^{2}+d^{2}$, where $\gamma.z$ denotes the usual Möbius action $\gamma.z=(az+b)(cz+d)^{-1}.$ Thus the above problem is equivalent to studying the asymptotic behaviour of the size of the set $\{\gamma\in\Gamma:d(\gamma.i,i)\leq x\}$, where $\Gamma$ is the group $SL(2,{\mathbb{Z}})$ and one may also consider a more general problem by taking $\Gamma$ to be any discrete subgroups of $SL(2,{\mathbb{R}})$ with finite co-volume. Unlike in the Euclidean case, even finding the asymptotic for lattice point counting problems is not easy and the elementary counting arguments no longer work due to the negative curvature of the underlying space. Lattice point counting problems are intimately related to the spectrum of the Laplace operator on appropriate manifolds. For example, the Gauss circle problem is nothing but the problem of improving the bound on the error term in the Weyl law for the flat torus ${{\mathbb{Z}}}^{2}\backslash{{\mathbb{R}}}^{2}$. The above hyperbolic lattice point counting problem is related to the spectrum of the hyperbolic Laplacian acting on $\ΓL^{2}(\Gamma\backslash\mathbb{H})$ (as a densely defined operator) and after initial advances made by Delsarte [Del42] and Huber, who solved the problem for cocompact groups (see [Hub56, Hub61]), the original problem for $\Gamma=SL(2,{\mathbb{Z}})$ could be tackled once Selberg established the complete spectral resolution of the Laplacian for groups with parabolic elements. Selberg’s result (see, e.g., [IK04, Corollary 15.12]) says
$\sum_{\begin{subarray}{c}\gamma\in SL(2,{\mathbb{Z}})\\ ||\gamma||\leq x\end{subarray}}1=6x^{2}+O(x^{4/3}).$ (1.1)
For an account of this beautiful and deep circle of ideas, see, e.g., [IK04, Chap. 15], [Iwa02], [LP82], [Pat75] and the original articles by Selberg, viz. [Sel56], [Sel63].
This result has been vastly generalized by Duke, Rudnick and Sarnak [DRS93] to varieties $V$ with an algebraic group action for which $V(\mathbb{R})$ is affine and symmetric by applying advanced results from harmonic analysis. A special case of their result says
$\sum_{\begin{subarray}{c}\gamma\in M(n,{\mathbb{Z}})\\ det(\gamma)=r\\ \\ ||\gamma||\leq x\end{subarray}}1=c_{n,r}x^{n^{2}-n}+O_{\varepsilon}(x^{n(n-1)-1/(n+1)+\varepsilon}),$ (1.2)
[…]
Theorem 1.1.
Let
[…]
where $a,b,c,d$ vary over the integers and $r$ is a fixed non-zero integer. We have,
$\mathpzc{S}_{r}(X)=\frac{16}{\zeta(2)}\frac{\sigma(|r|)}{|r|}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}),$ (1.3)
as long as $r=O\left(X^{{1/3}}\right)$. Here $\theta$ is any admissible exponent for the Generalized Ramanujan Conjecture, $\varepsilon$ is an arbitrarily small positive real number, and $\sigma(n)$ denotes the sum of the divisors of a positive integer $n$.
By the work of Kim and Sarnak [Kim03, Appendix 2] it is known that $\theta\leq\frac{7}{64}$.
Remark 1.2.
The reason we need to take $r=O(X^{1/3})$ is explained in Remark 16.1. However, with more work, it seems possible to extend the range of $r$. Also, at the cost of a worse exponent of $X$, we can make the dependence of $r$ on the error term disappear. See Remark 7.6 for a precise statement.
As mentioned earlier, we can put our problem in the general context of counting integral points on affine hypersurfaces defined over ${\mathbb{Z}}$ and more specifically to counting vectors $\mathbf{a}\in{{\mathbb{Z}}}^{n}$ to equations of the form
$Q(\mathbf{x})=r$ (1.4)
in an expanding box $[-X,X]^{n^{2}}$, where $Q$ is a quadratic form in $n$ variables over ${\mathbb{Z}}$ and $r$ is a fixed integer. This is a classical problem with a rich history and the circle method, developed and refined by Ramanujan, Hardy, Littlewood, Kloosterman and many others, has been employed here with great success. See the article [HB96] by Heath-Brown for an introduction with modern perspectives to this problem. Usually one counts integral solutions to (1.4) by attaching weights $w_{i}(a_{i}/X)$, where $w_{i}$’s are smooth functions with compact support and it is possible to get fairly strong bound on the error term relative to the main term in many cases (see, e.g., [HB96, Theorems 4-8]). If we count with sharp cuts, however, then the situation is very different and in general we can only get a saving of an arbitrary power of $\log X$ (see [Nie10]). It is interesting, therefore, to address this problem for forms of special shapes in a small number of variables and the determinant form $Q(x,y,z,w)=xy-zw$ is particularly suitable because of the structural advantage it affords. In fact, the only two results we could find in the literature that deal with unweighted counting are about determinant forms in two or more variables. Gorodnik and Nevo (see [GN10, Cor. 2.3]) studied the problem of counting matrices in $SL(n,{\mathbb{Z}})$ inside a ball with respect to a general norm using ergodic method and in the case $n=2$, the saving in the exponent in the error term is $1/16$. Roettger [Roe05] obtained an asymptotic formula for the number of unimodular matrices over the ring of integers in a number field where the elements have height up to some parameter $t$. When specialized to our case $K={\mathbb{Q}}$, the error term gives a saving of $1/15$ in the exponent. In comparison, we save $1/2$ in the exponent of the error term in Theorem 1.1.
If we count with smooth weights then we can save even more and can also take $r$ to be much larger as the following result shows.
Theorem 1.3.
Let
$S_{V}(X,r)=\mathop{\sum_{a}\sum_{b}\sum_{c}\sum_{d}}_{\begin{subarray}{c}ad-bc=r\end{subarray}}V\left(\frac{a}{X}\right)V\left(\frac{b}{X}\right)V\left(\frac{c}{X}\right)V\left(\frac{d}{X}\right),$
where $V$ is a fixed smooth function with compact support in $[1,2]$. Then for $0<r\ll X^{2}$, we have
$S_{V}(X,r)=M_{V}(X,r)+O_{\varepsilon}\left(r^{1/4}X^{1+\varepsilon}\right),$
where
$M_{V}(X,r)=\sum_{l|r}\sum_{k>0}\frac{\mu(k)}{k}\iiint\frac{1}{z}V\left(\frac{x}{X}\right)V\left(\frac{lky}{X}\right)V\left(\frac{lkz}{X}\right)V\left(\frac{r+lkxy}{zlkX}\right)\textnormal{d}x\textnormal{d}y\textnormal{d}z.$
It is not hard to see that $M_{V}(X,r)\asymp X^{2}$ if $r$ is not very large; e.g., if $0<|r|\leq 2X^{2}$. Moreover, by the mean value theorem of Calculus, we see that
$S_{V}(X,r)=KX^{2}+O\left(\sigma(r)\right)+O\left(r^{1/4}X^{1+\varepsilon}\right),$
where
$K=\frac{1}{\zeta(2)}\frac{\sigma(r)}{r}\iiint\frac{1}{t}V\left(u\right)V\left(v\right)V\left(t\right)V\left(\frac{uv}{t}\right)\textnormal{d}u\,\textnormal{d}v\,\textnormal{d}t.$
Although we use automorphic methods to prove this result, a result of comparable strength can also be obtained by employing the method of [HB96]. In fact, if we take $r$ to be of size $X^{2}$ then Theorem 1.3 gives an error term of the same strength as the one in [HB96, Theorem 4]. However, it is possible to improve the exponent of $r$ to $\theta$ with more work. It is of interest to consider more general quadratic forms and to investigate whether some of the results of [HB96] (see also [Get18]) can be refined in special cases by introducing the spectral methods of automorphic forms into the analysis after the application of the $\delta$-symbol method. In this context, see also the discussion after [HB96, Cor. 1]. We have included the above result to illustrate the impact of counting with smooth weights and have given only a brief sketch of the proof in §1.2 since this proof is much easier compared to the proof of Theorem 1.1. An interesting variation of the general counting problem is to attach arithmetic weights such the von Mangold function or coefficients of cusp forms to the variables of summation (see [Kum20]) and this is the topic of an upcoming work by the second author along with P. Sharma and S.K. Singh.
1.1. Extensions and applications of the results
1.1.1. Determinant equations in Analytic Number Theory
Although we are mainly motivated by the intrinsic interest in the problem, counting solutions to the determinant equations $ad-bc=r$, $r\neq 0$ with certain weights $\alpha(a),\alpha(b)$ etc. come up naturally in the subject, for example, in the study of amplified moments of $L$-functions (see, e.g., [DFI97], [DFI02, §16], [BC18]). In [BC18] and [DFI97], sums of the form
$\mathop{\sum_{n_{1}}\sum_{n_{2}}\sum_{n_{3}}\sum_{n_{4}}}_{\begin{subarray}{c}n_{i}\in\mathcal{N}_{i}1\leq i\leq 4\\ n_{1}n_{4}-n_{2}n_{3}=r\end{subarray}}\alpha(n_{1})\beta(n_{2})f(n_{3})g(n_{4})$ (1.5)
were studied and asymptotic formulae for such sums were obtained. Here, $\mathcal{N}_{i}=[N_{i}/2,N_{i}]$, $\alpha,\beta$ are arbitrary sequences, and $f,g$ are smooth functions satisfying suitable decay conditions on their derivatives. It is possible to allow arbitrary weights in some of the variables at the cost of an worse error term. For example, the following is a consequence of an intermediate step in the proof of Theorem 1.1.
Corollary 1.4.
For an arbitrary sequence of complex numbers $\alpha(a)=O(a^{\varepsilon})$, and for any non-zero integer $r$, we have
$\mathop{\sum_{a}\sum_{b}\sum_{c}\sum_{d}}_{\begin{subarray}{c}ad-bc=r\\ |a|,|b|,|c|,|d|\leq X\end{subarray}}\alpha(a)=A_{w,\alpha}(X,r)+O\left(X^{5/3+\varepsilon}\right),$
as long as $r=O\left(X^{5/3}\right)$, where
$A_{w,\alpha}(X,r)=\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a,c)=l\\ c\neq 0\\ a>1\end{subarray}}\frac{l\alpha(a)}{a}w(a)w(c)\int_{1}^{X}w\left(\frac{cx}{a}\right)\textnormal{d}x,$
and the function $w$ satisfies the conditions (3.4)-(3.7).
This result is a direct consequence of (3.2), (3.9), (3.10) and Remark 7.6.
Unlike (1.5), here only one variable is weighted by an arbitrary sequence.
If one chooses weights of arithmetic nature then further possibilities for improvements may open up. For example, if we take $\alpha(a)=\mathbf{1}_{a\equiv 0(\text{mod }q)}$, the indicator function of the integral multiples of $q$, where $q>1$ is a fixed integer and keep the other weights the constant function $1$ (or smooth weights $W(a/X),W(b/X)$ etc.), then also our method applies and this leads to the study of the spectral theory for the congruence group $\Gamma_{0}(q)$.
1.1.2. Extensions to other quadratic forms in four variables
The method of our proof does not distinguish between the determinant equation $xy-zw=r$ and the equation $pxy-qzw=r$ where $p,q$ are fixed integers such that $(p,q)|r$. One can easily obtain analogous results for the latter equation of the same strength as long as $p,q=O(X^{\varepsilon})$. In fact, by a simple change of variables, we can derive an asymptotic formula for the number of integer solutions to an equation of the form $f_{1}(x,y)-f_{2}(z,w)=r$, where $f_{1}$ and $f_{2}$ both factor over ${\mathbb{Z}}$ as products of two linear forms and the coefficients of both forms are assumed to be $O(X^{\varepsilon})$. However, the formula we obtain in this way counts the number of solutions in a new domain that depends on the coefficients of $f_{1}$ and $f_{2}$. To clarify the situation, let us give an illustrative example. Let us take $f_{1}(x,y)=f_{2}(x,y)=x^{2}-y^{2}$. Then we obtain the surface $x^{2}-y^{2}-z^{2}+w^{2}=r$. Now, by the change of variables $(x,y,z,w)\rightarrow(x^{\prime},y^{\prime},z^{\prime},w^{\prime})$, where $x+y=x^{\prime},x-y=y^{\prime},z+w=z^{\prime},z-w=z^{\prime}$, we obtain the surface $x^{\prime}y^{\prime}-z^{\prime}w^{\prime}=r$. By Theorem 1.1, we have a counting formula for the number of integer solutions to $x^{\prime}y^{\prime}-z^{\prime}w^{\prime}=r$, where $(x^{\prime},y^{\prime},z^{\prime},w^{\prime})\in[-X,X]^{4}$. Now the inverse change of variables transform this box to a four-dimensional tetrahedron $T(X)$ which can be described as the convex hull of the points
$(X,0,0,0),(-X,0,0,0),(0,X,0,0),(0,-X,0,0),\cdots,(0,0,0,X),(0,0,0,-X).$
Thus we obtain the following asymptotic formula:
Corollary 1.5.
We have,
$\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}a^{2}-b^{2}-c^{2}+d^{2}=r\\ a,b,c,d\in T(X)\end{subarray}}1=\frac{16}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}).$
1.1.3. Extensions to quadratic forms in three variables
If we consider quadratic forms of the shape $Q(x,y,z)=px^{2}-qyz$, then also our methods can be adopted to study integral solutions to $Q(x,y,z)=r$. Here, one encounters a sum of Salié sums instead of Kloosterman sums and Proskurin’s formula generalizing the Kuznetsov formula (see [Pro03]) is applicable here.
1.1.4. Relation to the shifted convolution sum problem for the restricted divisor function
The shifted convolution sum refers to the problem of the estimation or obtaining asymptotic formula for sums of the type $\sum_{n\leq x}\lambda(n)\lambda(n+h)$ or the corresponding smoothed sum, where $\lambda$ is an interesting arithmetical function, e.g., the divisor function, coefficients of a cusp form etc., and the shift $h$ is a fixed non-zero integer. Lattice point counting problems can sometimes be interpreted as the shifted convolution problem for coefficients of appropriate automorphic forms. For example, the formula (1.1) is equivalent to the formula
$\sum_{n\leq x}r(n)r(n+1)=6x^{2}+O(x^{4/3}),$
where $r(n)$ are the coefficients of ${\theta}^{2}(z)$, $\theta(z)$ being the classical $\theta$-function. If we take $\lambda(n)=\tau(n)$, the divisor function, then the corresponding problem (the additive divisor problem) has a rich history (see [Ing27], [Est32], [DI82], [Mot94], [Meu01]). Now, for a fixed positive integer $M$ we define the restricted divisor function ${\tau}^{\ast}_{M}$ by
${\tau}^{\ast}_{M}(n)=\#\{(a,b)\in{{\mathbb{Z}}}^{2}:1\leq a,b\leq M,ab=n\}.$
Truelsen (see [Tru10]) introduced this function in connection with the problem of determining the irrational numbers $\alpha$ for which the pair correlation for the fractional parts of $n^{2}\alpha$ is Poissonian. We can interpret (1.3)(or more accurately, (3.1)) as an asymptotic formula for the shifted convolution sum for ${\tau}^{\ast}_{\sqrt{X}}$:
Corollary 1.6.
We have, for $0<|r|\ll X^{{1/6}}$,
$\sum_{n\leq X}{\tau}^{\ast}_{\sqrt{X}}(n){\tau}^{\ast}_{\sqrt{X}}(n+r)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X+O({|r|}^{\theta}X^{3/4+\varepsilon}).$ (1.6)
Compared to the classical additive divisor problem, estimating the error term for this sum is harder and the reason is clear. Writing $n=bc$ and $n+r=ad$, here all the variables are individually restricted by sharp edges, whereas in the sum $\sum_{n\leq X}\tau(n)\tau(n+h)$, the four variables are subject to only one such restriction ($bc\leq X$) and it is possible to treat two of the three independent variables as “smooth” variables.
1.2. Method of the proofs of Theorem 1.1 and Theorem 1.3
The theorems above are proved by the so called “Kloostermania” techniques, i.e., we reduce the problems to that of estimation of sums of Kloosterman sums which is then carried out by employing the spectral theory of automorphic forms. Let us first consider the proof of Theorem 1.3 which serves as a warm-up for the proof of Theorem 1.1. The first step is to eliminate one of the four variables, say $d$, and to interpret the determinant equation as a congruence modulo one of the other variables, say, $a$. This is a natural and well-known idea. Then we apply Poisson summation (see Lemma 2.1) to one of the remaining two variables, say $b$. As usual, the 00-th frequency in the dual sum gives the main term and after one more application of Poisson summation in the variable $c$ to the remaining sum yields a sum of Ramanujan sums (coming from the $zero$-th term in new dual variable) that can be easily estimated, plus the following double sum:
$\sum_{\begin{subarray}{c}n\in{\mathbb{Z}}\\ n\neq 0\end{subarray}}\sum_{\begin{subarray}{c}m\in{\mathbb{Z}}\\ m\neq 0\end{subarray}}\iint V\left(\frac{x}{X}\right)V\left(\frac{y}{X}\right)\sum_{a}\frac{1}{a^{2}}V\left(\frac{a}{X}\right)V\left(\frac{r+yx}{aX}\right)e\left(\frac{-nx}{a}\right)e\left(\frac{-my}{a}\right)S(nr,-m,a)\textnormal{d}x\,\textnormal{d}y.$
At this stage, the Weil bound already gives an asymptotic formula but we get a better error term by applying spectral methods of automorphic forms via the Kuznetsov formula. Since the variable $a$ is weighted by a “flat” function $V(x/X)$ (i.e., $y^{j}V^{(j)}(y)\ll 1$), it is not difficult to obtain strong enough bounds for the Bessel transforms after application of the Kuznetsov formula so as to estimate the spectral side efficiently. For example, a direct application of [IK04, Theorem 16.9]) yields Theorem 1.3. The extra saving one gets by averaging over the variable $a$ is reminiscent of what Heath-Brown calls “double Kloosterman refinement” (see [HB96, §1]).
In contrast to the proof of Theorem 1.3, several new features emerge when we employ the above strategy towards proving Theorem 1.1. After the second application of Poisson summation, if we apply Kuznetsov formula directly then estimating the Bessel transforms that come up becomes quite difficult since we no longer have the advantage of a flat weight function. Instead we study the Fourier transform $\widehat{G}(n/a_{1},m/a_{1})$ (see (5.1)) carefully and extract stationary phases of the type $e\left(-2\sqrt{\frac{mnX}{a_{1}}}\right)$ (see Prop. 3.3). However, a straight-forward application of the stationary phase lemmas to the entire interval does not give good error terms and therefore, we need to carry out a delicate analysis involving subdivision of both the range of integration and the range of the stationary point which depends on the three variables of summation $a_{1}$, $m,$ and $n$ (see §5.2 to §7). When the stationary point is sufficiently localized, the variables are also under significant restrictions and exploiting these restrictions is crucial in estimating the contributions of the error terms arising out of the stationary phase analysis.
After we apply the Kuznetsov formula, the phase of the $K$-Bessel function combines with the already existing phase and in certain ranges of the variables we are able to extract a stationary phase that “looks like” $(mn)^{i\kappa_{j}}$ (see(LABEL:lt)). Now, a large sieve inequality (see Lemma 2.16) for the twisted coefficients $\rho_{j}(n)n^{i\kappa_{j}}$, that arose in a work of Deshouillers and Iwaniec [DI86] in connection with the Phillips-Sarnak theory on destruction of cusp forms for generic groups (see [PS85]), gives us an extra saving that is crucial. In complimentary ranges, the Weyl law and the local (spectral) large sieve of Jutila (Lemma 2.11) are essential. In the end we are able to exploit cancellations in each of the three independent variables $a,b$ and $c$. Apart from the results on automorphic forms and the Weil bound for Kloosterman sums, the stationary phase method in its different formulations plays a crucial role in the proof.
1.3. Notations and conventions
The letter $\varepsilon$ denotes a positive real number which can be taken to be as small as required and in different occurrences it may assume different values. The notation “$f(y)=O(g(y))$” or “$f(y)\ll g(y)$”, where $g$ is a positive function, means that there is a constant $c>0$ such that $|f(y)|\leq cg(y)$ for any $y$ in the concerned domain. The dependence of this implied constant on some parameter(s) may sometimes be displayed by a suffix (or suffixes) and may sometimes be suppressed. Also, $f(y)\asymp g(y)$ denotes $f(y)\ll g(y)$ as well as $g(y)\ll f(y)$. The symbol :=assign:=:= signifies that the left hand side is being defined. For a positive integer $n$, $\tau(n),\sigma(n),\phi(n)$ denotes, respectively the total number of positive divisors of $n$, the sum of the divisors of $n$, and the number of residue classes modulo $n$ that are prime to $n$. The symbol $\mathcal{L}(I)$ denotes the length of an interval $I$. The symbol $\sim$, as in $\sum_{n\sim N}$ denotes $n$ is varying in the interval $[N,4N)$.
Acknowledgements
The authors thank Tim Browning, Gergely Harcos, Vinay Kumaraswamy, Gopal Maiti, Ritabrata Munshi, D. Surya Ramana, Prahlad Sharma and Igor Shparlinski for helpful comments. This work is part of the second author’s PhD thesis. She thanks Indian Institute of Science Education and Research, Kolkata and all her teachers, especially Subrata Shyam Roy, for all the support. The authors also thank Indian Statistical Institute, Kolkata where this work was carried out for excellent working atmosphere.
2. Preliminaries
In this section we recall some standard results from the literature which will be required in the proofs of our results formula and some applications of the formula.
Lemma 2.1.
(Application of Poisson summation)
[…]
$\sum_{\begin{subarray}{c}n\equiv a(\textnormal{mod }q)\\ n\in{\mathbb{Z}}\end{subarray}}f\left(n\right)=\frac{1}{q}\sum_{n\in{\mathbb{Z}}}e\left(\frac{an}{q}\right)\hat{f}\left(\frac{n}{q}\right).$
Proof.
See, e.g. ,[IK04, Theorem 4.4 and Exercise 4]. ∎
Lemma 2.2.
Under the same assumptions on $a,q$ and $f$ as in Lemma 2.1, we have,
$\sum_{(n,q)=1}g(n)e\left(\frac{a\bar{n}}{q}\right)=\frac{1}{q}\sum_{n}\hat{g}(n/q)S(n,a,q),$
where
$S(m,n,q)=\sideset{}{{}^{\star}}{\sum}_{\beta(\textnormal{mod }q)}e\left(\frac{m\beta+n\bar{\beta}}{q}\right)$
is the Kloosterman sum (see [IK04]) and $\bar{\beta}$ denotes the multiplicative inverse of $\beta$ modulo $q$ (if $q=1$, $\bar{\beta}=S(m,a;q)=1$ by convention).
Proof.
This follows from subdividing the sum over $n$ into residue classes coprime to $q$ and then applying Lemma 2.1. ∎
Lemma 2.3.
(The Weil bound for Kloosterman sums)
Let $a,b,c$ be integers, $c$ positive. We have
$\left|S(a,b;c)\right|\leq\tau(c)(a,b,c)^{\frac{1}{2}}c^{\frac{1}{2}}.$
Proof.
See [IK04, Corollary (11.12)]. ∎
Lemma 2.4.
(The Perron formula)
$\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{x^{s}}{s}\textnormal{d}s=h(x)+O\left(x^{c}\min\left(1,\frac{1}{T|\log x|}\right)\right),$
where
$h(x)=\begin{cases}1,&\text{ if }x>1\\ \frac{1}{2},&\text{ if }x=1\\ 0,&\text{ if }x<1,\end{cases}$
for any $x>0$, $x\neq 1$, and any $T,c>0$, with an absolute implied constant. If $x=1$, the same equality holds with the error term replaced by $O(c/T)$.
Proof.
See [Dav80, Chap. 17]. ∎
2.1. Results on exponential integrals
Lemma 2.5.
(The $k$-th derivative bound) Suppose $\psi\in C^{1}[a,b]$. Let $\phi$ be a real valued and smooth in $(a,b)$ such that $\left|\phi^{(k)}\right|\geq\lambda_{k}$ for all $x\in(a,b)$. Then
$\left|\int\limits_{a}^{b}e^{i\phi(x)}\psi(x)\textnormal{d}x\right|\leq c_{k}\lambda_{k}^{-1/k}\left[\|\psi\|_{\infty}+\int\limits_{a}^{b}\left|\psi^{\prime}(x)\right|\textnormal{d}x\right]$
holds when $k\geq 2$, or $k=1$ and $\phi^{\prime}(x)$ is monotonic. Here $c_{k}>0$ is a constant that depends only on $k$.
Proof.
See [Ste93, Eq. (6) & Prop. 2, Chap. 8]. ∎
Lemma 2.6.
(The stationary phase lemma) Let $g(x)$ be real-valued function on $[a,b]$ with $g^{\prime}(x_{0})=0$ for some $x_{0}\in[a,b]$. Suppose
$0<\lambda_{2}\leq g^{\prime\prime}(x)$ (2.1)
holds throughout the interval, and in addition $|g^{(3)}(x)|\leq\lambda_{3}$ and that $|g^{(4)}(x)|\leq\lambda_{4}$ for all $x\in[a,b]$. Then
$\int\limits_{a}^{b}e(g(x))\textnormal{d}x=\frac{e(g(x_{0})+1/8)}{\sqrt{g^{\prime\prime}(x_{0})}}+O(R_{1})+O(R_{2}),$
where
$R_{1}=\min\left(\frac{1}{\lambda_{2}(x_{0}-a)},\frac{1}{\sqrt{\lambda_{2}}}\right)+\min\left(\frac{1}{\lambda_{2}(b-x_{0})},\frac{1}{\sqrt{\lambda_{2}}}\right)\textnormal{ and }R_{2}=\frac{(b-a)\lambda_{4}}{\lambda_{2}^{2}}+\frac{(b-a)\lambda_{3}^{2}}{\lambda_{2}^{3}}.$
If instead of (2.1) we have $g^{\prime\prime}(x)\leq-\lambda_{2}<0$ throughout the interval, then the main term becomes
$\frac{e(g(x_{0})-1/8)}{\sqrt{|g^{\prime\prime}(x_{0})|}}.$
Proof.
See [Mon94, Theorem 9]. ∎
Lemma 2.7.
(The weighted stationary phase lemma) Let $f(x)$ be a real function, four times continuously differentiable for $\alpha\leq x\leq\beta$, and let $g(x)$ be a real function, three times continuously differentiable for $\alpha\leq x\leq\beta$. Suppose that there are positive parameters $T,P,V,Q,$ with
$Q\geq\beta-\alpha,\quad V\geq\frac{Q}{\sqrt{P}},$
and positive constants $C_{r}$ such that, for $\alpha\leq x\leq\beta$,
$\left|f^{(r)}(x)\right|\leq\frac{C_{r}P}{Q^{r}},\quad\left|g^{(s)}(x)\right|\leq\frac{C_{s}T}{V^{s}}\quad\textnormal{and}\quad f^{\prime\prime}(x)\geq\frac{P}{C_{2}Q^{2}},$
for $r=2,3,4$, and $s=0,12,3$. Suppose also that $f^{\prime}(x)$ changes sign from negative to positive at a point $x=x_{0}$ with $\alpha<x_{0}<\beta$. If $P$ is sufficiently large in terms of $C_{2},C_{3}$, and $C_{4}$, then we have
$\int\limits_{\alpha}^{\beta}g(x)e(f(x))\textnormal{d}x=\frac{g(x_{0})e\left(f(x_{0})+\frac{1}{8}\right)}{\sqrt{f^{\prime\prime}(x_{0})}}+\frac{g(\beta)e(f(\beta))}{2\pi if^{\prime}(\beta)}-\frac{g(\alpha)e(f(\alpha))}{2\pi if^{\prime}(\alpha)}$
$+O\left(\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha)^{3}}+\frac{1}{(\beta-x_{0})^{3}}\right)\right)+O\left(\frac{QT}{P^{3/2}}\left(1+\frac{Q}{V}\right)^{2}\right).$
Proof.
See [Hux96, Lemma 5.5.6]. ∎
Lemma 2.8.
(The stationary phase lemma with smooth weights) Let $Y\geq 1,T,Q,U,R>0$ and suppose $w$ is a smooth function with support in $[\alpha,\beta]$ satisfying
$w^{(j)}(t)\ll_{j}TU^{-j}.$
Suppose that $h$ is a smooth function on $[\alpha,\beta]$ such that
$\left|h^{\prime}(t)\right|\geq R,\quad\textnormal{ and }\quad h^{(j)}(t)\ll_{j}YQ^{-j},\textnormal{ for }j=2,3\dots,$
for some $R>0$. Then the integral $I$ defined by
[…]
satisfies
$I\ll_{A}(\beta-\alpha)T\left[\left(\frac{QR}{\sqrt{Y}}\right)^{-A}+(RU)^{-A}\right].$
Proof.
See [BKY13, Lemma 8.1]. ∎
Lemma 2.9.
(Asymptotic expansion of exponential integrals with smooth weights) Let $0<\delta<\frac{1}{10},T,Y,V,V_{1},Q>0,$ $Z=Q+X+Y+V_{1}+1$ and assume that
$Y\geq Z^{3\delta},\quad\quad V_{1}\geq V\geq\frac{QZ^{\frac{\delta}{2}}}{\sqrt{Y}}.$ (2.3)
Suppose $w$ is a smooth function on ${\mathbb{R}}$ with support on an interval $J$ of length $V_{1}$ satisfying
$w^{(j)}\ll_{j}TV^{-j},$
for all $j\in{\mathbb{N}}_{0}$, and $h$ is a smooth function on $J$ such that there exists a unique point $t_{0}\in J$ such that $h^{\prime}(t_{0})=0$, and furthermore that
$h^{\prime\prime}(t)\gg YQ^{-2},\quad h^{(j)}(t)\ll_{j}YQ^{-j},\text{ for }j=1,2,3,\dots,t\in J.$
Suppose also that $h^{\prime}(t)$ changes sign from negative to positive at a point $t=t_{0}$. Then the integral $I$ defined in (2.2) has an asymptotic expansion of the form
$I=\frac{e^{ih(t_{0})}}{\sqrt{h^{\prime\prime}(t_{0})}}\sum\limits_{n\leq 3\delta^{-1}A}p_{n}(t_{0})+O_{A,\delta}\left(Z^{-A}\right),$ (2.4)
where
$p_{n}(t_{0})=\frac{\sqrt{2\pi}e^{\frac{\pi i}{4}}}{n!}\left(\frac{i}{2h^{\prime\prime}(t_{0})}\right)^{n}G^{(2n)}(t_{0}),\text{ }G(t)=w(t)e^{iH(t)},\textnormal{ and }H(t)=h(t)-h(t_{0})-\frac{h^{\prime\prime}(t_{0})}{2}(t-t_{0})^{2},$ (2.5)
for some arbitrary $A>0$. Furthermore, each $p_{n}$ is a rational function in $h^{\prime\prime},h^{\prime\prime\prime},\dots,$ satisfying
$\frac{d^{j}}{dt_{0}^{j}}p_{n}(t_{0})\ll_{j,n}T\left(V^{-j}+Q^{-j}\right)\left(\left(\frac{V^{2}Y}{Q^{2}}\right)^{-n}+Y^{-\frac{n}{3}}\right).$ (2.6)
Proof.
See [BKY13, Prop. 8.2]. ∎
2.2. Results from the theory of automorphic forms
We refer to [DI82] for the definitions and properties of Maass cusp forms and related topics. Let us fix an orthonormal basis of Maass cusp forms $\{u_{j}(z)\}_{1}^{\infty}$ for the full modular group $\text{SL}(2,{\mathbb{Z}})$, consisting of common eigenfunctions of all the Hecke operators $T_{n},n\geq 1$. Thus every $u_{j}(z)$ is an eigenfunction of the Laplace operator with the associated eigenvalue $s_{j}(1-s_{j})$, $s_{j}\in{\mathbb{C}}$. We write $s_{j}$ as $s_{j}=\frac{1}{2}+i\kappa_{j}$ with $\kappa_{j}\in{\mathbb{C}}$. It is known that $\kappa_{j}\in{\mathbb{R}}$ and in fact $|\kappa_{j}|>3.815$ (see [Mot97, Lemma 1.4]). Every $u_{j}(z)$ has the following Fourier expansion at the cusp at $\infty$:
$u_{j}(z)=\sqrt{y}\sum_{n\neq 0}\rho_{j}(n)K_{i\kappa_{j}}(2\pi|n|y)e(nx),$
where $z=x+iy,$ $x,y\in{\mathbb{R}}$. Note that $K_{i\kappa}(y)=K_{-i\kappa}(y)$. Hence, without any loss of generality, we may assume that $\kappa_{j}>0$. Suppose $\lambda_{j}(n)$ is the eigenvalue of $u_{j}(z)$ for the Hecke operator $T_{n}$. Then the following Hecke relation
$\lambda_{j}(n)\lambda_{j}(m)=\sum_{d|(m,n)}\lambda_{j}\left(\frac{mn}{d^{2}}\right)$ (2.7)
holds. We also have the relation
$\rho_{j}(\pm n)=\rho_{j}(\pm 1)\lambda_{j}(n).$ (2.8)
The Ramanujan–Petersson conjecture: Let us consider the statement that the Hecke eigenvalues at primes $p$ satisfy the bound
$H(\theta):|\lambda_{j}(p)|\leq 2p^{\theta},$ (2.9)
where $\theta$ is a positive real number. It is conjectured that $H(\theta)$ holds for any $\theta>0$.
Remark 2.10.
By the work of Kim and Sarnak [Kim03, Appendix 2], $H_{2}(\frac{7}{64})$ holds.
Lemma 2.11.
We have,
$\frac{|\rho_{j}(1)|^{2}}{\cosh(\pi\kappa_{j})}\ll_{\varepsilon}\kappa_{j}^{\varepsilon}.$
Proof.
See [HL94, Corollary 0.3]. ∎
Lemma 2.12.
(The Kuznetsov formula: opposite sign case) Let $f(t)$ be of $C^{2}$ class with compact support in $(0,\infty)$. Then for $n,m\geq 1$, we have
$\sum_{c}\frac{S(n,-m,c)}{c}f\left(\frac{4\pi\sqrt{mn}}{c}\right)$ $=\sum\limits_{j=1}^{\infty}\rho_{j}(n)\rho_{j}(m)\check{f}(\kappa_{j})+\frac{1}{\pi}\int\limits_{-\infty}^{\infty}(nm)^{-i\eta}\sigma_{2i\eta}(n)\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta,$
where $\check{f}(\eta)$ is the Bessel transform
$\check{f}(\eta)=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta}(t)f(t)\frac{\textnormal{d}t}{t}.$
Proof.
See [DI83, Eq. (1.20), Theorem 1]. ∎
Lemma 2.13.
(The Weyl law) We have,
$N_{SL(2,{\mathbb{Z}})}(K)=\left|\{j:|\kappa_{j}|\leq K\}\right|=\frac{\textnormal{Vol}(SL(2,{\mathbb{Z}})\backslash{\mathbb{H}})}{4\pi}K^{2}+O(K\log K).$
Proof.
See [Iwa02, Eq. (11.5)]. ∎
Lemma 2.14.
For any $\varepsilon>0$ and all $N\geq 1$, we have
$\sum_{n\leq N}|\lambda_{j}(n)|^{2}\ll_{\varepsilon}(\kappa_{j})^{\varepsilon}N.$
Proof.
See [Iwa92, Lemma 1]. ∎
Lemma 2.15.
(The local spectral large sieve inequality) For $1\leq\Delta\leq K$, we have
$\sum_{K\leq\kappa_{j}\leq K+\Delta}\frac{|\rho_{j}(1)|^{2}}{\cosh(\pi\kappa_{j})}\left|\sum_{n\leq N}a_{n}\lambda_{j}(n)\right|\ll(K\Delta+N)\|a\|^{2}(KN)^{\varepsilon}.$
Proof.
See [Jut00, Theorem 1.1]. ∎
Lemma 2.16.
(The large sieve inequality for the twisted coefficients $\rho_{j}(n)n^{i\kappa_{j}}$) For an arbitrary sequence of complex numbers $(a_{n})_{n\leq N}$, we have
$\sum_{0<\kappa_{j}\leq K}\frac{1}{\cosh(\pi\kappa_{j})}\left|\sum_{n\leq N}a_{n}\rho_{j}(n)n^{i\kappa_{j}}\right|^{2}\ll(K^{2}+N^{2})(NK)^{\varepsilon}\sum_{n\leq N}|a_{n}|^{2}.$
Proof.
See [DI86, Theorem 6]. ∎
3. Initial reduction and the main steps of the proof
To prove Theorem 1.3, it is enough to prove that
$S_{r}(X)=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c,d\leq X\\ ad-bc=r\end{subarray}}1=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O(r^{\theta}X^{3/2+\varepsilon}),$ (3.1)
for $0<|r|\ll X^{1/3}$, since $\mathpzc{S}_{r}(X)=8S_{r}(X)$ by symmetry. We first eliminate $d$ and interpret the equality as a congruence modulo $a$, and then express the resulting sum in a suitable form:
$S_{r}(X)$ $=\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c\leq X\\ bc\equiv-r(\textnormal{mod }a)\\ a-r\leq bc\leq aX-r\end{subarray}}1$
$=\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c\leq X\\ bc\equiv-r(\textnormal{mod }a)\\ a\leq bc\leq aX\end{subarray}}1+\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c\leq X\\ bc\equiv-r(\textnormal{mod }a)\\ a-r\leq bc\leq a\end{subarray}}1-\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c\leq X\\ bc\equiv-r(\textnormal{mod }a)\\ aX-r\leq bc\leq aX\end{subarray}}1$
$=S_{r}^{(0)}(X)+S_{r}^{(1)}(X)-S_{r}^{(2)}(X),$
say. We bound the last two sums trivially as follows:
$S_{r}^{(2)}(X)$ $\leq\sum_{1\leq a\leq X}\sum_{\begin{subarray}{c}aX-r\leq n\leq aX\\ n\equiv-r(\textnormal{mod }a)\end{subarray}}\tau(n)$ $\ll X^{\varepsilon}\sum_{1\leq a\leq X}\left(\frac{r}{a}+1\right)\ll rX^{\varepsilon}+O\left(X^{1+\varepsilon}\right),$
and the same bound holds for $S_{r}^{(1)}(X)$ as well. Thus
$S_{r}(X)=S_{r}^{(0)}(X)+O\left((r+X)X^{\varepsilon}\right).$ (3.2)
Now we attach a weight function to $S_{r}^{(0)}(X)$. Let $X>0$ be a large number and let $H$ be a parameter to be fixed later subject to the condition
$\sqrt{X}\leq H\leq X.$ (3.3)
Let $w:{\mathbb{R}}\to{\mathbb{R}}$ be a smooth function satisfying the following conditions:
$(1)\ w(x)=0\textnormal{ if }x\not\in(1-H,X+H),$ (3.4)
$(2)\ w(x)=1\textnormal{ if }x\in[1,X],$ (3.5)
(3) $(3)\ 0<w(x)<1\textnormal{ if }x\in(1-H,1)\cup(X,X+H),$ (3.6)
$(4)\ w^{(j)}(x)\ll_{j}H^{-j}\textnormal{ for every }j\geq 1.$ (3.7)
We define
$S_{w}(X,r)=\mathop{\sum\sum\sum}_{\begin{subarray}{c}bc\equiv-r(\textnormal{mod }a)\\ c\neq 0\\ a>0\end{subarray}}w(a)w(b)w(c)w\left(\frac{bc}{a}\right),$ (3.8)
where $a,b,c$ vary over the set of all integers.
Remark 3.1.
Finally, our choice will be $H=\sqrt{X}$ for the proof of Theorem 1.1.
The support of $w$ ensures that the contribution of the terms in which at least one of the variables $a,b,c,\frac{bc}{a}$ is outside the interval $[1,X]$ is $O\left(HX^{1+\varepsilon}\right)$, and therefore,
$S_{r}^{(0)}(X)=S_{w}(X,r)+O\left(HX^{1+\varepsilon}\right).$ (3.9)
Our task now is to evaluate $S_{w}(X,r)$ asymptotically. Let us first define $l=(a,c),r_{1}=r/l,c_{1}=c/l,a_{1}=a/l.$ Applying Lemma 2.1) to the $b$-sum and separating the zero frequency, we obtain
$S_{w}(X,r)$ $=\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a_{1},c_{1})=1\\ c_{1}\neq 0\\ a_{1}\geq 1\end{subarray}}w(la_{1})w(lc_{1})\sum_{\begin{subarray}{c}b\in{\mathbb{Z}}\\ b\equiv-r_{1}\bar{c_{1}}(\textnormal{mod }a_{1})\end{subarray}}w(b)w\left(\frac{bc_{1}}{a_{1}}\right)$ (3.10)
$=\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a_{1},c_{1})=1\\ c_{1}\neq 0\\ a_{1}>1\end{subarray}}\frac{1}{a_{1}}w(la_{1})w(lc_{1})\widehat{F_{a_{1},c_{1}}}\left(0\right)+\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a_{1},c_{1})=1\\ c_{1}\neq 0\\ a_{1}>1\end{subarray}}\frac{1}{a_{1}}w(la_{1})w(lc_{1})\sum_{\begin{subarray}{c}n\in{\mathbb{Z}}\\ n\neq 0\end{subarray}}e\left(\frac{-nr_{1}\bar{c_{1}}}{a_{1}}\right)\widehat{F_{a_{1},c_{1}}}\left(\frac{n}{a_{1}}\right)$
$=A_{w}(X,r)+B_{w}(X,r),$
say, where we define for any integer $a\neq 0$ and any integer $c$,
$F_{a,c}(x):=w(x)w\left(\frac{cx}{a}\right).$ (3.11)
3.1. The main term
The following proposition shows that we can isolate a main term of size $X^{2}$ from $A_{w}(X,r)$. This will be proved in the next section.
Proposition 3.2.
$A_{w}(X,r)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right).$ (3.12)
3.2. The error term
The following proposition expresses $B_{w}(X,r)$ as a sums of Kloosterman sums twisted by an oscillatory phase function where the sum over the modulus is smooth and thus it can be further analyzed by applying the Kuznetsov formula.
Proposition 3.3.
(The first main proposition)
We have,
$B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{\sqrt{2}\pi}\sum_{l|r}\mathop{\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}}_{nl\leq m}\sum_{a_{1}>0}\frac{S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\sin\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)w\left(\frac{ma_{1}}{n}\right)$
$+\frac{X^{\frac{1}{4}}}{\sqrt{2}\pi}\sum_{l|r}\mathop{\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}}_{m<nl}\sum_{a_{1}>0}\frac{S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\sin\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)w\left(\frac{nl^{2}a_{1}}{m}\right)$
$+\sqrt{2}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{nlX\leq mu}\sum_{a_{1}>0}\frac{S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\cos\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)w\left(\frac{ma_{1}u}{nX}\right)$
$+\sqrt{2}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{mu<nlX}\sum_{a_{1}>0}\frac{S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\cos\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)w\left(\frac{nl^{2}a_{1}X}{mu}\right)$
$+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right).$
The next proposition gives a strong bound for $B_{w}(X,r)$ obtained by applying the spectral methods of automorphic forms.
Proposition 3.4.
(The second main proposition)
We have,
$B_{w}(X,r)$ $\ll X^{\varepsilon}\left[X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}+r^{\theta}\left(X\left(\frac{X}{H}\right)+X^{\frac{1}{2}}\left(\frac{X}{H}\right)^{2}+X^{\frac{1}{3}+\frac{1}{4}}\left(\frac{X}{H}\right)^{\frac{3}{2}}\right)\right].$
In particular, for the choice $H=\sqrt{X}$,
$B_{w}(X,r)\ll r^{\theta}X^{\frac{3}{2}+\varepsilon}.$
3.3. Proof of Theorem 1.1
The theorem follows from Prop. 3.2 and Prop. 3.4, taking into account (3.10), (3.9) and (3.2) after making the choice $H=\sqrt{X}$.
4. The Main term and The Proof of Proposition 3.2
The assumptions on the function $w$ allow us to write
$A_{w}(X,r)$ $=\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a_{1},c_{1})=1\\ a_{1}c_{1}\neq 0\\ a_{1}>0\end{subarray}}\frac{1}{a_{1}}w(la_{1})w(lc_{1})\int\limits_{1}^{X}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x+O\left(H\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a_{1},c_{1})=1\\ a_{1}c_{1}\neq 0\\ a_{1}>0\end{subarray}}\frac{1}{a_{1}}w(la_{1})w(lc_{1})\right),$
where the $O$-term is easily seen to be $O\left(\frac{\sigma(r)}{r}HXX^{1+\varepsilon}\right)$; and as for the first sum, the idea is to isolate the part where $w(c_{1}x/a_{1})=1$. By (3.5), we have,
$w\left(\frac{c_{1}x}{a_{1}}\right)=1\iff\frac{a_{1}}{c_{1}}\leq x\leq\frac{a_{1}X}{c_{1}}.$
We compare this with the range of the $x$-integral and note that
$\left[\frac{a_{1}}{c_{1}},\frac{a_{1}X}{c_{1}}\right]\cap[1,X]=\begin{cases}[1,\frac{a_{1}X}{c_{1}}],&\text{ if }a_{1}\leq c_{1}\\ [\frac{a_{1}}{c_{1}},X],&\text{ if }a_{1}>c_{1}\end{cases}.$
In view of the above observations, it is natural to decompose $A_{w}(X,r)$ as
$A_{w}(X,r)=\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}a_{1}\leq c_{1}\leq\frac{X}{l}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}X}{c_{1}}}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}a_{1}\leq c_{1}\leq\frac{X}{l}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}X}{c_{1}}}^{X}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x$
$+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}}{c_{1}}}^{X}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}}{c_{1}}}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right),$
where the contribution of the terms with $a_{1}>X/l$ is absorbed in the error term. We write the above equation more simply as
$A_{w}(X,r)=A_{1,w}(X)+A_{2,w}(X)+A_{3,w}(X)+A_{4,w}(X)+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right).$ (4.1)
Executing the integral and removing the coprimality condition by Möbius inversion, we easily obtain
$A_{1,w}(X)$ $=\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}a_{1}\leq c_{1}\leq\frac{X}{l}\\ (a_{1},c_{1})=1\end{subarray}}\left(\frac{a_{1}X}{c_{1}}-1\right)=\frac{\sigma(r)}{r}\frac{X^{2}}{\zeta(2)}+O\left(\frac{\sigma(r)}{r}X(\log X)^{2}\right),$
and $A_{3,w}(X)$ satisfies the same asymptotic formula. Now we consider $A_{2,w}(X)$ and $A_{4,w}(X)$. Making the substitution $u=\frac{c_{1}x}{a_{1}}$ in $A_{2,w}(X)$, we note that the length of the $u$-integral is at most $O(H)$ due to the support of $w$ (see (3.4)) and hence $A_{2,w}(X)\ll\frac{\sigma(r)}{r}HX^{1+\varepsilon}$, and the same bound also holds for $A_{4,w}(X)$. This proves Prop. 3.2.
5. The error term: Initial Transformations
Now we begin analyzing the sum $B_{w}(X,r)$. Interchanging the order of summation and applying Lemma 2.2 to the sum over $c_{1}$,
$B_{w}(X,r)$ $=\sum_{l|r}\sum_{a_{1}>1}\frac{w(la_{1})}{a_{1}^{2}}\sum_{\begin{subarray}{c}n\in{\mathbb{Z}}\\ n\neq 0\end{subarray}}\sum_{m\in{\mathbb{Z}}}S(nr_{1},-m,a_{1})\widehat{G}(n/a_{1},m/a_{1}),$ (5.1)
where $G$ is defined by
$G(x,y)=w(x)w(ly)w\left(\frac{xy}{a_{1}}\right).$
The lemma below shows that $\widehat{F_{a_{1},c_{1}}}\left(\frac{n}{a_{1}}\right)$ is negligibly small if $n\gg\frac{X^{1+\varepsilon}}{lH}$ and $\frac{1}{l}\widehat{F_{la_{1},x}}\left(\frac{m}{la_{1}}\right)$ is negligibly small if $m\gg\frac{X^{1+\varepsilon}}{H}$.
Lemma 5.1.
We have
$\widehat{F_{a,c}}(y)\ll_{k}H^{1-k}y^{-k}\left(1+\left|\frac{c}{a}\right|^{k-1}\right),$ (5.2)
for every integer $k\geq 1$, where $F_{a,c}(x)$ is defined in (3.11). In particular, by taking $k=1$, we have the bound
$\widehat{F_{a,c}}(y)\ll_{k}y^{-1}.$ (5.3)
Furthermore. we have the bound
$\widehat{G}(u,v)\ll\min\left\{\frac{X}{v},\frac{X}{lu}\right\},$ (5.4)
where $u$ or $v$ are any two non-zero real numbers.
Proof.
Using (3.11) we integrate by parts $k$ times, getting
$\widehat{F_{a,c}}(y)=\frac{1}{(-2\pi iy)^{k}}\sum_{j=0}^{k}\int\limits_{-\infty}^{\infty}w^{(j)}(x){\left(\frac{c}{a}\right)}^{k-j}w^{(k-j)}\left(\frac{cx}{a}\right)e(-xy)\textnormal{d}x.$
Note that $w^{(k)}\left(\frac{cx}{a}\right)=0$ unless $\frac{cx}{a}\in(1-H,1)\cup(X,X+H)$, and so $w^{(k)}\left(\frac{cx}{a}\right)$ vanishes outside two intervals of length $O\left(\left|\frac{a}{c}\right|H\right)$ each and by (3.7), we obtain the bound $O({\left|\frac{c}{a}\right|}^{k-1}y^{-k}H^{1-k})$ for the term corresponding to $j=0$. Now we consider the other summands. Since $w^{(j)}(x)$ is non-zero only in the intervals $(1-H,1)$ and $(X,X+H)$ for every $j\geq 1$, all these terms together contribute only $O(y^{-k}H^{1-k})$, again by (3.7). Finally, by applying the bound (5.3) in one of the integrals and by trivially bounding the other integral using the support of the function $w$, we obtain the bound (5.4). ∎
Now we isolate the $m=0$ term in $B_{w}(X,r)$ and write
$B_{w}(X,r)=B_{0,w}(X)+B_{1,w}(X),$ (5.5)
where $B_{1,w}(X)$ denotes the sum over all $m\neq 0$.
5.1. Estimation of $B_{0,w}(X)$
Note that the Kloosterman sum appearing in $B_{0,w}(X)$ is nothing but the Ramanujan sum $r_{a_{1}}(nr_{1})$, and by the bound $|r_{q}(n)|\leq(q,n)$ (see [IK04, Eq. (3.5)]), we easily obtain the following:
Proposition 5.2.
We have,
$B_{0,w}(X)\ll X^{1+\varepsilon}.$ (5.6)
5.2. Analysis of $B_{1,w}(X)$: initial steps
At this stage, the Weil bound for Kloosterman sums yields $B_{1,w}(X)\ll\tau(r)X^{5/2+\varepsilon}/H$ and one obtains the asymptotic formula (with the optimal choice $H=X^{3/4}$)
${S_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(X^{7/4+\varepsilon}\right).$ (5.7)
Our goal is to improve upon the size of the error term in terms of $X$. Note that the contribution of the terms $a_{1}$ so that $X<la_{1}\leq X+H$ is quite small. We consider a more general sum
$\mathcal{E}(Y)$ $:=\sum_{l|r}\sum_{0\neq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{0\neq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{\begin{subarray}{c}a_{1}>0\\ la_{1}\in(Y,Y+H]\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}^{2}}\widehat{G}(n/a_{1},m/a_{1}).$
Proposition 5.3.
If $Y\asymp X$, then
$\mathcal{E}(Y)\ll X^{\frac{3}{2}+\varepsilon}.$
Proof.
This follows easily from (5.4) and the Weil bound (Lemma 2.3). ∎
Remark 5.4.
In view of the above proposition, whenever we have a sum over $a_{1}$, we can disregard the portion $\sum_{la_{1}>X}$ by adding the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$. Thus, from now onwards, whenever we encounter a sum of the above form, we may assume that the variable $a_{1}$ lies in the interval $[1,X]$ but we may not display the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ on every occasion since this term is already present in the error term in the final result, i.e., Prop. 3.3.
Now we state a result that will be applied many times. We omit the proof since it follows easily from the Weil bound.
Proposition 5.5.
Let $l\geq 1$, and $\alpha,\beta,\gamma\in{\mathbb{R}}$, such that $\alpha\leq\frac{3}{2}$ and $\beta,\gamma\leq 1$. Then for $0<U\leq X$ we have
$\Delta(\alpha,\beta,\gamma,l,U)$ $:=\sum_{\begin{subarray}{c}0\neq n\ll\frac{X^{1+\varepsilon}}{lH}\end{subarray}}\frac{1}{n^{\gamma}}\sum_{\begin{subarray}{c}0\neq m\ll\frac{X^{1+\varepsilon}}{H}\end{subarray}}\frac{1}{m^{\beta}}\sum_{\begin{subarray}{c}1\leq la_{1}\leq U\end{subarray}}\frac{w(la_{1})\left|S(nr_{1},-m,a_{1})\right|}{a_{1}^{\alpha}}$
$\ll U^{-\alpha+\frac{3}{2}+\varepsilon}\left(\frac{X}{H}\right)^{-\gamma-\beta+2}.$
5.3. Decomposition of $\widehat{G}(n/a_{1},m/a_{1})$
Separating the contribution of $x$ in the range $1\leq x\leq X$, we have
$\widehat{G}(n/a_{1},m/a_{1})=\int\limits_{1}^{X}\int w(ly)w\left(\frac{yx}{a_{1}}\right)e\left(\frac{-my}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\,\textnormal{d}y\,\textnormal{d}x$
$+\left[\int\limits_{1-H}^{1}+\int\limits_{X}^{X+H}\right]w(x)\int w(ly)w\left(\frac{yx}{a_{1}}\right)e\left(\frac{-my}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\,\textnormal{d}y\,\textnormal{d}x.$
By integration by parts in the $y$-integral followed by a trivial estimation of the $x$-integral and noting that the derivative of the function $w$ is supported only on union of two intervals of length $H$ and satisfies the bound $w^{\prime}(t)\ll H^{-1}$, the last two terms are $O(a_{1}H/m)$. Now we decompose the first integral using the property (3.5) of $w$. Note that
$w\left(\frac{yx}{a_{1}}\right)=1\iff\frac{a_{1}}{x}\leq y\leq\frac{a_{1}X}{x},\text{ and }w(ly)=1\iff\frac{1}{l}\leq y\leq\frac{X}{l}.$
Comparing the above ranges of the variable $y$, we see that $w(ly)=w\left(\frac{yx}{a_{1}}\right)=1$ if and only if
$y\in\left[\frac{a_{1}}{x},\frac{a_{1}X}{x}\right]\cap\left[\frac{1}{l},\frac{X}{l}\right]=\begin{cases}[\frac{1}{l},\frac{a_{1}X}{x}],&\text{ if }x>la_{1}\\ [\frac{a_{1}}{x},\frac{X}{l}],&\text{ if }x\leq la_{1}\end{cases}.$
In view of the above observations, we write, using (3.4),
$\widehat{G}(n/a_{1},m/a_{1})=\int\limits_{1}^{la_{1}}\int\limits_{\frac{a_{1}}{x}}^{\frac{X}{l}}e\left(\frac{-nx-my}{a_{1}}\right)\textnormal{d}y\textnormal{d}x+\int\limits_{1}^{la_{1}}\left[\int\limits_{\frac{X}{l}}^{\frac{X+H}{l}}+\int\limits_{\frac{a_{1}}{x}(1-H)}^{\frac{a_{1}}{x}}\right]w(ly)w\left(\frac{xy}{a_{1}}\right)e\left(\frac{-nx-my}{a_{1}}\right)\textnormal{d}y\textnormal{d}x$ (5.8)
$+\int\limits_{la_{1}}^{X}\int\limits^{\frac{a_{1}X}{x}}_{\frac{1}{l}}e\left(\frac{-nx-my}{a_{1}}\right)\textnormal{d}y\textnormal{d}x+\int\limits_{la_{1}}^{X}\left[\int\limits_{\frac{a_{1}X}{x}}^{\frac{a_{1}}{x}(X+H)}+\int\limits_{\frac{1}{l}(1-H)}^{\frac{1}{l}}\right]w(ly)w\left(\frac{xy}{a_{1}}\right)e\left(\frac{-nx-my}{a_{1}}\right)\textnormal{d}y\textnormal{d}x+O\left(\frac{a_{1}H}{m}\right)$
$=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6}+O(a_{1}H/m),$
say. Accordingly, we decompose $B_{1,w}(X)$ as
$B_{1,w}(X)=\sum_{j=1}^{6}B_{1,w}(X,I_{j})+O\left(X^{3/2+\varepsilon}\right),$
where we use Prop. 5.5 to estimate the contribution of the error term. From $I_{4}$ and $I_{5}$ we extract stationary phases for further analysis. The other terms can be handled by the Weil bound and they contribute only to the error term. Thus we obtain the following intermediate proposition, where we have used the notation
$\mathop{\sum_{l}\sum_{m}\sum_{n}}:=\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}.$ (5.9)
Proposition 5.6.
We have,
$B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)$
$-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}$
$+\frac{1}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)$
$+e\left(2\sqrt{\frac{mnu}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right).$
$=T_{1,w}(X)-T_{2,w}(X)+T_{3,w}(X)+T_{4,w}(X)+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right),$
say.
We proceed to prove the proposition below and the proof ends in §7.8.
6. Contributions of $I_{1},I_{2},I_{3},$ and $I_{6}$
Proposition 6.1.
We have,
$(i)B_{1,w}(X,I_{1})\ll\left(X^{\frac{5}{4}}(X/H)^{1/2}+(X/H)^{3}+X^{\frac{3}{2}}\right)X^{\varepsilon},$
$(ii)B_{1,w}(X,I_{3})\ll\left(X^{3/2}+X^{5/2}/H^{2}\right)X^{\varepsilon},$
$(iii)B_{1,w}(X,I_{2}),B_{1,w}(X,I_{6})\ll X^{\frac{3}{2}+\varepsilon}.$
Proof.
It is enough to estimate the integrals $I_{1},I_{2},I_{3},$ and $I_{6}$ since their contributions can be estimated easily by the Weil bound or using Prop. 5.5.
We first prove (iii) and consider only $I_{2}$ since $I_{6}$ can be handled in a similar manner. To bound $I_{2}$, we first interchange the $x$ and $y$ integrals and then use integration by parts in the $x$-integral, obtaining
$I_{2}$ $=-\frac{1}{2\pi i}\frac{a_{1}}{n}\int\limits_{\frac{X}{l}}^{\frac{X+H}{l}}w(ly)e\left(\frac{-my}{a_{1}}\right)\int\limits_{x\leq la_{1}}\left(w^{\prime}(x)w\left(\frac{xy}{a_{1}}\right)+\frac{y}{a_{1}}w(x)w^{\prime}\left(\frac{xy}{a_{1}}\right)\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y+O\left(\frac{a_{1}H}{nl}\right)$
$\ll\frac{a_{1}H}{nl},$
where we have used the bound (3.7) and the fact that $w^{\prime}(x)\neq 0$ only on two intervals of length $H$ (see (3.4), (3.5), and (3.6)).
Now we consider (ii). From (LABEL:hatGintegral), we have
$I_{3}$ $=\int\limits_{1}^{la_{1}}e\left(\frac{-nx}{a_{1}}\right)\int\limits_{\frac{a_{1}}{x}(1-H)}^{\frac{a_{1}}{x}}w(ly)w\left(\frac{xy}{a_{1}}\right)e\left(\frac{-my}{a_{1}}\right)\textnormal{d}y\textnormal{d}x=\int\limits_{1}^{la_{1}}\left[\int\limits_{\frac{a_{1}}{x}(1-H)}^{0}+\int\limits_{0}^{1}+\int\limits_{1}^{\frac{a_{1}}{x}}\right].$
In the resulting three double-integrals, the last two can be bounded trivially and they contribute $O(a_{1}\log X)$. In the first one, let us denote it by ${I_{3}}^{\ast}$, we interchange the $x$ and the $y$ integrals and the change of variable $y\rightarrow-y$ yields
${I_{3}}^{\ast}$ $:=\int\limits_{0}^{a_{1}(H-1)}w(-ly)e\left(\frac{my}{a_{1}}\right)\int\limits_{1}^{\min\{la_{1},a_{1}(H-1)/y)\}}w\left(\frac{-xy}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y$
$=\int\limits_{0}^{(H-1)/l}\int\limits_{1}^{la_{1}}+\int\limits_{(H-1)/l}^{a_{1}(H-1)}\int\limits_{1}^{a_{1}(H-1)/y}={I_{3,1}}^{\ast}+{I_{3,2}}^{\ast},$
say. By integration by parts in the $x$-variable followed by trivial estimation of the $y$-integral, we have
${I_{3,1}}^{\ast}\ll\frac{a_{1}H}{nl}.$
For ${I_{3,2}}^{\ast}$, we note that from the support of the function $w$ (see (3.4)), $w(-ly)=0$ if $y>(H-1)/l$. Hence,
${I_{3,2}}^{\ast}=0.$
Combining the above bounds,
$I_{3}\ll a_{1}\left(\frac{H}{nl}+\log X\right).$
Next we consider (i). We carry out the integration over the $y$-variable (see (LABEL:hatGintegral)) and write, using (3.5),
$I_{1}$ $=-\frac{1}{2\pi i}\frac{a_{1}}{m}\int\limits_{1}^{la_{1}}e\left(\frac{-nx}{a_{1}}\right)\left[e\left(\frac{-mX}{la_{1}}\right)-e\left(\frac{-m}{x}\right)\right]\textnormal{d}x+O\left(\frac{a_{1}H}{m}\right).$
Integrating over $x$ in the first integral, we get
$I_{1}=\frac{1}{2\pi i}\frac{a_{1}}{m}\int\limits_{1}^{la_{1}}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-m}{x}\right)\textnormal{d}x+O\left(\frac{a_{1}H}{m}\right)+O\left(\frac{a_{1}^{2}}{mn}\right).$ (6.1)
The contribution of the $O$-terms is $O(X^{3/2+\varepsilon})$ by Prop. 5.5. We now analyze the exponential integral
$I^{\prime}_{m,n,a_{1}}=\int\limits_{1}^{la_{1}}e(h_{1}(x))\textnormal{d}x,\quad\text{where}\quad h_{1}(x)=-\frac{nx}{a_{1}}-\frac{m}{x}.$
If $mn<0$, there is no stationary point, and $|h^{\prime}_{1}(x)|\geq\frac{n}{a_{1}}$ throughout the interval and hence $I^{\prime}_{m,n,a_{1}}\ll\frac{a_{1}}{n},$ by the first derivative bound (i.e., Lemma 2.5 with $k=1$). If $mn>0$, there is a stationary point given by
$x_{0}=\sqrt{\frac{a_{1}m}{n}}.$
Let us assume further $m,n>0$ since the analysis in the other case is similar and we get the same bounds. We consider two cases.
Case I: $x_{0}\in[2,2la_{1}/3]$. Here we split the interval $[1,la_{1}]$ in three parts as follows:
$I^{\prime}_{m,n,a_{1}}=\left(\int_{1}^{x_{0}-\delta x_{0}}+\int_{x_{0}-\delta x_{0}}^{x_{0}+\delta x_{0}}+\int_{x_{0}+\delta x_{0}}^{la_{1}}\right)e(h_{1}(x))\textnormal{d}x=J_{1}+J_{2}+J_{3},$
say, where $\delta=1/2$. By the first derivative bound, $J_{1},J_{3}\ll a_{1}/n$, and the second derivative bound (i.e., Lemma 2.5 with $k=2$) yields
$J_{2}=\int_{x_{0}-\delta x_{0}}^{x_{0}+\delta x_{0}}e(h_{1}(x))\textnormal{d}x\ll\frac{a_{1}^{\frac{3}{4}}m^{\frac{1}{4}}}{n^{\frac{3}{4}}}.$
Case II: $x_{0}\notin[2,2la_{1}/3]$. In this case, (i) follows from the second derivative bound applied to $I^{\prime}_{m,n,a_{1}}$ and the Weil bound, once we observe that the conditions $x_{0}<2$ and $x_{0}>2la_{1}/3$ correspond to the conditions $a_{1}\leq\frac{4n}{m}$ and $a_{1}\leq\frac{9m}{4nl^{2}}$, respectively. ∎
7. Contributions of $I_{4}$ and $I_{5}$
7.1. Contributions of $I_{4}$
Here we shall assume that $a_{1}$ satisfies $la_{1}\leq X-1$ and this is justified by Remark 5.4. Analogously to (6.1), here we have the expression
$I_{4}=\frac{1}{2\pi i}\frac{a_{1}}{m}\int_{la_{1}}^{X}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mX}{x}\right)\textnormal{d}x+O\left(\frac{a_{1}H}{m}\right)+O\left(\frac{a_{1}^{2}}{mn}\right).$
By Prop. 5.5, the contributions of the $O$-terms are absorbed into the error term. Define
$I_{m,n,a_{1}}:=\int_{la_{1}}^{X}e(h(x))\,\textnormal{d}x,\quad\textnormal{ where }\quad h(x):=-\frac{nx}{a_{1}}-\frac{mX}{x}.$
Differentiating $h(x)$, we get
$h^{\prime}(x)=-\frac{n}{a_{1}}+\frac{mX}{x^{2}},\quad\text{and}\quad h^{\prime\prime}(x)=-\frac{2mX}{x^{3}}.$ (7.1)
If $mn<0$, then there is no stationary point and $|h^{\prime}(x)|\geq\frac{n}{a_{1}}$ throughout the interval and in this case we use the first derivative bound directly to the integral $I_{m,n,a_{1}}$. This gives $I_{m,n,a_{1}}\ll\frac{a_{1}}{n}$ and the contribution of this is sufficiently small by Prop. 5.5. The case left to consider is $mn>0$. In this case the stationary point is given by
$x_{0}=x_{0}(m,n,a_{1}):=\sqrt{\frac{ma_{1}X}{n}}.$ (7.2)
We will assume that $m,n>0$ in our analysis since the other case is identical. Now, if we give an upper bound for exponential integrals and then sum over the three variables $m,n,$ and $a_{1}$ as in the case of $I_{1}$ then the error term we get is not good enough. We, therefore, refine our analysis by localizing the stationary point $x_{0}$ and extracting the stationary phases wherever they occur and estimating the ensuing error terms carefully. This leads to the following:
Proposition 7.1.
We have,
$B_{1,w}(X,I_{4})$ $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}\sum_{a_{1}>0}}_{\sqrt{\frac{ma_{1}X}{n}}\in[la_{1},X]}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\Biggl{\{}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)$
$-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\Biggr{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+\frac{X^{2}}{H}\right)\right).$
Before we begin the proof, let us first observe that from (7.2), we record the following crude range for $x_{0}$: $x_{0}\ll X\sqrt{X/H}\ll X^{5/4}$.
7.2. Localizing the stationary point and decomposition of $B_{1,w}(X,I_{4})$
Let us define, for a fixed $a_{1}$,
$k=k(a_{1})$ $:=\max\left\{j\geq 0:\frac{X}{2^{j}}\geq la_{1}+1\right\}=\left[\frac{\log(X/(la_{1}+1))}{\log 2}\right].$ (7.3)
Now, (7.3) gives the following upper bound on the variable $j$:
for $\text{for }0\leq j\leq k=k(a_{1})\leq J(X),\text{ we have }X/2^{j}\geq la_{1},\text{ where }J(X):=\left[\frac{\log(X/2)}{\log 2}\right].$ (7.4)
We make a dyadic subdivision of the range of $x_{0}$, i.e., the interval $[1,X^{5/4}]$ in the following manner:
$\mathcal{I}_{0}:=\left[X+\delta,X^{\frac{5}{4}}\right],\mathcal{I}_{j}:=\left[\frac{X}{2^{j}}+\delta,\frac{X}{2^{j-1}}-\delta\right]\text{ for }1\leq j\leq k,\text{ and }\mathcal{I}_{k+1}:=[la_{1}+\delta,\frac{X}{2^{k}}-\delta],\mathcal{I}_{k+2}:=[1,la_{1}-\delta].$ (7.5)
Let us define
$\mathcal{I}^{\prime}_{j}:=\left(\frac{X}{2^{j-1}}-\delta,\frac{X}{2^{j-1}}+\delta\right)\text{ for }1\leq j\leq k+1,\mathcal{I}^{\prime}_{k+2}:=(la_{1}-\delta,la_{1}+\delta).$ (7.6)
Here $\delta>0$ is to be chosen later (we will see that $\delta=1/100$ suffices). Clearly, we can write the interval $[1,X^{5/4}]$ as a disjoint union of the $\mathcal{I}_{j}$’s and $\mathcal{I}^{\prime}_{j}$s, and thus we express $B_{1,w}(X,I_{4})$ as
$B_{1,w}(X,I_{4})$ $=\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{0\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$ (7.7)
$+\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{1\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}^{\prime}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$
$=\mathcal{S}+\mathcal{T},$
say.
7.3. Subdividing the range of integration and decomposition of the sum $\mathcal{S}$
We first subdivide the range of the integral $[la_{1},X]$. Let
$\mathcal{F}_{j}:=\left[\frac{X}{2^{j}},\frac{X}{2^{j-1}}\right],\text{ for }1\leq j\leq k\text{ and }\mathcal{F}_{k+1}:;=\left[la_{1},\frac{X}{2^{k}}\right].$ (7.8)
Clearly, $[la_{1},X]=\bigcup\limits_{1\leq j\leq k+1}\mathcal{F}_{j},$ where an interval $\mathcal{F}_{j}$ can intersect its neighbour only at the boundary point. Let us define
$y_{j}:=\frac{X}{2^{j}},\text{ for }0\leq j\leq k+1,\text{ and }y_{k+2}:=\frac{la_{1}}{2}.$ (7.9)
From (7.3) and (7.9), we have,
$y_{k+1}\leq la_{1}\leq y_{k}.$ (7.10)
Note that for $11\leq j\leq k+$ 1, the interval $\mathcal{I}_{j}$ is contained in $\mathcal{F}_{j}$ and the edges of $\mathcal{I}_{j}$ are away from the edges of $\mathcal{F}_{j}$. For isolating the stationary phase when the stationary point $x_{0}\in\mathcal{I}_{j}$ we express the sum $\mathcal{S}$ as
$\mathcal{S}$ $=\frac{1}{2\pi i}\sum_{l|r}\mathop{\sum_{a_{1}>0}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{0}}\frac{1}{m}\frac{S(nr_{1},-m,a_{1})w(la_{1})}{a_{1}}\int_{la_{1}}^{y_{0}}e(h(x))\textnormal{d}x$ (7.11)
$+\frac{1}{2\pi i}\sum_{l|r}\mathop{\sum_{a_{1}>0}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{1}}\frac{1}{m}\frac{S(nr_{1},-m,a_{1})w(la_{1})}{a_{1}}\left[\int_{la_{1}}^{y_{1}}+\int_{\mathcal{F}_{1}}\right]e(h(x))\textnormal{d}x$
$+\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{1<j\leq k}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}}\frac{1}{m}\frac{S(nr_{1},-m,a_{1})w(la_{1})}{a_{1}}\left[\int_{la_{1}}^{y_{j}}+\int_{\mathcal{F}_{j}}+\int_{2y_{j}}^{X}\right]e(h(x))\textnormal{d}x$
$+\frac{1}{2\pi i}\sum_{l|r}\mathop{\sum_{a_{1}>0}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{k+1}}\frac{1}{m}\frac{S(nr_{1},-m,a_{1})w(la_{1})}{a_{1}}\left[\int_{\mathcal{F}_{k+1}}+\int_{2y_{k+1}}^{X}\right]e(h(x))\textnormal{d}x$
$+\frac{1}{2\pi i}\sum_{l|r}\mathop{\sum_{a_{1}>0}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{k+2}}\frac{1}{m}\frac{S(nr_{1},-m,a_{1})w(la_{1})}{a_{1}}\int_{2y_{k+2}}^{X}e(h(x))\textnormal{d}x$
$=\mathcal{S}_{1}+(\mathcal{S}_{2}+\mathcal{S}_{3})+(\mathcal{S}_{4}+\mathcal{S}_{5}+\mathcal{S}_{6})+(\mathcal{S}_{7}+\mathcal{S}_{8})+\mathcal{S}_{9},$
say.
Remark 7.2.
Note that $la_{1}>X/2$ iff $k=0$ and in that case, the only sums that occur in the right hand side of the above equation are $\mathcal{S}_{1}$, $\mathcal{S}_{8}$ and $\mathcal{S}_{9}$.
7.4. Estimation of $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, $\mathcal{S}_{4}$, $\mathcal{S}_{6}$, $\mathcal{S}_{8}$ and $\mathcal{S}_{9}$
Let $Q_{1}(X):=\mathcal{S}_{1}+\mathcal{S}_{2}+\mathcal{S}_{4}$ and $Q_{2}(X):=\mathcal{S}_{6}+\mathcal{S}_{8}+\mathcal{S}_{9}$.
Proposition 7.3.
We have,
$Q_{1}(X),Q_{2}(X)\ll X^{\frac{3}{2}+\varepsilon}.$
Proof.
We only consider the case of $Q_{1}(X)$ as $Q_{2}(X)$ can be tackled in the same way. We have,
$Q_{1}(X)=\frac{1}{2\pi i}\sum_{l|r}\mathop{\sum_{a_{1}>0}\sum_{0\leq j\leq k}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}}_{x_{0}\in\mathcal{I}_{j}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I^{+}_{m,n,a_{1}}(y_{j}),$ (7.12)
where
$I^{+}_{m,n,a_{1}}(y_{j}):=\int\limits_{la_{1}}^{y_{j}}e(h(x))\textnormal{d}x.$
Here the main idea is to subdivide the ranges of $m$, $n$ and $a_{1}$ according to the position of $x_{0}$. First we interchange the sum over $j$ and the sum over $a_{1}$ using (7.4) and (7.9). Then we bring the sum over $j$ outside the sums over the variables $m$ and $n$. Thus we have,
$Q_{1}(X)\ll\sum_{l|r}\sum_{1\leq j\leq J(X)}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{\begin{subarray}{c}0<la_{1}\leq y_{j}\\ \sqrt{\frac{ma_{1}X}{n}}>y_{j}+\delta\end{subarray}}\frac{1}{m}\frac{|S(nr_{1},-m,a_{1})|w(la_{1})}{a_{1}}|I_{m,n,a_{1}}^{+}(y_{j})|.$ (7.13)
Note that the sum over $a_{1}$ is vacuous if $j$ is so large that $X/2^{j}<la_{1}$. We now make a dyadic subdivision of the range of $x_{0}$ and write
$[y_{j}+\delta,X^{3/2}]=\bigcup_{\begin{subarray}{c}Y=2^{j}\delta\\ 0\leq j\leq L(X)\end{subarray}}[y_{j}+Y,y_{j}+2Y],$
where $L(X)=O(\log X)$. We consider two cases.
Case I: $\delta\leq Y\leq\frac{y_{j}}{2}$. Suppose $y_{j}+Y\leq x_{0}\leq y_{j}+2Y,$ where $\delta\leq Y\leq\frac{y_{j}}{2}$. This implies
$\frac{n}{m}\leq\frac{a_{1}X}{(y_{j}+Y)^{2}}\ll\frac{1}{l}\left(\frac{X}{y_{j}}\right)^{2},$ (7.14)
and
$\frac{n}{mX}(y_{j}+Y)^{2}\leq a_{1}\leq\frac{n}{mX}(y_{j}+2Y)^{2},$ (7.15)
and hence, the size of $a_{1}$ is given by $a_{1}\asymp\frac{n{y_{j}}^{2}}{mX}$ and the number of such $a_{1}$ is $O\left(\frac{ny_{j}Y}{mX}+1\right)$. Moreover, from (7.15)
$\frac{n}{a_{1}}\left(\frac{(y_{j}+Y)^{2}}{y_{j}^{2}}-1\right)\leq\frac{mX}{y_{j}^{2}}-\frac{n}{a_{1}}\leq\frac{n}{a_{1}}\left(\frac{(y_{j}+2Y)^{2}}{y_{j}^{2}}-1\right).$ (7.16)
Hence, by the first derivative bound, (7.1) and the first inequality of the last equation, it follows that
$I^{+}_{m,n,a_{1}}(y_{j})\ll\frac{a_{1}y_{j}}{nY}.$ (7.17)
Note that from (7.14), we have the condition $nl\ll m(X/y_{j})^{2}$. Therefore, using (7.13), the contribution of the terms considered in Case I is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$, by the Weil bound, (7.17) and (7.14), where we have extended the definition $y_{j}=X/2^{j}$ to $j\geq k+3$ by positivity.
Case II: $Y>\frac{y_{j}}{2}$. By (7.1), the first inequality of (7.16) and Prop. 5.5, this case contributes $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ in $Q_{1}(X)$ (see (7.12)). Combining this with the bound in Case I, we prove Prop. 7.3. ∎
7.5. Estimation of $\mathcal{T}$
To estimate $\mathcal{T}$ (see (7.7)), we take advantage of the fact that the length of each of the intervals is small: $\mathcal{L}(\mathcal{I}_{j}^{\prime})=\delta$.
Proposition 7.4.
For $\delta=\frac{1}{100}$, we have,
$\mathcal{T}\ll X^{\varepsilon}\left(\sqrt{X}\left(\frac{X}{H}\right)^{3/2}+X^{3/2}\right).$
Proof.
Recalling (7.2), (7.6), (7.7) and (7.9), we write $\mathcal{T}$ as
$\mathcal{T}=\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{1\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{\sqrt{\frac{ma_{1}X}{n}}\in(2y_{j}-\delta,2y_{j}+\delta)}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left[\int\limits_{la_{1}}^{y_{j}}+\int\limits_{y_{j}}^{4y_{j}}+\int\limits_{4y_{j}}^{X}\right]e(h(x))\textnormal{d}x$
$=\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3},$
say. By an argument similar to the one used in proving Prop. 7.3, we can show that
$\mathcal{T}_{1},\mathcal{T}_{3}\ll X^{3/2+\varepsilon}.$
For the estimation of $\mathcal{T}_{2}$, we bring the $j$-sum outside using (7.4) and (7.9) as before. Thus, we obtain
$\mathcal{T}_{2}\ll\sum_{l|r}\sum_{1\leq j\leq J(X)+2}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{\begin{subarray}{c}0<la_{1}\leq 2y_{j}\\ \sqrt{\frac{ma_{1}X}{n}}\in(2y_{j}-\delta,2y_{j}+\delta)\end{subarray}}\frac{w(la_{1})}{m}\frac{|S(nr_{1},-m,a_{1})|}{a_{1}}\left|\int\limits_{y_{j}}^{4y_{j}}e(h(x))\textnormal{d}x\right|.$
Since $x_{0}\asymp y_{j}$, by Lemma 2.5 with $k=2$, and recalling (7.1),
$\int\limits_{y_{j}}^{4y_{j}}e(h(x))\textnormal{d}x\ll\frac{x_{0}^{3/2}}{\sqrt{mX}}=\frac{m^{\frac{1}{4}}X^{\frac{1}{4}}a_{1}^{\frac{3}{4}}}{\sqrt{2}n^{\frac{3}{4}}}.$ (7.18)
Now the location of $x_{0}$ determines the size of $a_{1}$ (as in (7.15)) and now, by an application of the Weil bound, we get
$\mathcal{T}_{2}\ll\sqrt{X}\left(\frac{X}{H}\right)^{3/2+\varepsilon}.$
7.6. Analysis of $\mathcal{S}_{3}$, $\mathcal{S}_{5}$, $\mathcal{S}_{7}$ and Extraction of The Stationary Phase
Since $x_{0}\in\mathcal{I}_{j}$, and $\mathcal{I}_{j}\subset\mathcal{F}_{j}$, we extract the stationary phases by applying Lemma 2.6 and thus obtain that for $11\leq j\leq k+$ 1,
$\int\limits_{\mathcal{F}_{j}}e(h(x))\textnormal{d}x=\frac{m^{\frac{1}{4}}X^{\frac{1}{4}}a_{1}^{\frac{3}{4}}}{\sqrt{2}n^{\frac{3}{4}}}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)+O(R_{1})+O(R_{2}),$ (7.19)
where
$R_{1}=\frac{1}{\lambda_{2}(x_{0}-y_{j})}+\frac{1}{\lambda_{2}(2y_{j}-x_{0})},\text{ and }R_{2}=\frac{\mathcal{L}(\mathcal{F}_{j})\lambda_{4}}{\lambda_{2}^{2}}+\frac{\mathcal{L}(\mathcal{F}_{j})\lambda_{3}^{2}}{\lambda_{2}^{3}},$
except in case of $\mathcal{F}_{k+1}$. In this case, the lower end point of the integral is $la_{1}$, and consequently
$R_{1}=\frac{1}{\lambda_{2}(x_{0}-la_{1})}+\frac{1}{\lambda_{2}(2y_{k+1}-x_{0})}$
and everything else remains unchanged. As in the proof of Prop. 7.3, by introducing dyadic subdivision of the range of the stationary point and analyzing the ensuing restrictions on the variables $m,n$ and $a_{1}$, we show that the contributions of the error terms in $\mathcal{S}_{3}$, $\mathcal{S}_{5}$ and $\mathcal{S}_{7}$ is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the Weil bound. Furthermore, from (7.19), we observe that the stationary phase analysis of the integrals in $\mathcal{S}_{3},\mathcal{S}_{5},\mathcal{S}_{7}$ give the same main term. Combining all these estimates in one place, we conclude that
$Q_{3}(X):=\mathcal{S}_{3}+\mathcal{S}_{5}+\mathcal{S}_{7}$
$=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}\sum_{a_{1}>0}}_{x_{0}\in\cup_{1\leq j\leq k+1}\mathcal{I}_{j}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)+O\left(X^{\frac{3}{2}+\varepsilon}\right).$
Now we extend the range of $x_{0}$ to the whole interval $[la_{1},X]$ and subtract the contribution of the extra terms that have added. Note that the total contribution of these extra terms is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the same kind of argument used in the proof of Prop. 7.4. Thus, we have
$Q_{3}(X)$ $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}\sum_{a_{1}>0}}_{\sqrt{\frac{ma_{1}X}{n}}\in[la_{1},X]}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)+O\left(X^{\frac{3}{2}+\varepsilon}\right).$ (7.20)
Combining Prop. 7.3, Prop. 7.4, (7.20) and also considering the case $m,n<0$, we prove Prop. 7.1.
7.7. Contribution of $I_{5}$
Proposition 7.5.
We have,
$B_{1,w}(X,I_{5})$ $=\frac{1}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}\sum_{a_{1}>0}}_{la_{1}\leq\sqrt{\frac{ma_{1}u}{n}}\leq X}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)$
$+e\left(2\sqrt{\frac{mnu}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}+O\left(X^{3/2+\varepsilon}\right).$
Proof.
Here we only give a sketch as the proof is very similar to the proof of Prop. 7.1. From (LABEL:hatGintegral) and by change of variable $y=ua_{1}/x$, we have
$I_{5}$ $=a_{1}\int\limits_{X}^{X+H}w(u)\int\limits_{la_{1}}^{X}\frac{1}{x}w\left(\frac{a_{1}ul}{x}\right)e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mu}{x}\right)\textnormal{d}x\textnormal{d}u.$
Next, by (3.5) of $w$, we have
$w\left(\frac{la_{1}u}{x}\right)=1\iff\frac{la_{1}u}{X}\leq x\leq la_{1}u,\quad\text{and}\quad\bigg{[}\frac{la_{1}u}{X},la_{1}u\bigg{]}\bigcap\bigg{[}la_{1},X\bigg{]}=\bigg{[}\frac{la_{1}u}{X},X\bigg{]},$
since $u\geq X$. Thus we write $I_{5}$ as
$I_{5}$ $=a_{1}\int\limits_{X}^{X+H}w(u)\left[\int\limits_{la_{1}u/X}^{X}\frac{1}{x}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mu}{x}\right)+\int\limits_{la_{1}}^{la_{1}u/X}\frac{1}{x}w\left(\frac{a_{1}ul}{x}\right)e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mu}{x}\right)\right]\textnormal{d}x\textnormal{d}u.$
Let us consider the second integral first. Since $u\leq X+H$, the length of the $x$-integral is bounded by $la_{1}H/X$. Thus the $x$-integral is $O(H/X)$ and hence the second integral is $(a_{1}H^{2}/X)$. For the first integral, we write
$\int\limits_{la_{1}u/X}^{X}=\int\limits_{la_{1}}^{X}-\int\limits_{la_{1}}^{la_{1}u/X}.$
Estimating the second integral trivially,
$\int\limits_{la_{1}}^{la_{1}u/X}\frac{1}{x}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mu}{x}\right)\textnormal{d}x\ll u/X-1=H/X,$
and hence estimating the $u$-integral trivially, we finally get
$I_{5}=a_{1}\int\limits_{X}^{X+H}w(u)\int\limits_{la_{1}}^{X}\frac{1}{x}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mu}{x}\right)\textnormal{d}x\textnormal{d}u+O\left(\frac{a_{1}H^{2}}{X}\right).$ (7.21)
By Prop. 5.5, the contribution of the $O$-term in $B_{1,w}(X)$ is $O\left(X^{3/2+\varepsilon}\right)$.
Let us now compare the above expression for $I_{5}$ with that of $I_{4}$. The outside integral over $u$ plays practically no role and we can insert the sum over $m,n,$ and $a_{1}$ inside and treat $u$ as a constant of size $X$. Thus our phase function is essentially the same as before except that $X$ has been replaced by $u$. However, when we execute the outside integral at the end, this gives rise to a factor of size $H$. Also, there is a factor $1/m$ in the case of $I_{4}$ which is not present here and there is an extra factor $1/x$ in $I_{5}$ where $x$ varies from $la_{1}$ to $X$. In the crucial ranges of the variables; i.e., for $a_{1}\asymp X/l$ and $m\asymp X/H$, we see that the two integrals $I_{4}$ and $I_{5}$ have the same shape. The same analysis that was carried out in the case of $I_{4}$7.1 to §7.6) applies here too, with the same kind of subdivisions of the range of the stationary point $x_{0}=\sqrt{\frac{ma_{1}u}{n}}$ and the range of the $x$-integral. However, due to the presence of the factor $1/x$, we need to use Lemma 2.7 instead of Lemma 2.6. ∎
7.8. Proof of Proposition 5.6
Bringing together Prop. 6.1, Prop. 7.1 and Prop. 7.5, we arrive at Prop. 5.6, except for the conditions on $a_{1}$. Now, the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ appearing in the sum over $a_{1}$ in Prop. 7.1 is equivalent to saying $0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m},\frac{X}{l}\}$. However, we can drop the condition $a_{1}\leq\frac{X}{l}$ because of the presence of the factor $w(la_{1})$ in our sum and in view of Remark 5.4. A similar consideration shows that the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ is equivalent to saying $0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}$.
Remark 7.6.
At this stage, a direct application of the Weil bound for Kloosterman sum (see Lemma 2.3), yields the bound
$B_{w}(X,r)\ll X^{\varepsilon}\left(\frac{X^{2}}{\sqrt{H}}+X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right).$
Now, from (3.2), (3.9), (3.10), and Prop. 3.2, the optimal choice for $H$ is given by $H=X^{2/3}$ and this gives us
${S_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(X^{5/3+\varepsilon}\right).$
8. Removal of the sharp-cut condition and the proof of Proposition 3.3
To simplify notation, let us, henceforth, denote the variable $a_{1}$ by $c$.
In Prop. 5.6, we expressed $\widehat{G}(n/c,m/c)$ in a form that enabled us to see its oscillating behaviour and its order of magnitude. The next step is to apply the Kuznetsov formula to the sums over the variable $c$, but first we need to make the variable smooth. We show how to do it for $T_{1,w}(X)$. The condition $c\leq\min\left\{\frac{mX}{nl^{2}},\frac{nX}{m}\right\}$ can be restated as
$c\leq\frac{mX}{nl^{2}}\textnormal{ when }m\leq nl,\textnormal{ and }c\leq\frac{nX}{m}\textnormal{ when }nl\leq m.$ (8.1)
We have,
Proposition 8.1.
$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{0<c\leq\frac{nX}{m}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{mc}{n}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
An analogous statement holds in the complimentary situation, i.e.,
$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m\leq nl}\sum_{0<c\leq\frac{mX}{nl^{2}}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m\leq nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{cnl^{2}}{m}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
Proof.
In the case $nl\leq m$, $nX/m\leq X/l$ and hence, by comparing the range of $c$ with the support of $w$ (see (3.4), (3.5) and (3.6)), we have (suppressing the sums over $l,m,n$)
$\sum_{0<c\leq\frac{nX}{m}}w(lc)\cdots=\sum_{0<c\leq\frac{nX}{m}}1.\cdots+O\left(\sum_{\frac{X}{l}\leq c\leq\frac{X+H}{l}}w(lc)\cdots\right),$
and the error term here is $O\left(X^{1+\varepsilon}\sqrt{H}\right)$ by the Weil bound. We remove the sharp-cut condition $c\leq nX/m$ in the first sum by reintroducing the function $w$, up to an acceptable error term as follows:
$\sum_{0<c<\frac{nX}{m}}1.\cdots=\sum_{c>0}w\left(\frac{mc}{n}\right)\cdots+O\left(\sum_{X<\frac{mc}{n}\leq X+H)}w\left(\frac{mc}{n}\right)\cdots\right).$
One notes that this error term is also $O\left(X^{1+\varepsilon}\sqrt{H}\right)$ by the Weil bound. This finishes the proof in this case. The proof in the other case is similar and we omit it. ∎
Thus we write $T_{1,w}(X)$ as:
$T_{1,w}(X)=R_{1,w}(X)+{R^{\prime}}_{1,w}(X)+O\left(X^{1+\varepsilon}\sqrt{H}\right),$ (8.2)
where
$R_{1,w}(X)$ $:=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right),$
and
${R^{\prime}}_{1,w}(X)$ $:=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m<nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{nl^{2}c}{m}\right).$
We can decompose $T_{3,w}(X)$ in a similar manner as
$T_{3,w}(X)=R_{3,w}(X)+{R^{\prime}}_{3,w}(X)+O\left(X^{1+\varepsilon}\sqrt{H}\right),$ (8.3)
where
$R_{3,w}(X)$ $:=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right),$ (8.4)
and
${R^{\prime}}_{3,w}(X)$ $:=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{nl^{2}c}{m}\frac{X}{u}\right).$
The sum $T_{2,w}(X)$ (resp. $T_{4,w}(X)$) is exactly analogous to $T_{1,w}(X)$ (resp. $T_{3,w}(X)$) and we have similar decompositions for $T_{2,w}(X)$ and $T_{4,w}(X)$. Combining these expressions for $T_{i,w}(X),i=1,2,3,4$, we obtain Prop. 3.3.
9. Proof of Proposition 3.4: initial steps
We only discuss the cases of $T_{1,w}(X)$ and $T_{3,w}(X)$ below as the other two sums can be handled similarly. Moreover, it is enough to consider $R_{1,w}(X)$ and $R_{3,w}(X)$ as ${R^{\prime}}_{1,w}(X)$ and ${R^{\prime}}_{3,w}(X)$ are similar to $R_{1,w}(X)$ and $R_{3,w}(X)$, respectively.
9.1. Analysis of $R_{1,w}(X)$
We introduce smooth dyadic partitions of unity to the sums over $m,n,$ and $c$ in $R_{1,w}(X)$. Let $b_{1},b_{2},$ and $V$ be smooth functions with support inside the interval $[1,4]$ such that
$\sum_{j\geq 0}V(x/2^{j})=\sum_{j\geq 0}b_{i}(x/2^{j})=1\ (i=1,2),\quad\text{and}\quad b_{i}^{(k)}(x),V^{(k)}(x)\ll_{k}1\ (i=1,2)$
for every $x>0$. Thus we have,
$R_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{C\text{ dyadic}}}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (9.1)
×$\times\sum_{c}V\left(\frac{c}{C}\right)\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right).$
When $C$ is small enough, the Weil bound already gives a good bound (see Prop. 5.5) and we collect all such terms in the error term and write
$R_{1,w}(X)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{\begin{subarray}{c}C\text{ dyadic}\\ C\geq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\end{subarray}}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots+O\left(X^{\frac{5}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{1}{2}}\right).$ (9.2)
Now we fix a divisor $l$ of $r$ and consider the sum
$R_{1}(M,N,C,X):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right)V\left(\frac{c}{C}\right)g(m,n,c),$
where
$g(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{3}{4}}.$ (9.3)
We define
$C_{0}:=\max\left\{2^{j}:2^{j}\leq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\right\},$ (9.4)
so that the contribution of the terms corresponding to $C\leq C_{0}$ is absorbed in the error term after we make the choice $H=\sqrt{X}$ at the end. Note that in the sum $R_{1}(M,N,C,X)$, the dyadic variables $M$ and $N$ satisfy $1\leq M\ll\frac{X^{1+\varepsilon}}{H}$, $1\leq N\ll\frac{X^{1+\varepsilon}}{lH}$ and from the support of the function $w$ (see (3.6)), we have $C\leq\frac{8NX}{M}$ as $N\leq n\leq 4N$ and $M\leq m\leq 4M$. For estimating $R_{1,w}(X)$, it is enough to estimate $R_{1}(M,N,C,X)$ and here is the main result we need.
Proposition 9.1.
For fixed $M,N,$ and $C$ such that
$C_{0}\leq C\leq\frac{8NX}{M},1\leq M\ll\frac{X^{1+\varepsilon}}{H},\textnormal{ and }1\leq N\ll\frac{X^{1+\varepsilon}}{lH},$ (9.5)
we have
$R_{1}(M,N,C,X)\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}+r^{\theta}X^{\frac{3}{4}+\varepsilon}N$
$+r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left[MN\left(\sqrt{\frac{MC}{NH}}+\sqrt{M}+\sqrt{N}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}+\sqrt{\frac{X}{H}}\right)+(MN)^{\frac{3}{4}}M\right]+r_{1}^{\theta}X^{3/4+\varepsilon}M^{23/32}.$
In particular,
$R_{1}(M,N,C,X)\ll r^{\theta}X^{\frac{3}{4}+\varepsilon}\left(\frac{X}{H}\right).$
9.2. Analysis of $R_{3,w}(X)$
The analysis of this sum is similar to that of $R_{1,w}(X)$. Note that the variable $u$ in the sum $R_{3,w}(X)$ plays a role analogous to that of the parameter $X$ inside the phase function in $T_{1,w}(X)$ and we can work with a fixed value of $u$ and then integrate trivially. Moreover, the size of $u$ is the same as that of $X$, in fact $u=X+O(H)$ and $H$ is chosen to be $\sqrt{X}$ in the end. Therefore, the $u$-integral outside contributes a factor of size $O(X^{-1/4}H)$ which has the same effect as multiplying $X^{1/4}$ to the inner sum over $l,m,n,$ and $c$. Thus, the essential difference between $R_{1,w}(X)$ and $R_{3,w}(X)$ is that in $R_{3,w}(X)$, we have the term $(c/mn)^{1/4}$ in the place of $(c/mn)^{3/4}$ in $R_{1,w}(X)$. Note that the condition $nl\leq m$ is now replaced by $nl\leq mu/X$, but since $u$ and $X$ are of the same size, this change does not make much of a difference. The same comment applies to the argument of the function $w$. We estimate the initial part of the sum over $c$ by the Weil bound and we truncate the sum at the point
$C_{0}^{\star}:=\max\left\{2^{j}:2^{j}\leq\frac{X^{\frac{2}{3}+\varepsilon}}{l}\right\},$ (9.6)
for a suitable $\varepsilon>0$. By Prop. 5.5, we have,
$\frac{e\left(\frac{1}{8}\right)}{\sqrt{2}}$ $\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{0<c\leq C_{0}^{\star}}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)\ll X^{\frac{3}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{3}{2}}.$ (9.7)
The choice of $C_{0}^{\star}$ is dictated by the above bound, since the optimal choice for $H$ is $\sqrt{X}$. For larger values of $c$ we follow the same approach as above and we get the following sum after dyadic subdivision which is analogous to $R_{1}(M,N,C,X)$:
$R_{3}(M,N,C,X,u)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)V\left(\frac{c}{C}\right)g^{\star}(m,n,c),$
where
$g^{\star}(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{1}{4}}.$ (9.8)
We then have the following proposition which can be compared with Prop. 9.1.
Proposition 9.2.
For fixed $M,N,$ and $C$ such that
$C_{0}^{\star}\leq C\leq\frac{8NX}{M},1\leq M\ll\frac{X^{1+\varepsilon}}{H},\textnormal{ and }1\leq N\ll\frac{X^{1+\varepsilon}}{lH},$
and for a fixed $u\in[X,X+H]$, where $\sqrt{X}\leq H\leq X$, we have
$R_{3}(M,N,C,X,u)\ll r^{\theta}X^{\varepsilon}\left(X^{\frac{1}{4}}\left(\frac{X}{H}\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right).$
We give the details of the proof of Prop. 9.1 below and the proof of Prop. 9.2 is almost identical and we only give a very brief sketch of it at the end.
10. Application of The Kuznetsov formula
10.1. Decomposition of $R_{1}(M,N,C,X)$
Let us define
$f_{m,n,C}(t):=w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(-\frac{X^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right),$ (10.1)
where $g(m,n,c)$ is defined in (9.3). Applying the Kuznetsov formula (Lemma 2.12) to the $c$-sum inside $R_{1}(M,N,C,X)$, we obtain
$R_{1}(M,N,C,X)=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C}(\kappa_{j})$
$+\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\frac{1}{\pi}\int\limits_{-\infty}^{\infty}(nmr_{1})^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}_{m,n,C}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta$
$=\Sigma_{\text{disc.}}+\Sigma_{\text{cont.}},$
say, where
$\check{f}_{m,n,C}(\eta)=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$
From the support of $V$, the argument $t$ of $f_{m,n,C}(t)$ (see (10.1)) is very small. Indeed, it lies in the the interval
$\frac{\pi\sqrt{mnr_{1}}}{C}<t\leq\frac{4\pi\sqrt{mnr_{1}}}{C}\leq\frac{16\pi\sqrt{MNr_{1}}}{C}\ll 1,$ (10.2)
since $C>C_{0}$ (see (9.4)), $H\geq\sqrt{X}$ (see (3.3)) and $r=O(X^{1/3})$ (see Remark 16.1). We now expand $K_{2i\eta}(t)$ as a power series:
$K_{2i\eta}(t)$ $=\frac{\pi}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\left(\frac{\left(\frac{t}{2}\right)^{-2i\eta+2\nu}}{\Gamma(1-2i\eta)\nu!(1-2i\eta)\dots(\nu-2i\eta)}-\frac{\left(\frac{t}{2}\right)^{2i\eta+2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}\right)$
$=K_{2i\eta,+}(t)-K_{2i\eta,-}(t),$
say. Accordingly, we write
$\Sigma_{\text{disc.}}=\mathcal{D}_{+}(M,N,C)-\mathcal{D}_{-}(M,N,C),\quad\text{and}\quad\Sigma_{\text{cont.}}=\mathcal{C}_{+}(M,N,C)-\mathcal{C}_{-}(M,N,C),$
where
$\mathcal{D}_{\pm}(M,N,C):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,\pm}(\kappa_{j}),$ (10.3)
$\mathcal{C}_{\pm}(M,N,C)$ $:=\frac{1}{\pi}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{-\infty}^{\infty}(nmr_{1})^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}_{m,n,C,\pm}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta,$ (10.4)
and
$\check{f}_{m,n,C,\pm}(\eta)$ $:=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta,\pm}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$ (10.5)
Thus we decompose $R_{1}(M,N,C,X)$ as
$R_{1}(M,N,C,X)=\mathcal{D}_{+}(M,N,C)-\mathcal{D}_{-}(M,N,C)+\mathcal{C}_{+}(M,N,C)-\mathcal{C}_{-}(M,N,C),$ (10.6)
and our goal now is to estimate the above four sums separately. First we analyze $\check{f}_{m,n,C,\pm}(\eta)$, by writing it more explicitly as follows:
$\check{f}_{m,n,C,\pm}(\eta)$ $=\frac{4}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\Gamma(1\mp 2i\eta)\nu!(1\mp 2i\eta)\dots(\nu\mp 2i\eta)}I_{\pm}(m,n,C,\nu,\eta,r_{1}),$ (10.7)
where
$I_{\pm}(m,n,C,\nu,\eta,r_{1})$ $:=\int t^{\mp 2i\eta+2\nu}w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(\frac{-X^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right)\frac{\textnormal{d}t}{t}.$
We collect various estimates for the above integrals. By the change of variables $t=u^{2}\sqrt{r_{1}}$,
$I_{\pm}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu\mp i\eta}\int A(u)e(-h_{\pm}(u))\frac{\textnormal{d}u}{u},$ (10.8)
where
$A(u)=A_{m,n,\nu,C}(u):=\frac{u^{4\nu}}{u}w\left(\frac{4\pi m\sqrt{mn}}{nu^{2}}\right)V\left(\frac{4\pi\sqrt{mn}}{u^{2}C}\right)g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right),$ (10.9)
and
$h_{\pm}(u):=\frac{X^{\frac{1}{2}}u(mn)^{\frac{1}{4}}}{\sqrt{\pi}}\pm\frac{2\eta\log u}{\pi}.$ (10.10)
Recall that $g(m,n,c)=(\frac{c}{mn})^{\frac{3}{4}}$ and under the condition $nl\leq m$, we have the bound $g(m,n,c)\leq(\frac{c}{n^{2}l})^{\frac{3}{4}}$. We record the following bounds that will be required later:
$g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right)\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}},\text{ and }\|A\|_{\infty}\ll\frac{\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}}{\left(\frac{\sqrt{MN}}{C}\right)^{\frac{1}{2}}}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}.$ (10.11)
11. Estimation of $\mathcal{D}_{+}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$
Here, the phase function $-h_{+}$ has no stationary point and the first derivative bound together with the Weyl law suffices.
Proposition 11.1.
We have, for any $\varepsilon>0$
$\check{f}_{m,n,C,+}(\eta)\ll\begin{cases}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta^{\frac{3}{2}}}e^{-\pi|\eta|}&\text{if }\eta>0,\\ X^{-100},&\text{if }\eta\gg\frac{MCX^{\varepsilon}}{NH}.\end{cases}$
Proof.
We have,
$h^{\prime}_{+}(u)=\frac{X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}}+\frac{2\eta}{\pi u}\geq\frac{2\eta}{\pi u}\gg\frac{\eta}{\left(\frac{\sqrt{MN}}{C}\right)^{\frac{1}{2}}}.$ (11.1)
In the case $\eta\gg\frac{MCX^{\varepsilon}}{NH}$, we recall the condition (10.2) and conclude using Lemma 2.8 that the integral $I_{+}(m,n,C,\nu,\eta,r_{1})$ is negligibly small, say $O(X^{-100})$.
Now we consider the more general case $\eta>0$. Applying the first derivative bound, we have,
[…]
The proposition follows once we note that
$\Gamma(1\pm 2i\eta)\gg(|\eta|+1)^{\frac{1}{2}}e^{-\pi|\eta|},$ (11.2)
which follows from the Stirling formula. ∎
11.1. Estimation of $\mathcal{D}_{+}(M,N,C)$
First we prove a preparatory lemma.
Lemma 11.2.
For any integer $q\geq 1$, the following bound holds.
$\sum_{1\leq n\leq N}\frac{|\rho_{j}(nq)|^{2}}{\cosh(\pi\kappa_{j})}\ll q^{\theta+\varepsilon}N(N\kappa_{j})^{\varepsilon}.$ (11.3)
Proof.
by Lemma 2.11, (2.8) and the Hecke relation (2.7), we have
$\sum_{1\leq n\leq N}\frac{|\rho_{j}(nq)|^{2}}{\cosh(\pi\kappa_{j})}\ll\kappa_{j}^{\varepsilon}\sum_{1\leq n\leq N}\left|\sum_{k|(n,q)}\mu(k)\lambda_{j}\left(\frac{q}{k}\right)\lambda_{j}\left(\frac{n}{k}\right)\right|\ll\kappa_{j}^{\varepsilon}q^{\theta+\varepsilon}\sum_{k|q}\sum_{1\leq n\leq\frac{N}{k}}\left|\lambda_{j}\left(n\right)\right|^{2}\ll q^{\theta+\varepsilon}N(N\kappa_{j})^{\varepsilon},$ (11.4)
by Remark 2.10 and Lemma 2.14. ∎
Proposition 11.3.
Let $1\ll K_{1},K_{2}\ll X^{2}$, and $\lambda\geq 0$. For any fixed $c_{1},c_{2}>0$ and any positive integer $q$, we have,
$\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $:=\sum_{\begin{subarray}{c}c_{1}K_{1}\leq\kappa_{j}\leq c_{2}K_{2}\end{subarray}}\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)|\rho_{j}(nq)\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\lambda}}$
$\ll\begin{cases}q^{\theta+\varepsilon}(MN)^{1+\varepsilon}K_{2}^{2-\lambda+\varepsilon}&\text{if}\quad 0\leq\lambda\leq 2\\ q^{\theta+\varepsilon}(MN)^{1+\varepsilon}\text{if}\quad\lambda>2.\end{cases}$
Proof.
We subdivide the $j$-sum into dyadic pieces with $\kappa_{j}\asymp K$ such that $c_{1}K_{1}\leq K\leq c_{2}K_{2}$, and apply the Cauchy-Schwarz inequality to the $m$-sum and the $n$-sum, getting
$\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $\ll\sum_{c_{1}K_{1}\leq K\leq c_{2}K_{2}}\frac{1}{K^{\lambda}}\sum_{\kappa_{j}\asymp K}\frac{\cosh(\pi\kappa_{j})}{e^{\pi\kappa_{j}}}\sqrt{MN}\left(\sum_{n\sim N}\frac{|\rho_{j}(nq)|^{2}}{\cosh(\pi\kappa_{j})}\right)^{\frac{1}{2}}\left(\sum_{m\sim M}\frac{|\rho_{j}(m)|^{2}}{\cosh(\pi\kappa_{j})}\right)^{\frac{1}{2}}$
$\ll q^{\theta+\varepsilon}(MN)^{1+\varepsilon}\sum_{c_{1}K_{1}\leq K\leq c_{2}K_{2}}\frac{1}{K^{\lambda}}\sum_{\kappa_{j}\asymp K}(\kappa_{j})^{\varepsilon},$
by applying (11.3) to sums over $m$ and $n$. Finally, an application of the Weyl law (see Lemma 2.13) applied to the $\kappa_{j}$-sum yields the bound
$\Theta_{c_{1},c_{2}}(M,N,K,\vartheta,\lambda)\ll q^{\theta+\varepsilon}(MN)^{1+\varepsilon}\sum_{c_{1}K_{1}\leq K\leq c_{2}K_{2}}K^{2-\lambda+\varepsilon}.$ (11.5)
The proposition follows. ∎
Now we estimate $\mathcal{D}_{+}(M,N,C)$ (see (10.3)) using the above proposition.
Proposition 11.4.
We have,
$\mathcal{D}_{+}(M,N,C)\ll r_{1}^{\theta}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\left(\frac{MCX^{\varepsilon}}{NH}\right)^{\frac{1}{2}}.$
Proof.
Since we know from Prop. 11.1 that the integral $I_{+}(m,n,C,\nu,\eta,r_{1})$ (see (10.8)) is negligibly small for $\kappa_{j}\gg\frac{MCX^{\varepsilon}}{NH}$, it is enough to estimate $\mathcal{D}_{+}(M,N,C)$ (see (10.3) and (10.7)) for $1\ll\kappa_{j}\ll\frac{MCX^{\varepsilon}}{NH}$. Using (10.3), Prop. 11.1, Prop. 11.3, and (9.5), we easily get
$\mathcal{D}_{+}(M,N,C)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\mathop{\sum_{m\sim M}\sum_{n\sim N}}_{nl\leq m}\sum_{1\ll\kappa_{j}\ll\frac{MCX^{\varepsilon}}{NH}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\frac{3}{2}}}\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}X^{\varepsilon}\left(r_{1}^{\theta}MN\left(\frac{MC}{NH}\right)^{\frac{1}{2}}\right).$
11.2. Estimation of $\mathcal{C}_{+}(M,N,C)$
We start with an easy proposition.
Proposition 11.5.
For $1<\lambda\leq 2$, we have
$\Theta(M,N,K,\lambda)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{0}^{\infty}(nr_{1}m)^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta$
$\ll MNX^{\varepsilon}.$
Proof.
We break the integral into two parts and write
$\Theta(M,N,K,\lambda)$ $\ll\sum_{m\asymp M}\sum_{n\asymp N}(mnr_{1})^{\varepsilon}\left[\int_{0}^{\varepsilon}+\int_{\varepsilon}^{\infty}\right]\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta.$
We know from [Tit86, (2.1.16)] that for $0<\eta<\varepsilon$,
$\eta^{\lambda}\zeta(1+2i\eta)^{2}=\eta^{\lambda}\left(\frac{1}{2i\eta}+\gamma+O\left(\eta\right)\right)^{2}\gg\eta^{\lambda-2}.$
Also, from [Tit86, (3.6.5)], for $\eta\geq\varepsilon$,
$\frac{1}{\left|\zeta(1+2i\eta)\right|}\ll_{\varepsilon}|\log\eta|^{7}.$
The proposition follows since both the integrals are bounded by constants that depend only on $\varepsilon$. ∎
Proposition 11.6.
We have,
$\mathcal{C}_{+}(M,N,C)\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MNX^{\varepsilon}.$
Proof.
This is clear from Prop. 11.1, (10.4) and Prop. 11.5. ∎
12. Estimation of $\mathcal{D}_{-}(M,N,C)$ for $C_{0}\leq C\leq C_{1}$
Our next task is to estimate $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$ (see (10.3), (10.4) and (10.6)). The sum $\mathcal{D}_{-}(M,N,C)$ (see (10.3)) is considerably more difficult to estimate than $\mathcal{D}_{+}(M,N,C)$ because the phase function $-h_{-}(u)$ (see (10.10)) can have a stationary point inside the range of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ (see (10.8)). Differentiating $h_{-}(u)$ (see (10.10)),
$h^{\prime}_{-}(u)=\frac{X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}}-\frac{2\eta}{u\pi},\quad\text{and}\quad h^{\prime\prime}_{-}(u)=\frac{2\eta}{\pi u^{2}}.$ (12.1)
Hence, the stationary point is given by
$u_{0}=u_{0}(m,n,X,\eta):=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}.$ (12.2)
12.1. Further truncation of the $c$-sum
We shall consider two cases, depending on whether or not the factor $w(\frac{4\pi m\sqrt{mn}}{nu^{2}})$ of $A(u)$ (see (10.9)) takes the value $1$. Let us define
$C_{1}:=\max\left\{2^{j}:2^{j}\leq\frac{XN}{100M}\right\}.$ (12.3)
From (3.5) and the fact that $V$ is supported in $[1,4]$, it follows that $w(\frac{4\pi m\sqrt{mn}}{nu^{2}})=1$ if $C_{0}<C\leq C_{1}$. In this range of $C$, the integral we need to analyze becomes
$I_{-}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu+i\eta}\int_{{\mathbb{R}}}A(u)e(-h_{-}(u))\textnormal{d}u,$
where
$A(u)=\frac{u^{4\nu}}{u}V\left(\frac{4\pi\sqrt{mn}}{u^{2}C}\right)g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right).$
When $C$ is in this range we are able to apply Lemma 2.9 since the condition (2.3), which translates to the condition $\eta\gg(mC/nH)X^{\varepsilon}$, is satisfied. The main result of this section is the following proposition.
Proposition 12.1.
For $C_{0}\leq C\leq C_{1}$, we have the bound
$\mathcal{D}_{-}(M,N,C)$ $\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}$
$+r^{\theta}X^{\frac{3}{4}+\varepsilon}N+r_{1}^{\theta}X^{\varepsilon}MN\bigg{(}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}+1\bigg{)}.$
Unlike the previous case of $\mathcal{D}_{+}(M,N,C)$, an application of the Weyl law is not sufficient here. We need to obtain an explicit asymptotic expansion of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ by applying Lemma 2.9. This is essential because if we estimate $I_{-}(m,n,C,\nu,\eta,r_{1})$ then our final bound will not be satisfactory. We will see that it is enough to consider the leading term $\mathcal{L}(m,n,\kappa_{j},\nu)$ of this asymptotic expansion (see (LABEL:lt). The phase $e(-h(u_{0}))$ occurring in $\mathcal{L}(m,n,\kappa_{j},\nu)$ simplifies to $(mn)^{-i\kappa_{j}}$, up to multiplication by a non-oscillatory function that does not involve $m$ and $n$, and we need to take advantage of the cancellations arising out of this oscillatory function. In order to implement this idea we will first have to separate the variables $m$ and $n$ which are at the moment entangled in the function $V$ and they are also constrained by the condition $nl\leq m$. Dropping the condition $nl\leq m$ by positivity as we had done earlier will be too costly here. We remove this condition by approximating the indicator function $\mathbf{1}_{[1,\infty)}(m/nl)$ by the Perron formula (Lemma 2.4). Furthermore, the variables $m$ and $n$ occurring in the argument of the function $V$ are separated by Mellin inversion. Finally, after the two variables are separated, we apply the Cauchy-Schwarz inequality and the the Large Sieve inequality Lemma 2.16 for twisted Fourier coefficients $\rho_{j}(n)n^{-i\kappa_{j}}$.
12.2. Beginning of the proof of Proposition 12.1
By Lemma 2.9, we have the expansion
$I_{-}(m,n,C,\nu,\eta,r_{1})=\frac{2r_{1}^{\nu+i\eta}}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)\sum\limits_{k\leq 3\delta^{-1}A}p_{k}(u_{0})+O_{A,\delta}\left({\eta}^{-A}\right),$ (12.4)
where
$p_{k}(u_{0})=\frac{1}{k!}\left(\frac{i}{4\pi h^{\prime\prime}_{-}(u_{0})}\right)^{k}G^{(2k)}(u_{0}),$ (12.5)
where $A$ and $\delta$ are arbitrary subject to the conditions $A>0$ and $100<\delta<1/$ 10. We define
$G(u)=A(u)e(-H(u)),\quad\text{where}\quad H(u)=h_{-}(u)-h_{-}(u_{0})-\frac{1}{2}h^{\prime\prime}_{-}(u_{0})(u-u_{0})^{2}.$ (12.6)
Now we write $\mathcal{D}_{-}(M,N,C)$ explicitly using (10.3), (10.7) and (12.4) as
$\mathcal{D}_{-}(M,N,C)=\sum\limits_{k\leq 3\delta^{-1}A}\mathcal{D}_{-,k}(M,N,C)+O\left(\mathcal{D}_{-,Z}(M,N,C)\right),$ (12.7)
where
$\mathcal{D}_{-,k}(M,N,C)$ $:=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\frac{1}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}$ (12.8)
×$\times\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\left[\frac{2r_{1}^{\nu+i\eta}}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)p_{k}(u_{0})\right],$
and
$\mathcal{D}_{-,Z}(M,N,C)$ $:=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\left|\frac{\rho_{j}(nr_{1})\rho_{j}(m)\kappa_{j}^{-A}}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}\right|$ $\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}r_{1}^{\nu}}{\nu!\left|(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})\right|}.$ (12.9)
We first consider the case $k=0$ in (12.8). Let us first define
$\mathcal{L}(m,n,\kappa_{j},\nu,C):=2\frac{A(u_{0})}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)$ (12.10)
$=\sqrt{\frac{2\pi}{\kappa_{j}}}e\left(-\frac{2\kappa_{j}}{\sqrt{\pi}}-\frac{1}{8}\right)u_{0}^{4i\eta}{u_{0}}^{4\nu}V\left(\frac{\pi^{2}Xmn}{\kappa_{j}^{2}C}\right)\left(\frac{\pi^{2}X}{\kappa_{j}^{2}}\right)^{3/4}$
$=\sqrt{2}\pi^{2}X^{3/4}\frac{1}{\kappa_{j}^{2}}e\left(-\frac{2\kappa_{j}}{\sqrt{\pi}}-\frac{1}{8}\right)\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{1/2}}\right)^{4i\kappa_{j}}(mn)^{-i\kappa_{j}}\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{1/2}(mn)^{1/4}}\right)^{4\nu}V\left(\frac{\pi^{2}Xmn}{\kappa_{j}^{2}C}\right),$
by (12.2) and (9.3). From the support of $V$, the expression for $u_{0}$ (see (12.2)), and taking into account the condition $nl\leq m$, we see that the effective range of $\kappa_{j}$ is given by
$\frac{\pi}{2}\left(\frac{mnX}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq\pi\left(\frac{mnX}{C}\right)^{\frac{1}{2}};$ (12.11)
in particular, we have
$\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}.$ (12.12)
Even though the presence of the function $V$ restricts the range of $\kappa_{j}$, we, nevertheless, add the condition (12.12) to the sum over $\kappa_{j}$. This condition seems redundant at this stage, but later, when we express the function $V$ as a Mellin inverse of its Mellin transform, then recovering the range of $\kappa_{j}$ becomes easier if we keep this condition. Hence, we write $\mathcal{D}_{-,0}(M,N,C)$ in the following way.
$\mathcal{D}_{-,0}(M,N,C):=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)T(m,n,C),$ (12.13)
where
$T(m,n,C):=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C),$
where
$\mathcal{V}(\kappa_{j},\nu,r_{1})$ $:=\frac{1}{\sinh(2\pi\kappa_{j})}\frac{2^{-2i\kappa_{j}-2\nu}r_{1}^{\nu+i\kappa_{j}}}{\Gamma(1+2i\kappa_{j})\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\ll\frac{e^{-\pi\kappa_{j}}r_{1}^{\nu}}{\nu!\sqrt{\kappa_{j}}},$ (12.14)
which follows from the Stirling asymptotic (see [IK04, Eq. (5.113)]). Let us also define
$\tilde{T}(m,n,C):=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C)|.$ (12.15)
12.3. Removal of the condition $nl\leq m$
Our next step is to capture the condition $nl\leq m$ by Lemma 2.4. So, we write
$\mathbf{1}_{nl\leq m}=\frac{1}{2\pi i}\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\left(\frac{m}{nl}\right)^{s_{1}}\frac{\textnormal{d}s_{1}}{s_{1}}-\frac{1}{2}\mathbf{1}_{nl=m}+O\left(\left(\frac{m}{nl}\right)^{\sigma_{1}}\min\left(1,\frac{1}{T|\log(m/nl)|}\right)+\frac{1}{T}\right),$
where $\sigma_{1}=\Re s_{1}=\varepsilon$ and $T$ is a parameter that will be chosen later. Note that when $m=nl$, the corresponding sum will be weighted by the factor $1/2$ instead of $1$ and this will result in an extra error term. Thus we get,
$\mathcal{D}_{-,0}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}\cdots\frac{1}{2\pi i}\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}T(m,n,C)\left(\frac{m}{nl}\right)^{s_{1}}\frac{\textnormal{d}s_{1}}{s_{1}}+O(E_{1}+E_{2}+E_{3}),$ (12.16)
where
$E_{1}:=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m\neq nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{\sigma_{1}}\frac{1}{|\log(m/nl)|}\tilde{T}(m,n,C),$
$E_{2}:=\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\tilde{T}(m,n,C)\frac{\textnormal{d}|s_{1}|}{|s_{1}|},$
and
$E_{3}:=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\tilde{T}(m,n,C).$
Proposition 12.2.
We have,
$E_{1}$ $\ll r_{1}^{\theta}X^{3/4}\frac{X^{1+\varepsilon}}{TH}MN,$ (12.17)
$E_{2}$ $\ll{r_{1}}^{\theta}X^{3/4+\varepsilon}\sqrt{MN},$
$E_{3}$ $\ll r_{1}^{\theta}\frac{X^{3/4+\varepsilon}}{T}\sqrt{MN}.$
Proof.
First we consider $E_{1}$. Since $m\neq nl$, we have the bound $|\log(m/nl)|\gg 1/m$ by the Taylor series expansion of $\log(1+x)$ for $|x|<1$. Hence
$E_{1}$ $\ll\frac{X^{1+\varepsilon}}{TH}\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\tilde{T}(m,n,C).$
From the support of $V$, $u_{0}\asymp\sqrt{MN}/C$ and hence from (LABEL:lt) and (10.2), we have,
$\tilde{T}(m,n,C)$ $\ll X^{3/4}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi\kappa_{j}}}{\kappa_{j}^{5/2}},$ (12.18)
which yields the bound for $E_{1}$ once we recall Prop. 11.3.
For the sum $E_{2}$, using (12.18) and estimating the integral trivially, we have,
$E_{2}$ $\ll X^{3/4+\varepsilon}\sum_{n}b_{1}\left(\frac{nl}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nl)\rho_{j}(nr_{1})|\frac{e^{-\pi\kappa_{j}}}{\kappa_{j}^{5/2}}.$
Now, by the Cauchy-Schwarz inequality applied to the $n$-sum followed by an application of Lemma 11.2 and Lemma 2.13, we estimate $E_{2}$. The sum $E_{3}$ is estimated in the same manner. ∎
12.4. Analysis of the main term in (12.16)
Now we introduce a dyadic subdivision of the sum over $\kappa_{j}$ and separate the variables $m$ and $n$ in the argument of the function $V$ by Mellin inversion, i.e., by writing
$V(y)=\frac{1}{2\pi i}\int_{(\sigma_{2})}\tilde{V}(s_{2})y^{-s_{2}}\textnormal{d}s_{2}.$ (12.19)
Thus, using (12.16), (LABEL:lt), (10.9), (9.3), (12.2), and (12.19), we get
$\mathcal{D}_{-,0}(M,N,C)\ll X^{\frac{3}{4}}\int\limits_{\sigma-iT}^{\sigma+iT}\int\limits_{(\sigma_{2})}\left(\frac{X}{C}\right)^{-\sigma_{2}}|\tilde{V}(s_{2})|U(M,N,C)\frac{\textnormal{d}|s_{1}|}{|s_{1}|}\textnormal{d}|s_{2}|+E_{1}+E_{2}+E_{3},$ (12.20)
where
$U(M,N,C):=\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\sum_{\kappa_{j}\sim K}\frac{\kappa_{j}^{2\sigma_{2}}}{\kappa_{j}^{2}}\Bigg{|}\sum_{\nu=0}^{\infty}\mathcal{V}(\kappa_{j},\nu,r_{1})\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (12.21)
×$\times\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\right)^{4\nu}\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\Bigg{|}$
$\ll\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\frac{K^{2\sigma_{2}}}{K^{\frac{5}{2}}}\sum_{\nu=0}^{\infty}\left(\frac{2K}{\sqrt{\pi}X^{\frac{1}{2}}}\right)^{4\nu}\frac{r_{1}^{\nu}}{\nu!}$
×$\times\sum_{\kappa_{j}\sim K}e^{-\pi\kappa_{j}}\left|\sum_{m}\sum_{n}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}-\nu}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\right|$
by (12.14) and integration by parts. Observing that $e^{-\pi\kappa_{j}}\asymp 1/\cosh\pi\kappa_{j}$, the above bound yields
$U(M,N,C)$ $\ll\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\frac{K^{2\sigma_{2}}}{K^{\frac{5}{2}}}\sum_{\nu=0}^{\infty}\left(\frac{2K}{\sqrt{\pi}X^{\frac{1}{2}}}\right)^{4\nu}\frac{r_{1}^{\nu}}{\nu!}\sum_{\kappa_{j}\sim K}\frac{1}{\cosh\pi\kappa_{j}}\left|\mathop{\sum_{m}\sum_{n}}\cdots\right|.$ (12.22)
By the Cauchy–Schwarz inequality applied to the $\kappa_{j}$-sum of (12.22), followed by an application of the large sieve inequality (2.16), and taking $\sigma_{2}=\varepsilon$, we obtain
$U(M,N,C)\ll r_{1}^{\theta}X^{\varepsilon}\sqrt{MN}\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\frac{(K^{2}+M^{2})^{\frac{1}{2}}(K^{2}+N^{2})^{\frac{1}{2}}}{K^{\frac{5}{2}}}\sum_{\nu=0}^{\infty}\left(\frac{16K^{4}r_{1}}{\pi^{2}X^{2}MN}\right)^{\nu}.$
Here we have used Hecke multiplicativity (see (2.7)) to separate the factor $r_{1}$ in $\rho_{j}(nr_{1})$ in the manner of the proof of Lemma 11.2. Note that since $\sqrt{MNr_{1}}/C\ll 1$ (see (10.2)), the sum over $\nu$ is absolutely bounded. Now we execute the integrals in (LABEL:cs) over $s_{1}$ and $s_{2}$. Because of the decay of $\tilde{V}$, the integral over $s_{2}$ presents no difficulty and the integral over $s_{1}$ contributes a factor which is $O(\log X)$ if we choose $T=X^{10}$ (see (12.16)). Hence, by Prop. 12.2 with this choice of $T$, we have the bound
$\mathcal{D}_{-,0}(M,N,C)$ $\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}+{r_{1}}^{\theta}X^{\frac{3}{4}+\varepsilon}N,$ (12.23)
since $K\geq\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}$.
Remark 12.3.
If we did not use the large sieve (Lemma 2.16) here, we would have obtain the bound $U(M,N,C)\ll(MN)^{3/4},$ and putting this in (LABEL:cs), Prop. 12.2 shows that the error term can be no better than $O(X^{8/5+\varepsilon})$. The large sieve improves the bound to $O((MN)^{-1/4}(M+N))$ and this gives us a saving which is of size $(X/H)^{3/4}\asymp X^{3/8}$ for our final choice $H=\sqrt{X}$.
12.5. Estimation of $\mathcal{D}_{-,k}(M,N,C)$ for $k>0$
A routine calculation shows that for $k=1,2$, we have
$||p_{k}||_{\infty}\ll\frac{1}{\eta}||p_{0}||_{\infty},\quad\text{and}\quad||p_{0}||_{\infty}\ll||A||_{\infty}\ll\frac{\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}}{\left(\frac{\sqrt{MN}}{C}\right)^{1/2}}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}},$ (12.24)
where we use (10.11), (12.5) and (12.6). Now, (12.24), (12.4) and (2.6) with $j=0$ imply that the terms in the asymptotic expansion of $I_{-}(m,n,C,\nu,\eta,r_{1})$ (see (12.4)) corresponding to $k\geq 1$ are bounded by
$\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta^{3/2}}.$ (12.25)
Combining (12.25), (12.8), (11.2), (10.2), ignoring the condition $nl\leq m$ by positivity and by Prop. 11.3, we have for any $k\geq 1$,
$\mathcal{D}_{-,k}(M,N,C)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left(r_{1}^{\theta+\varepsilon}MN\right).$ (12.26)
12.6. Estimation of $\mathcal{D}_{-,Z}(M,N,C)$
As before, taking $A=2$, and using (9.5) (10.2), (11.3), (11.2) and (12.9),
$\mathcal{D}_{-,Z}(M,N,C)$ $\ll r^{\theta}\left(\frac{X}{H}\right)^{2+\varepsilon}.$ (12.27)
Bringing together (12.7), (12.23), (12.26), (12.27), and taking $\delta=1/20$ and $A=2$, we complete the proof of Prop. 12.1.
13. Estimation of $\mathcal{D}_{-}(M,N,C)$ for $C>C_{1}$
In this case, we cannot apply Lemma 2.9 since the condition (2.3) need not hold here and instead, we apply Lemma 2.7 which is suitable for the weighted stationary phase analysis of integrals over finite intervals. The main result here is:
Proposition 13.1.
For $C_{1}<C\leq\frac{8NX}{M}$, we have
$\mathcal{D}_{-}(M,N,C)\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}+r^{\theta}X^{\frac{3}{4}+\varepsilon}N$
$+r_{1}^{\theta}X^{3/4+\varepsilon}M^{23/32}+r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left(MN\left(\sqrt{M}+\sqrt{N}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}+\sqrt{\frac{X}{H}}\right)+(MN)^{\frac{3}{4}}M\right).$
The integral we need to analyze now is
$I_{-}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu+i\eta}\int_{{\mathbb{R}}}A(u)e(-h_{-}(u))\textnormal{d}u.$ (13.1)
From the support of $w$, the integrand vanishes outside these two points
$\alpha:=\frac{\sqrt{4\pi}\sqrt{m}(mn)^{\frac{1}{4}}}{\sqrt{(X+H)n}}\quad\text{and}\quad\beta:=\frac{\sqrt{4\pi}(mn)^{\frac{1}{4}}}{\sqrt{C_{1}}},$ (13.2)
First we will apply Lemma 2.8 to our integral to observe that when $\eta$ is sufficiently large the integral is negligibly small.
13.1. Restricting the range of $\eta$ up to $\frac{X^{1+\varepsilon}}{H}$
Proposition 13.2.
For any $\varepsilon>0$, if $\eta\gg X^{1+\varepsilon}/{H}$, then we have the bound
$I_{-}(m,n,C,\nu,\eta,r_{1})\ll_{\varepsilon}X^{-100}.$ (13.3)
Proof.
Without loss of generality, we may assume $M\ll X^{1+\frac{\varepsilon}{2}}/H$. Note that since $C_{1}\asymp NX/M$ (see (12.3)), $\beta\asymp(MN)^{1/4}/\sqrt{NX/M}$ and hence
$\beta\sqrt{X}(MN)^{1/4}\asymp M.$ (13.4)
Now, by (12.2),
$\eta>\sqrt{\pi}\beta X^{\frac{1}{2}}(mn)^{\frac{1}{4}}\iff u_{0}>2\beta.$
By our assumption that $\eta\gg\frac{X^{1+\varepsilon}}{H}$, we see that $\eta\gg MX^{\frac{\varepsilon}{2}}$ and hence by (13.4), $\eta>\sqrt{\pi}\beta X^{\frac{1}{2}}(mn)^{\frac{1}{4}}$. In this case, it is easy to verify that
$|h^{\prime}_{-}(u)|>|h^{\prime}_{-}(\beta)|=\frac{2\eta}{\beta\pi}-\frac{X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}}>\frac{\eta}{\beta\pi},$
and hence, by Lemma 2.8, the result follows. ∎
13.2. Decomposition of $\mathcal{D}_{-}(M,N,C)$
Here we may assume that $\eta\ll\frac{X^{1+\varepsilon}}{H}$. We divide the integral into two parts as follows:
$I_{-}(m,n,C,\nu,\eta,r_{1})$ $=2r_{1}^{\nu+i\eta}\left[\int_{\frac{2\sqrt{\pi}\sqrt{m}(mn)^{1/4}}{\sqrt{n(X+H)}}}^{\frac{2\sqrt{\pi}\sqrt{m}(mn)^{1/4}}{\sqrt{Xn}}}+\int_{\frac{\sqrt{4\pi}\sqrt{m}(mn)^{1/4}}{\sqrt{Xn}}}^{\frac{\sqrt{4\pi}(mn)^{1/4}}{\sqrt{C_{1}}}}\right]A(u)e(h_{-}(u))\textnormal{d}u$ (13.5)
$=\sum\limits_{\mu=0,1}I_{-,\mu}(m,n,C,\nu,\eta,r_{1}),$
say. Note that in the second integral, the function $w$ which occurs inside $A(u)$ (see (10.9)) is identically equal to $1$. Correspondingly, we write
$\check{f}_{m,n,C,-}(\eta)=\sum\limits_{\mu=0,1}\check{f}_{m,n,C,-,\mu}(\eta),\quad\text{and}\quad\mathcal{D}_{-}(M,N,C)=\sum\limits_{\mu=0,1}\mathfrak{D}_{-,\mu}(M,N,C),$ (13.6)
where we define (see (10.7))
$\check{f}_{m,n,C,-,\mu}(\eta):=\frac{4}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}I_{-,\mu}(m,n,C,\nu,\eta,r_{1}),$ (13.7)
and by (13.6),
$\mathfrak{D}_{-,\mu}(M,N,C)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,\mu}(\kappa_{j}).$ (13.8)
Now our aim is to estimate the above two sums separately. First we begin with a proposition that will be required on multiple occasions. This is an application of the local (spectral) large sieve (Lemma 2.15).
Proposition 13.3.
For $1\ll\vartheta\leq\frac{K}{2}$, $1\ll K\ll X^{2}$, $\lambda\geq 0$ and $nl\leq m$, we have
$\sum_{\begin{subarray}{c}K+\vartheta\leq\kappa_{j}\leq K+2\vartheta\end{subarray}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\lambda}}\ll r_{1}^{\theta+\varepsilon}\frac{1}{K^{\lambda}}\left(K\vartheta+m\right)\left(mK\right)^{\varepsilon}.$
Proof.
By the Cauchy-Schwarz inequality, the left hand side is
$\ll\frac{1}{K^{\lambda}}\left(\sum_{\begin{subarray}{c}K+\vartheta\leq\kappa_{j}\leq K+2\vartheta\end{subarray}}\frac{|\rho_{j}(nr_{1})|^{2}}{\cosh(\pi\kappa_{j})}\right)^{\frac{1}{2}}\left(\sum_{\begin{subarray}{c}K+\vartheta\leq\kappa_{j}\leq K+2\vartheta\end{subarray}}\frac{|\rho_{j}(m)|^{2}}{\cosh(\pi\kappa_{j})}\right)^{\frac{1}{2}},$
and using (2.8),
$\sum_{\begin{subarray}{c}K+\vartheta\leq\kappa_{j}\leq K+2\vartheta\end{subarray}}\frac{|\rho_{j}(nr_{1})|^{2}}{\cosh(\pi\kappa_{j})}\ll r_{1}^{\theta+\varepsilon}\sum_{k|(n,r_{1})}\sum_{\begin{subarray}{c}K+\vartheta\leq\kappa_{j}\leq K+2\vartheta\end{subarray}}\frac{|\rho_{j}(1)|^{2}}{\cosh(\pi\kappa_{j})}\left|\lambda_{j}\left(\frac{n}{k}\right)\right|^{2}\quad,$
where we use (2.7) and (2.9). Now the proof follows from Lemma 2.15 and the condition $nl\leq m$. ∎
We first estimate $\mathfrak{D}_{-,0}(M,N,C)$.
Proposition 13.4.
For $C_{1}<C\leq\frac{8NX}{M}$, we have
$\mathfrak{D}_{-,0}(M,N,C)\ll r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\sqrt{\frac{X}{H}}.$
Proof.
Using (10.11), (10.2), and (11.2),
$\check{f}_{m,n,C,-,0}(\eta)\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{H}{X}\frac{1}{\sqrt{\eta}}e^{-\pi|\eta|};$
and therefore, using (13.8) and recalling the discussion in §13.1,
[…]
14. Analysis of $I_{-,1}(m,n,C,\nu,\eta,r_{1})$ and estimation of $\mathfrak{D}_{-,1}(M,N,C)$
Proposition 14.1.
For $C_{1}<C\leq\frac{8NX}{M}$, we have,
$\mathfrak{D}_{-,1}(M,N,C)$ $\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}$
$+r_{1}^{\theta}X^{3/4+\varepsilon}M^{23/32}+r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\left(\sqrt{M}+\sqrt{N}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}\right).$
In this range of the variables, we have $w\left(\frac{4\pi m\sqrt{mn}}{nu^{2}}\right)=1$ and the integral in question becomes
$I_{-,1}(m,n,C,\nu,\eta,r_{1}):=2r_{1}^{\nu+i\eta}\int\limits_{\alpha_{1}}^{\beta_{1}}A(u)e(h_{-}(u))\textnormal{d}u,\text{ and }\alpha_{1}:=\frac{\sqrt{4\pi}\sqrt{m}(mn)^{1/4}}{\sqrt{Xn}},\text{ }\beta_{1}:=\frac{\sqrt{4\pi}(mn)^{1/4}}{\sqrt{C_{1}}}$ (14.1)
are the end points. Before proceeding with the calculation, we first note that $\beta_{1}=\beta$ (see (12.3) and (13.2)).
14.1. Decomposition of $\mathfrak{D}_{-,1}(M,N,C)$
Recall that the stationary point is given by $u_{0}=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}$. Here the analysis is somewhat similar to that of $I_{4}$ in §7 but the main differences are that because of the presence of the weight function $A(u)$, we need to use the weighted stationary phase lemma (Lemma 2.7). Also, apart from the large sieve inequality for twisted coefficients (Lemma 2.16), here we also require the local (spectral) large sieve inequality of Jutila (Lemma 2.15). In order to analyze the integral $I_{-,1}(m,n,C,\nu,\eta,r_{1})$ (see (14.1)) and estimate $\mathfrak{D}_{-,1}(M,N,C)$ we consider five intervals in which the stationary point $u_{0}$ can possibly lie, namely, $u_{0}<\alpha_{1}-\delta,u_{0}>\beta_{1}+\delta,|u_{0}-\alpha_{1}|\leq\delta,|u_{0}-\beta_{1}|\leq\delta,\alpha_{1}+\delta<u_{0}<\beta_{1}-\delta$, where we choose
$\delta=\frac{1}{100}\sqrt{\frac{M}{NX}},$ (14.2)
and note that $\beta_{1}>100\delta$ by our choice. We subdivide the sum $\mathfrak{D}_{-,1}(M,N,C)$ accordingly as
$\mathfrak{D}_{-,1}(M,N,C)$ $=\mathfrak{D}_{-,1,1}(M,N,C)+\cdots+\mathfrak{D}_{-,1,5}(M,N,C).$ (14.3)
Furthermore, let us recall that the condition $\kappa_{j}\ll X^{1+\varepsilon}/H$ holds here (see §13.1). Finally we introduce the notation
$c_{m,n}:=\frac{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{2}.$ (14.4)
14.2. Estimation of $\mathfrak{D}_{-,1,1}(M,N,C),\mathfrak{D}_{-,1,2}(M,N,C),\mathfrak{D}_{-,1,3}(M,N,C)$ and $\mathfrak{D}_{-,1,4}(M,N,C)$
Proposition 14.2.
For $C_{1}<C\leq\frac{8NX}{M}$, we have
$(i)\quad\mathfrak{D}_{-,1,1}(M,N,C),\mathfrak{D}_{-,1,2}(M,N,C)\ll r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\left(\sqrt{\frac{X}{H}}+\sqrt{M}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}\right).$
$(ii)\quad\mathfrak{D}_{-,1,3}(M,N,C),\mathfrak{D}_{-,1,4}(M,N,C)\ll r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left((MN)^{\frac{3}{4}}M+MN\right).$
Proof.
Let us consider (i) first and we only estimate $\mathfrak{D}_{-,1,2}(M,N,C)$ below as the other one can be estimated in the same way. The proof is similar to the proof of Prop. 7.3 and we only give a sketch. First note that since $\eta\ll X^{1+\varepsilon}/H$, $u_{0}=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\ll\frac{X^{\frac{1}{2}+\varepsilon}}{H}.$ We write the range $\left[\beta_{1}+\delta,\frac{X^{\frac{1}{2}+\varepsilon}}{H}\right]$ of $u_{0}$ as a union of dyadic subintervals of the form $[\beta_{1}+Y,\beta_{1}+2Y]$, where the dyadic variable $Y$ can take at most $O(\log X)$ values. Let us consider two cases separately: $\delta\leq Y\leq\frac{\beta_{1}}{2}$ and $Y>\frac{\beta_{1}}{2}$.
Case I: $\delta\leq Y\leq\frac{\beta_{1}}{2}$. Suppose $\beta_{1}+Y\leq u_{0}\leq\beta_{1}+2Y$, where $\delta\leq Y\leq\frac{\beta_{1}}{2}$. By (12.1) and the first derivative bound, we have
$I_{-,1}(m,n,C,\nu,\eta,r_{1})$ $\ll r_{1}^{\nu}\|A\|_{\infty}\frac{\beta_{1}}{YX^{\frac{1}{2}}(MN)^{\frac{1}{4}}}\ll\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{YX^{\frac{1}{2}}(MN)^{\frac{1}{4}}}.$
Recall that $C\asymp C_{1}$ here. Using (13.7), (11.2), (10.2), and the last estimate, we have
$\check{f}_{m,n,C,-,1}(\eta)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{YX^{\frac{1}{2}}(MN)^{\frac{1}{4}}}\frac{1}{\sqrt{\eta}}e^{-\pi|\eta|}$ (14.5)
Now we estimate the contribution of case I to the sum $\mathfrak{D}_{-,1,2}(M,N,C)$. We have
$\mathfrak{D}_{-,1,2}(M,N,C,I):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\begin{subarray}{c}Y\textnormal{ dyadic }\\ \delta\leq Y\leq\frac{\beta_{1}}{2}\end{subarray}}\sum_{\begin{subarray}{c}\kappa_{j}\\ \beta_{1}+Y\leq u_{0}\leq\beta_{1}+2Y\end{subarray}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,1}(\kappa_{j})$
$\ll\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\sum_{\begin{subarray}{c}Y\textnormal{ dyadic}\\ \delta\leq Y\leq\frac{\beta_{1}}{2}\end{subarray}}\frac{1}{YX^{\frac{1}{2}}(MN)^{\frac{1}{4}}}\sum_{c_{m,n}(\beta_{1}+Y)\leq\kappa_{j}\leq c_{m,n}(\beta_{1}+2Y)}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\sqrt{\kappa_{j}}}$
by (14.5). Applying Prop. 13.3, we now get
$\mathfrak{D}_{-,1,2}(M,N,C,I)\ll r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\sqrt{M}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}\right),$ (14.6)
by (14.2). Now we come to the other case.
Case II: $Y>\frac{\beta_{1}}{2}$. In this case, the range of $\eta$ we need to consider is (by (14.1) and (12.2))
$\frac{3}{4}\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}\beta_{1}\leq\eta\ll\frac{X^{1+\varepsilon}}{H}.$ (14.7)
By the first derivative bound, (10.11) and (14.1), we have
$I_{-,1}(m,n,C,\nu,\eta,r_{1})\ll\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta},\quad\text{and}\quad\check{f}_{m,n,C,-,1}(\eta)\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta^{\frac{3}{2}}}e^{-\pi|\eta|},$
where we use (11.2), (10.2) and (13.7) to bound $\check{f}_{m,n,C,-,1}(\eta)$. Therefore, the contribution of $\mathfrak{D}_{-,1,2}(M,N,C)$ for $\kappa_{j}$ in this range is given by
$\mathfrak{D}_{-,1,2}(M,N,C,II):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\begin{subarray}{c}Y\textnormal{ dyadic }\\ Y>\frac{\beta_{1}}{2}\end{subarray}}\sum_{\begin{subarray}{c}\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}\\ \beta_{1}+Y\leq u_{0}\leq\beta_{1}+2Y\end{subarray}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,1}(\kappa_{j})$ (14.8)
by (14.7). We we apply Prop. 11.3 to estimate this. We first bring the $Y$-sum and the $\kappa_{j}$-sum outside of the sums over $m$ and $n$. Since $\beta_{1}$ depends on $m,n$, we extend the range of $Y$ and $\kappa_{j}$ by positivity, as follows. We have,
$Y>\frac{\beta_{1}}{2}\implies Y\geq\beta_{0},\quad\text{where}\quad\beta_{0}:=\frac{1}{1000}\frac{\sqrt{M}(MN)^{\frac{1}{4}}}{\sqrt{XN}}.$ (14.9)
Thus, using (14.9) and (14.7), we have
$\mathfrak{D}_{-,1,2}(M,N,C,II)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\sum_{X^{\frac{1}{2}}(MN)^{\frac{1}{4}}\beta_{0}\leq\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}}\sum_{\begin{subarray}{c}Y\text{ dyadic}\\ Y\geq\beta_{0}\end{subarray}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\frac{3}{2}}}$ (14.10)
[…]
The proof is finished in view of (14.6) and (14.10).
Now we consider part (ii). Here, $\eta$ is highly localized and Lemma 2.5 with $k=2$ suffices and we apply Prop. 13.3 to handle the sum over $\kappa_{j}$. We omit the details. ∎
14.3. Estimation of $\mathfrak{D}_{-,1,5}(M,N,C)$ and the weighted stationary phase analysis
Proposition 14.3.
For $C_{1}<C\leq\frac{8NX}{M}$, we have,
$\mathfrak{D}_{-,1,5}(M,N,C)\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}+{r_{1}}^{\theta}X^{\frac{3}{4}+\varepsilon}N$
$+r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}M^{\frac{23}{32}}+r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\left(\sqrt{M}+\sqrt{N}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}\right)+r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left((MN)^{\frac{3}{4}}M+MN\right).$
Here we extract the stationary phase of the integral $I_{-,1}(m,n,C,\nu,\eta,r_{1})$ (see (14.1)) by Lemma 2.7 as follows:
$I_{-,1}(m,n,C,\nu,\eta,r_{1})$ $=2r_{1}^{\nu+i\eta}\int\limits_{\alpha_{1}}^{\beta_{1}}A(u)e(h_{-}(u))\textnormal{d}u=2r_{1}^{\nu+i\eta}\frac{A(u_{0})}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)+O(R_{1}+R_{2}+R_{3}),$
where $\alpha_{1}$ and $\beta_{1}$ are defined in (14.1) and
$R_{1}:=\frac{|A(\beta_{1})|}{|h_{-}^{\prime}(\beta_{1})|}+\frac{|A(\alpha_{1})|}{|h_{-}^{\prime}(\alpha_{1})|},\text{ }R_{2}:=\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha_{1})^{3}}+\frac{1}{(\beta_{1}-x_{0})^{3}}\right),\text{ and }R_{3}:=\frac{QT}{P^{\frac{3}{2}}}\left(1+\frac{Q}{V}\right)^{2},$ (14.11)
where $Q,T,P$ and $V$ are defined in Lemma 2.7. It is clear that the leading term here is $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ (see (LABEL:lt)). Therefore,
$\mathfrak{D}_{-,1,5}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\nu=0}^{\infty}\sum_{\begin{subarray}{c}\kappa_{j}\\ \alpha_{1}+\delta<u_{0}<\beta_{1}-\delta\end{subarray}}\rho_{j}(nr_{1})\rho_{j}(m)$
×$\times\mathcal{V}(\kappa_{j},\nu,r_{1})\Big{[}\mathcal{L}(m,n,\kappa_{j},\nu,C)+O(R_{1}+R_{2}+R_{3})\Big{]}.$
Since we can handle the contribution of the portion of the sum where the stationary point $u_{0}$ varies over a small interval easily, we rewrite the above sum as
$\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\dots\sum_{\nu=0}^{\infty}\Biggl{\{}\sum_{\begin{subarray}{c}\kappa_{j}\\ \alpha_{1}<u_{0}<\beta_{1}\end{subarray}}-\Bigl{(}\sum_{\begin{subarray}{c}\kappa_{j}\\ \alpha_{1}<u_{0}<\alpha_{1}+\delta\end{subarray}}+\sum_{\begin{subarray}{c}\kappa_{j}\\ \beta_{1}-\delta<u_{0}<\beta_{1}\end{subarray}}\Bigr{)}+O\Bigl{(}\sum_{\begin{subarray}{c}\kappa_{j}\\ \alpha_{1}+\delta<u_{0}<\beta_{1}-\delta\end{subarray}}(R_{1}+R_{2}+R_{3})\Bigr{)}\Biggr{\}}$ (14.12)
$=G(M,N,C)-H(M,N,C)+O(J(M,N,C)),$
say.
14.3.1. Contribution of $G(M,N,C)$
Here, the condition $\alpha_{1}<u_{0}<\beta_{1}$ translates to:
$\pi m<\kappa_{j}<\pi\sqrt{\frac{mnX}{C_{1}}},$ (14.13)
by (12.2) and (14.1). This sum is similar to the sum in (12.13) except for the condition (14.13). In (12.13), the corresponding condition on $\kappa_{j}$ was automatic from the support of the function $V$. Here also the presence of the function $V$ in $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ limits $\kappa_{j}$ to lie in the interval given by (12.11) and hence the above upper bound for $\kappa_{j}$ is automatically satisfied. So we have only one extra condition $\kappa_{j}>\pi m$ from which we need to separate the variables $m$ and $\kappa_{j}$. By Lemma 2.4,
$G(M,N,C)$ $=\sum_{\kappa_{j}>\pi m}(\cdots)=\sum_{\kappa_{j}}(\cdots)\frac{1}{2\pi i}\int\limits_{\sigma-iT}^{\sigma+iT}\left(\frac{\kappa_{j}}{\pi m}\right)^{s}\frac{\textnormal{d}s}{s}-\frac{1}{2}\sum_{\kappa_{j}}(\cdots)\mathbf{1}_{\kappa_{j}=\pi m}$ (14.14)
$+O\left(\sum_{\kappa_{j}}|(\cdots)|\left(\left(\frac{\kappa_{j}}{\pi m}\right)^{\sigma}\min\left(1,\frac{1}{T|\log(\frac{\kappa_{j}}{\pi m})|}\right)+\frac{1}{T}\right)\right).$
Here,
$(\cdots)=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\nu=0}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C),$
and recall that the support of $V$ inside $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ forces $\kappa_{j}$ to lie in the interval given by (12.12) which is independent of $m$ and $n$. Now, for estimating the contribution of the first term, we insert the sum over $\kappa_{j}$ inside the integral and estimate the sum as in §12.4. The only difference is the presence of the factors ${\kappa_{j}}^{s}$ and $m^{-s}$ which does not make any difference since we will take $\sigma$ to be a small positive real number $\varepsilon$. Thus the contribution of the integral over $s$ in the sum $G(M,N,C)$ will be the same as the contribution of the leading term $\mathcal{D}_{-,0}(M,N,C)$ in $\mathcal{D}_{-}(M,N,C)$ (see (12.23)) for $C_{0}\leq C\leq C_{1}$, up to a factor of size $O(X^{\varepsilon})$. Also, we choose $T$ to be a large power of $X$ (say, $X^{10}$) as in §12.16 and this only adds a factor of size $\log X$. Thus, we have the bound (from (12.23))
$4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\nu=0}^{\infty}\sum\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C)\frac{1}{2\pi i}\int\limits_{\sigma-iT}^{\sigma+iT}\left(\frac{\kappa_{j}}{\pi m}\right)^{s}\frac{\textnormal{d}s}{s}$ (14.15)
$\ll r_{1}^{\theta}X^{\frac{3}{4}+\varepsilon}\sqrt{MN}\bigg{(}\left(\frac{C}{Xl}\right)^{\frac{1}{4}}\frac{1}{\sqrt{N}}+\left(\frac{C}{Xl}\right)^{\frac{3}{4}}\frac{(M+N)}{N^{\frac{3}{2}}}+\left(\frac{C}{Xl}\right)^{\frac{5}{4}}\frac{(MN)}{N^{\frac{5}{2}}}\bigg{)}+r^{\theta}X^{\frac{3}{4}+\varepsilon}N.$
By (LABEL:lt), (12.14) and (10.2), the contribution of the term $\frac{1}{2}\mathbf{1}_{\kappa_{j}=\pi m}$ is bounded by
$X^{3/4}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\begin{subarray}{c}\kappa_{j}\\ \kappa_{j}=\pi m\end{subarray}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi\kappa_{j}}}{\kappa_{j}^{5/2}}$ (14.16)
[…]
Now we estimate the contribution of the error terms in (14.14). The term $O(1/T)$ presents no problem, and as for the other error term, we subdivide the sum over $\kappa_{j}$ into two parts: $|\kappa_{j}-\pi m|>1$ and $|\kappa_{j}-\pi m|\leq 1$. In the first case, we observe that
$1/|\log(\kappa_{j}/\pi m)|\ll m,$
with an absolute implied constant. Since $T$ is very large, the contribution of this part is negligibly small. For the other case, we take the minimum in the $O$-term to be $1$ and use the Weyl law to bound the resulting sum and the calculation is identical to that of the case $\kappa_{j}=\pi m$ treated above. Thus the contribution of the $O$-terms is also $O\left(r_{1}^{\theta}X^{3/4+\varepsilon}M^{23/32}\right)$.
14.3.2. Contribution of $H(M,N,C)$
We first note that using (LABEL:lt) and (10.11),
$\mathcal{L}(m,n,\kappa_{j},\nu,C)\ll\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\sqrt{\kappa_{j}}},$
since from the support of $V$, $u_{0}\asymp\frac{(MN)^{1/4}}{\sqrt{C}}$. Therefore, using the same analysis as in the proof of part (ii) of Prop. 14.2 and using (12.14), we conclude that the contribution of these two terms is
$O\left(r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left((MN)^{\frac{3}{4}}M+MN\right)\right).$ (14.17)
14.3.3. Contribution of $J(M,N,C)$
We consider $R_{1}$ first and only estimate the first term since the second term can be handled in a similar way. By making a dyadic subdivision of the range $\left(\alpha_{1}+\delta,\beta_{1}-\delta\right)$ of $u_{0}$ into $O(\log X)$ many subintervals and proceeding as before, and using (10.11) and (10.9), we see that the contribution of the term $\frac{|A(\beta_{1})|}{|h_{-}^{\prime}(\beta_{1})|}$ in $\mathfrak{D}_{-,1,5}(M,N,C)$ is bounded by
$\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots\sum_{\begin{subarray}{c}Y\text{ dyadic}\\ \delta\leq Y\leq\frac{\beta_{1}}{2}\end{subarray}}\sum_{c_{m,n}(\beta_{1}-2Y)<\kappa_{j}<c_{m,n}(\beta_{1}-Y)}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{1}{YX^{\frac{1}{2}}(MN)^{\frac{1}{4}}}\frac{1}{\sqrt{\kappa_{j}}}e^{-\pi|\kappa_{j}|}$ (14.18)
[…]
For $R_{2}$ and $R_{3}$, we first note that
$Q=V=\beta_{1},P=\kappa_{j}\text{ and }T=\frac{\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}}{\beta_{0}}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}},$
where $\beta_{0}$ and $\beta_{1}$ is defined in (14.9) and (14.1) respectively. The contribution of $R_{2}$ is handled in the same way as above and we make use of the facts that $u_{0}\asymp\beta_{1}\asymp\alpha_{1}$ and $nl\leq m$ and the contribution of $R_{2}$ in $\mathfrak{D}_{-,1,5}(M,N,C)$ is
$O\left(r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\left(\sqrt{N}+1\right)\right).$ (14.19)
It is easy to see that the contribution of $R_{3}$ in $\mathfrak{D}_{-,1,5}(M,N,C)$ is
$O\left(r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MN\right).$ (14.20)
Thus we complete the proof of Prop. 14.3 by clubbing together (LABEL:illish), (LABEL:G-leading), (LABEL:kappa=m), (LABEL:error1), (14.19), (14.20), (14.17) and recalling (9.5). Combining (14.3), Prop. 14.2, and Prop. 14.3, Prop. 14.1 follows.
15. Esitimation of $\mathcal{C}_{-}(M,N,C)$ for $C\geq C_{0}$
Proposition 15.1.
For $C\geq C_{0}$, we have
$\mathcal{C}_{-}(M,N,C)\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MNX^{\varepsilon}.$
Proof.
We consider two cases.
Case I: $C_{0}\leq C\leq C_{1}$. From (12.12), we recall that
$\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\eta\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}.$
Also, from (12.1), (12.2), (12.4), (12.24) and (12.25), and taking $\delta=1/20$ and $A=2$,
$I_{-}(m,n,C,\nu,\eta,r_{1})\ll\frac{r_{1}^{\nu}}{\sqrt{h_{-}^{\prime\prime}(u_{0})}}\sum_{0\leq k\leq 3\delta^{-1}A}|p_{k}(u_{0})|+\frac{1}{\eta^{A}}\ll\frac{\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}}{\sqrt{\eta}}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}.$ (15.1)
Using (10.7), (11.2), (15.1) and (10.2),
$\check{f}_{m,n,C,-}(\eta)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{e^{-\pi|\eta|}}{\eta}\sum_{\nu=0}^{\infty}\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}\frac{1}{\nu!}\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{e^{-\pi|\eta|}}{\eta}.$
Thus, proceeding as in the proof of Prop. 11.5 and by (10.4), we have,
$\mathcal{C}_{-}(M,N,C)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MNX^{\varepsilon}.$
Case II: $C_{1}<C\leq\frac{8NX}{M}$. Here, we have
$\mathcal{C}_{-}(M,N,C)\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MNX^{\varepsilon},$
and the proof is similar to that Proposition 11.6 and estimates from §13.1 are used. ∎
16. Proof of Proposition 3.4: final steps
16.1. Proof of Proposition 9.1
Bringing together (10.6) and the estimates for $\mathcal{D}_{+}(M,N,C)$, $\mathcal{C}_{+}(M,N,C)$, $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$, i.e., Prop. 11.4, Prop. 11.6, Prop. 12.1, Prop. 13.1 and Prop. 15.1, we have proved Prop. 9.1.
16.2. A sketch of the proof of Proposition 9.2
Note that here the role of the function $f_{m,n,C}(t)$ (see (10.1)) is now played by the function
$f_{m,n,C}^{\star}(t)=w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\frac{u}{X}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g^{\star}\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(-\frac{u^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right).$
Note that the phase function now is essentially the same as before with the only change being that the parameter $X$ has now been replaced by $u$, which is of the same size as $X$. Because the function $g$ is now replaced by $g^{\star}$ (see (9.8)), instead of the function $A(u)$ we now have a corresponding amplitude function $A^{\star}(u)$ which satifies the bound
$\|A^{\star}\|_{\infty}\ll\frac{\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}}{\left(\frac{\sqrt{MN}}{C}\right)^{\frac{1}{2}}}\left(\frac{C}{N^{2}l}\right)^{\frac{1}{4}}.$
The proofs of the analogues of Prop. 11.4, Prop. 11.6, Prop. 12.1, Prop. 13.1, and Prop. 15.1 are along the same lines but with some minor changes. For example, the bound for $\mathcal{D}_{-}(M,N,C)$ in Prop. 12.1 (for $C\leq C_{1}$) is now replaced by $\left(r^{\theta}\left(X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)\right)$. To see this, note that much of the proof of the proposition carries over but now $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ (see (LABEL:lt)) is changed to $\mathcal{L}^{\star}(m,n,\kappa_{j},\nu,C)$ where the factor $\left(\frac{{\pi}^{2}X}{{\eta}^{2}}\right)^{\frac{3}{4}}$ is now replaced by $\left(\frac{{\pi}^{2}X}{{\eta}^{2}}\right)^{\frac{1}{4}}$. Note that the main contribution to the bound in Prop. 12.1 comes from the term $E_{2}$ (see Prop. 12.2), $\mathcal{D}_{-,0}(M,N,C)$ (see (12.23)), and $\mathcal{D}_{-,k}(M,N,C)$ (see (12.26)). The term analogous to $E_{2}$ now contributes $O\left(r^{\theta}X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ and the same contribution also comes from the analogue of $\mathcal{D}_{-,0}(M,N,C)$. In both the cases, the exponent of $X$ changes from $\frac{3}{4}$ to $\frac{1}{4}$ and the exponent of $\kappa_{j}$ changes from $\frac{5}{2}$ to $\frac{3}{2}$. The analogue of $\mathcal{D}_{-,k}(M,N,C)$ now contributes $\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$ since the factor $\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}$ is now changed to $\left(\frac{C}{N^{2}l}\right)^{\frac{1}{4}}$. We obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(X^{{\frac{1}{4}}}\left(X/H\right)^{2}+X^{\frac{1}{4}}\left(X/H\right)^{\frac{3}{2}}\right)\right)$ for the analogue of $\mathcal{D}_{-}(M,N,C)$ in Prop. 13.1 (for $C>C_{1}$) as well. For the analogue of $\mathcal{D}_{+}(M,N,C)$ in Prop. 11.4, we now have the bound $O\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$. Together, the analogous of $\mathcal{C}_{-}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$ in Prop.s 15.1 and 11.6 now contribute $O\left(X^{{\frac{1}{4}}+\varepsilon}{\left(X/H\right)}^{\frac{3}{2}}\right)$.
16.3. Proof of Proposition 3.4
Now, from Prop. 9.1, (9.1), and (9.2), we obtain the bound $O\left(r^{\theta}\frac{X^{2+\varepsilon}}{H}\right)$ for $R_{1,w}(X)$, and as discussed earlier, the bound for ${R^{\prime}}_{1,w}(X)$ is the same. Hence we obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(\frac{X^{2}}{H}+X\sqrt{H}\right)\right)$ for $T_{1,w}(X)$ (see (8.2)) and the same bound holds for $T_{2,w}(X)$. We have, $R_{3,w}(X)\ll r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X^{\frac{1}{4}}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ by Prop. 9.2 and (8.4), and the same bound holds for ${R^{\prime}}_{3,w}(X)$. Hence, by (8.3), $T_{3,w}(X)\ll r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X^{\frac{1}{4}}\left(X/H\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}+X\sqrt{H}\right)$ and the same bound holds for $T_{4,w}(X)$. Therefore, Prop. 3.4 follows from the above discussion and Prop. 5.6.
Remark 16.1.
We explain here why we need to choose $r=O\left(X^{1/3}\right)$. Note that from (10.2), we have the condition $r_{1}\ll{C_{0}}^{2}/MN$. For $R_{3}(M,N,C,X,u)$, we similarly have the condition $r_{1}\ll{C_{0}^{\star}}^{2}/MN$. Since we need to choose $C_{0}^{\star}\leq\frac{X^{2/3}+\varepsilon}{l}$ (see (9.6)), we get $r\ll X^{1/3}$.

### v2

@@
 $E(r)=O(r^{131/208+\varepsilon}),\ \text{for any }\varepsilon>0,$
-due to Huxley [Hux96]  [^1] Recently, Xiaochun Li and Xuerui Yang have posted an article (see https://arxiv.org/abs/2308.14859) that claims to improve this bound of Huxley., while the conjectured bound $O(r^{1/2+\varepsilon})$ remains hopelessly out of reach of the current technology.
+due to Huxley [Hux96]  [^1] Recently, Xiaochun Li and Xuerui Yang have announced (see https://arxiv.org/abs/2308.14859) an improvement over the bound of Huxley.
+and the conjectured bound $O(r^{1/2+\varepsilon})$ remains hopelessly out of reach of the current technology.
 If we replace the circle by a hyperbola $xy=N$, with $N\longrightarrow\infty$, the problem of counting lattice points with positive coordinates under the hyperbola is known as Dirichlet’s divisor problem and the situation here is completely analogous to the Gauss circle problem (see [IK04, Chap.
 1]).
-An interesting lattice point counting problem concerns counting elements $\gamma\in SL(2,{\mathbb{Z}})$; i.e., integer matrices $\gamma=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ with determinant $ad-bc=1$, such that $||\gamma||\leq x$, where we take the norm to be the standard Euclidean norm on ${{\mathbb{R}}}^{4}$ after identifying the matrix $\gamma$ with the point $(a,b,c,d)$; i.e.,
-$||\gamma||^{2}=a^{2}+b^{2}+c^{2}+d^{2}.$
+An interesting lattice point counting problem concerns counting elements $\gamma\in SL(2,{\mathbb{Z}})$; i.e., integer matrices $\gamma=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ with determinant adbc=1ad-bc=1, such that $\gamma||\leq x$, where we take the norm to be the standard Euclidean norm on ${{\mathbb{R}}}^{4}$ after identifying the matrix $\gamma$ with the point $(a,b,c,d)$; i.e.,
+$\gamma||^{2}=a^{2}+b^{2}+c^{2}+d^{2}.$
 This problem can be interpreted as the analogue of the Gauss circle problem in the hyperbolic plane $\mathbb{H}$.
@@
 For example, the Gauss circle problem is nothing but the problem of improving the bound on the error term in the Weyl law for the flat torus ${{\mathbb{Z}}}^{2}\backslash{{\mathbb{R}}}^{2}$.
-The above hyperbolic lattice point counting problem is related to the spectrum of the hyperbolic Laplacian acting on $\ΓL^{2}(\Gamma\backslash\mathbb{H})$ (as a densely defined operator) and after initial advances made by Delsarte [Del42] and Huber, who solved the problem for cocompact groups (see [Hub56, Hub61]), the original problem for $\Gamma=SL(2,{\mathbb{Z}})$ could be tackled once Selberg established the complete spectral resolution of the Laplacian for groups with parabolic elements.
+The above hyperbolic lattice point counting problem is related to the spectrum of the hyperbolic Laplacian acting on $\)L^{2}(\Gamma\backslash\mathbb{H}$) (as a densely defined operator) and after initial advances made by Delsarte [Del42] and Huber, who solved the problem for cocompact groups (see [Hub56, Hub61]), the original problem for $\Gamma=SL(2,{\mathbb{Z}})$ could be tackled once Selberg established the complete spectral resolution of the Laplacian for groups with parabolic elements.
 Selberg’s result (see, e.g., [IK04, Corollary 15.12]) says
-$\sum_{\begin{subarray}{c}\gamma\in SL(2,{\mathbb{Z}})\\ ||\gamma||\leq x\end{subarray}}1=6x^{2}+O(x^{4/3}).$ (1.1)
-For an account of this beautiful and deep circle of ideas, see, e.g., [IK04, Chap. 15], [Iwa02], [LP82], [Pat75] and the original articles by Selberg, viz. [Sel56], [Sel63].
-This result has been vastly generalized by Duke, Rudnick and Sarnak [DRS93] to varieties $V$ with an algebraic group action for which $V(\mathbb{R})$ is affine and symmetric by applying advanced results from harmonic analysis.
+[…]
+For an account of this beautiful and deep circle of ideas, see, e.g., [IK04, Chap.
+15], [Iwa02], [LP82], [Pat76] and the original articles by Selberg, viz.
+[Sel56], [Sel63].
+For an interesting twisted version of Selberg’s result, see [Goo83].
+Selberg’s work has been vastly generalized by Duke, Rudnick and Sarnak [DRS93] to varieties $V$ with an algebraic group action for which $V(\mathbb{R})$ is affine and symmetric by applying advanced results from harmonic analysis.
 A special case of their result says
-$\sum_{\begin{subarray}{c}\gamma\in M(n,{\mathbb{Z}})\\ det(\gamma)=r\\ \\ ||\gamma||\leq x\end{subarray}}1=c_{n,r}x^{n^{2}-n}+O_{\varepsilon}(x^{n(n-1)-1/(n+1)+\varepsilon}),$ (1.2)
 […]
+where $M(n,{\mathbb{Z}})$ is the set of $n\times n$ integer matrices, $\|\cdot\|$ denotes the Euclidean norm $\sqrt{tr(\gamma{\gamma}^{t})}$, $r\neq 0$ is a fixed integer, (i.e., $r=O(1)$), $c_{n,r}>0$ is an explicit constant and $\varepsilon>0$ is arbitrary.
+The saving $\delta_{n}=1/(n+1)$ in the exponent of the error term in (1.2) was later improved by Gorodnik, Nevo and Yehoshua [GNY17] to, roughly, $2\delta_{n}$ and very recently it has been improved further by Blomer and Lutsko [BL24] who have shown that the saving is at least $1+1/\sqrt{2}+O(1/n)$ (for $r=1$ and $n\geq 3$).
+In the case $n=2$, however, the bound $O(X^{4/3})$ for the error term due to Selberg remains unbeaten to this day.
 […]
@@
 where $a,b,c,d$ vary over the integers and $r$ is a fixed non-zero integer. We have,
-$\mathpzc{S}_{r}(X)=\frac{16}{\zeta(2)}\frac{\sigma(|r|)}{|r|}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}),$ (1.3)
-as long as $r=O\left(X^{{1/3}}\right)$.
-Here $\theta$ is any admissible exponent for the Generalized Ramanujan Conjecture, $\varepsilon$ is an arbitrarily small positive real number, and $\sigma(n)$ denotes the sum of the divisors of a positive integer $n$.
-By the work of Kim and Sarnak [Kim03, Appendix 2] it is known that $\theta\leq\frac{7}{64}$.
+$\mathcal{S}_{r}(X)=\frac{16}{\zeta(2)}\frac{\sigma(|r|)}{|r|}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}),$ (1.3)
+as long as $r=O\left(X^{{1/3}}\right)$. Here $\theta$ is an admissible exponent towards the Ramanujan Conjecture.
+Conjecturally $\theta$ can be taken to be zero and by the work of Kim and Sarnak [Kim03, Appendix 2] it is known that $\theta\leq\frac{7}{64}$.
+Here $\varepsilon$ is any positive real number that can be arbitrarily small and $\sigma(n)$ denotes the sum of all divisors of a positive integer $n$.
+Remark 1.1.
+The question we address here can be put in the general context of counting integral points on surfaces of the type $Q(x1,x_{2},\cdots,x_{n})=r$ where $Q$ is a quadratic form.
+See [HB96] for an introduction to this topic.
+However, counting with smooth weights seems more natural from that point of view and we have considered the “smooth” analogue of this problem in our companion article [GG25].
+The interest in our question from this angle has been discussed in some detail in the introduction to [GG25].
+As one should expect, we get a much stronger error term in this situation.
+Indeed, if $V$ is a fixed smooth function with compact support, say, in $[1,2]$ and if $r$ is a fixed non-zero integer, then we show
+$\mathop{\sum_{a}\sum_{b}\sum_{c}\sum_{d}}_{\begin{subarray}{c}ad-bc=r\end{subarray}}V\left(\frac{a}{X}\right)V\left(\frac{b}{X}\right)V\left(\frac{c}{X}\right)V\left(\frac{d}{X}\right)=M_{V}(X,r)+O_{\varepsilon}\left(|r|^{\theta}X^{1+\varepsilon}\right),$
+where
+$M_{V}(X,r)=\sum_{l|r_{1}}\sum_{k>0}\frac{\mu(k)}{k}\iiint\frac{1}{z}V\left(\frac{x}{X}\right)V\left(\frac{lky}{\ X}\right)V\left(\frac{lkz}{X}\right)V\left(\frac{r_{1}+lkxy}{zlkX}\right)\textnormal{d}x\textnormal{d}y\textnormal{d}z,$
+and for $0\neq r\in(-3X^{2},X^{2})$, we may write $M_{V}(X,r)=c_{V}\frac{\sigma(|r|)}{|r|}X^{2}+O(\sigma(|r|)$, where $c_{V}>0$ is a constant depending only on $V$.
+The proof of this result, though non-trivial, is considerably simpler than the proof of Theorem 1.3. See §1.2 below or [GG25] for further details.
 Remark 1.2.
 The reason we need to take $r=O(X^{1/3})$ is explained in Remark 15.1.
-However, with more work, it seems possible to extend the range of $r$.
-Also, at the cost of a worse exponent of $X$, we can make the dependence of $r$ on the error term disappear.
-See Remark 7.6 for a precise statement.
-As mentioned earlier, we can put our problem in the general context of counting integral points on affine hypersurfaces defined over ${\mathbb{Z}}$ and more specifically to counting vectors $\mathbf{a}\in{{\mathbb{Z}}}^{n}$ to equations of the form
-$Q(\mathbf{x})=r$ (1.4)
-in an expanding box $[-X,X]^{n^{2}}$, where $Q$ is a quadratic form in $n$ variables over ${\mathbb{Z}}$ and $r$ is a fixed integer.
-This is a classical problem with a rich history and the circle method, developed and refined by Ramanujan, Hardy, Littlewood, Kloosterman and many others, has been employed here with great success.
-See the article [HB96] by Heath-Brown for an introduction with modern perspectives to this problem.
-Usually one counts integral solutions to (1.4) by attaching weights $w_{i}(a_{i}/X)$, where $w_{i}$’s are smooth functions with compact support and it is possible to get fairly strong bound on the error term relative to the main term in many cases (see, e.g., [HB96, Theorems 4-8]).
-If we count with sharp cuts, however, then the situation is very different and in general we can only get a saving of an arbitrary power of $\log X$ (see [Nie10]).
-It is interesting, therefore, to address this problem for forms of special shapes in a small number of variables and the determinant form $Q(x,y,z,w)=xy-zw$ is particularly suitable because of the structural advantage it affords.
-In fact, the only two results we could find in the literature that deal with unweighted counting are about determinant forms in two or more variables.
+This restriction can possibly be relaxed with more work but we have not tried to do so in this paper.
+Also, at the cost of a worse exponent of $X$, we can enlarge the admissible range of $r$.
+In particular, we have
+${\mathcal{S}_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(\max\left(X^{5/3},|r|\right)X^{\varepsilon}\right).$
+See Remark 3.2 for a proof.
+Muhammad Afifurrahman [Afi24] has also independently obtained the same result recently by a totally different method.
+We employ an elaborate stationary phase analysis and the Weil bound for Kloosterman sums whereas Afifurrahman uses clever counting arguments and earlier results on counting integer points on modular hyperbolas.
+Remark 1.3.
+What about the case $r=0?$ The best known result in this direction is due to Ayyad, Cochrane, and Zhang [ACZ96] who have shown that
+$\sum_{\begin{subarray}{c}x_{1}x_{2}-x_{3}x_{4}=0\\ 1\leq x_{1},x_{2},x_{3},x_{4}\leq X\end{subarray}}1=\frac{12}{{\pi}^{2}}X^{2}\log X+O\left(X^{19/{13}}(\log X)^{\frac{7}{13}}\right)$
+(see [ACZ96, Thm. 3]). The main ingredient in the proof is the clever observation that
+$\sum_{\begin{subarray}{c}x_{1}x_{2}-x_{3}x_{4}=0\\ 1\leq x_{i}\leq X\end{subarray}}1=\sum_{\begin{subarray}{c}(a,b)=1\\ 1\leq a,b\leq X\end{subarray}}\left(\sum_{x_{1}/x_{3}=a/b}1\right)^{2}.$
+It is an interesting and challenging problem to improve on this error term.
+The only other results we could find in the literature that deal with unweighted counting for indefnite forms are [GN10] and [Roe05] and both consider determinant forms.
 Gorodnik and Nevo (see [GN10, Cor.
@@
 In comparison, we save $1/2$ in the exponent of the error term in Theorem 1.1.
-If we count with smooth weights then we can save even more and can also take $r$ to be much larger as the following result shows.
-Theorem 1.3.
-Let
-$S_{V}(X,r)=\mathop{\sum_{a}\sum_{b}\sum_{c}\sum_{d}}_{\begin{subarray}{c}ad-bc=r\end{subarray}}V\left(\frac{a}{X}\right)V\left(\frac{b}{X}\right)V\left(\frac{c}{X}\right)V\left(\frac{d}{X}\right),$
-where $V$ is a fixed smooth function with compact support in $[1,2]$. Then for $0<r\ll X^{2}$, we have
-$S_{V}(X,r)=M_{V}(X,r)+O_{\varepsilon}\left(r^{1/4}X^{1+\varepsilon}\right),$
-where
-$M_{V}(X,r)=\sum_{l|r}\sum_{k>0}\frac{\mu(k)}{k}\iiint\frac{1}{z}V\left(\frac{x}{X}\right)V\left(\frac{lky}{X}\right)V\left(\frac{lkz}{X}\right)V\left(\frac{r+lkxy}{zlkX}\right)\textnormal{d}x\textnormal{d}y\textnormal{d}z.$
-It is not hard to see that $M_{V}(X,r)\asymp X^{2}$ if $r$ is not very large; e.g., if $0<|r|\leq 2X^{2}$. Moreover, by the mean value theorem of Calculus, we see that
-$S_{V}(X,r)=KX^{2}+O\left(\sigma(r)\right)+O\left(r^{1/4}X^{1+\varepsilon}\right),$
-where
-$K=\frac{1}{\zeta(2)}\frac{\sigma(r)}{r}\iiint\frac{1}{t}V\left(u\right)V\left(v\right)V\left(t\right)V\left(\frac{uv}{t}\right)\textnormal{d}u\,\textnormal{d}v\,\textnormal{d}t.$
-Although we use automorphic methods to prove this result, a result of comparable strength can also be obtained by employing the method of [HB96].
-In fact, if we take $r$ to be of size $X^{2}$ then Theorem 1.3 gives an error term of the same strength as the one in [HB96, Theorem 4].
-However, it is possible to improve the exponent of $r$ to $\theta$ with more work.
-It is of interest to consider more general quadratic forms and to investigate whether some of the results of [HB96] (see also [Get18]) can be refined in special cases by introducing the spectral methods of automorphic forms into the analysis after the application of the $\delta$-symbol method.
-In this context, see also the discussion after [HB96, Cor.
-1].
-We have included the above result to illustrate the impact of counting with smooth weights and have given only a brief sketch of the proof in §1.2 since this proof is much easier compared to the proof of Theorem 1.1.
-An interesting variation of the general counting problem is to attach arithmetic weights such the von Mangold function or coefficients of cusp forms to the variables of summation (see [Kum20]) and this is the topic of an upcoming work by the second author along with P.
-Sharma and S.K.
-Singh.
-1.1. Extensions and applications of the results
-1.1.1. Determinant equations in Analytic Number Theory
-Although we are mainly motivated by the intrinsic interest in the problem, counting solutions to the determinant equations $ad-bc=r$, $r\neq 0$ with certain weights $\alpha(a),\alpha(b)$ etc.
-come up naturally in the subject, for example, in the study of amplified moments of $L$-functions (see, e.g., [DFI97], [DFI02, §16], [BC18]).
+1.1. Some applications and further remarks
+Although we are mainly motivated by the intrinsic interest in the problem, the problem of counting solutions to the determinant equations adbc=rad-bc=r, $r\neq 0$ with certain weights $\alpha(a),\alpha(b)$ etc.
+come up naturally, for example, in the study of amplified moments of $L$-functions (see, e.g., [DFI97], [DFI02, §16], [BC18]).
 In [BC18] and [DFI97], sums of the form
@@
 Here, $\mathcal{N}_{i}=[N_{i}/2,N_{i}]$, $\alpha,\beta$ are arbitrary sequences, and $f,g$ are smooth functions satisfying suitable decay conditions on their derivatives.
-It is possible to allow arbitrary weights in some of the variables at the cost of an worse error term.
+It is possible to allow arbitrary weights in some of the variables at the cost of a worse error term.
 For example, the following is a consequence of an intermediate step in the proof of Theorem 1.1.
@@
 This result is a direct consequence of (3.2), (3.9), (3.11) and Remark 3.2.
-Unlike (1.5), here only one variable is weighted by an arbitrary sequence.
-If one chooses weights of arithmetic nature then further possibilities for improvements may open up.
-For example, if we take $\alpha(a)=\mathbf{1}_{a\equiv 0(\text{mod }q)}$, the indicator function of the integral multiples of $q$, where $q>1$ is a fixed integer and keep the other weights the constant function $1$ (or smooth weights $W(a/X),W(b/X)$ etc.), then also our method applies and this leads to the study of the spectral theory for the congruence group $\Gamma_{0}(q)$.
-1.1.2. Extensions to other quadratic forms in four variables
-The method of our proof does not distinguish between the determinant equation $xy-zw=r$ and the equation $pxy-qzw=r$ where $p,q$ are fixed integers such that $(p,q)|r$.
-One can easily obtain analogous results for the latter equation of the same strength as long as $p,q=O(X^{\varepsilon})$.
-In fact, by a simple change of variables, we can derive an asymptotic formula for the number of integer solutions to an equation of the form $f_{1}(x,y)-f_{2}(z,w)=r$, where $f_{1}$ and $f_{2}$ both factor over ${\mathbb{Z}}$ as products of two linear forms and the coefficients of both forms are assumed to be $O(X^{\varepsilon})$.
-However, the formula we obtain in this way counts the number of solutions in a new domain that depends on the coefficients of $f_{1}$ and $f_{2}$.
-To clarify the situation, let us give an illustrative example.
-Let us take $f_{1}(x,y)=f_{2}(x,y)=x^{2}-y^{2}$.
-Then we obtain the surface $x^{2}-y^{2}-z^{2}+w^{2}=r$.
-Now, by the change of variables $(x,y,z,w)\rightarrow(x^{\prime},y^{\prime},z^{\prime},w^{\prime})$, where $x+y=x^{\prime},x-y=y^{\prime},z+w=z^{\prime},z-w=z^{\prime}$, we obtain the surface $x^{\prime}y^{\prime}-z^{\prime}w^{\prime}=r$.
-By Theorem 1.1, we have a counting formula for the number of integer solutions to $x^{\prime}y^{\prime}-z^{\prime}w^{\prime}=r$, where $(x^{\prime},y^{\prime},z^{\prime},w^{\prime})\in[-X,X]^{4}$.
-Now the inverse change of variables transform this box to a four-dimensional tetrahedron $T(X)$ which can be described as the convex hull of the points
-$(X,0,0,0),(-X,0,0,0),(0,X,0,0),(0,-X,0,0),\cdots,(0,0,0,X),(0,0,0,-X).$
-Thus we obtain the following asymptotic formula:
-Corollary 1.5.
-We have,
-$\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}a^{2}-b^{2}-c^{2}+d^{2}=r\\ a,b,c,d\in T(X)\end{subarray}}1=\frac{16}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}).$
-1.1.3. Extensions to quadratic forms in three variables
-If we consider quadratic forms of the shape $Q(x,y,z)=px^{2}-qyz$, then also our methods can be adopted to study integral solutions to $Q(x,y,z)=r$.
-Here, one encounters a sum of Salié sums instead of Kloosterman sums and Proskurin’s formula generalizing the Kuznetsov formula (see [Pro03]) is applicable here.
-1.1.4. Relation to the shifted convolution sum problem for the restricted divisor function
-The shifted convolution sum refers to the problem of the estimation or obtaining asymptotic formula for sums of the type $\sum_{n\leq x}\lambda(n)\lambda(n+h)$ or the corresponding smoothed sum, where $\lambda$ is an interesting arithmetical function, e.g., the divisor function, coefficients of a cusp form etc., and the shift $h$ is a fixed non-zero integer.
+Our question is also related to the problem of the estimation or obtaining asymptotic formula for sums of the type $\sum_{n\leq x}\lambda(n)\lambda(n+h)$ or the corresponding smoothed sum, where $\lambda$ is an interesting arithmetical function, e.g., the divisor function, coefficients of a cusp form etc., and the shift $h$ is a fixed non-zero integer.
+Such sums are known in the literature as shifted convolution sums and they play a fundamental role in modern Analytic Number Theory, especially in studying moments of $L$-functions.
+See [Ing27], [Est32], [DI82], [Mot94], [Meu01], [DFI94], [Blo04], [HM06], to name a few.
 Lattice point counting problems can sometimes be interpreted as the shifted convolution problem for coefficients of appropriate automorphic forms.
@@
 $\sum_{n\leq X}{\tau}^{\ast}_{\sqrt{X}}(n){\tau}^{\ast}_{\sqrt{X}}(n+r)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X+O({|r|}^{\theta}X^{3/4+\varepsilon}).$ (1.5)
-Compared to the classical additive divisor problem, estimating the error term for this sum is harder and the reason is clear.
-Writing $n=bc$ and $n+r=ad$, here all the variables are individually restricted by sharp edges, whereas in the sum $\sum_{n\leq X}\tau(n)\tau(n+h)$, the four variables are subject to only one such restriction ($bc\leq X$) and it is possible to treat two of the three independent variables as “smooth” variables.
-1.2. Method of the proofs of Theorem 1.1 and Theorem 1.3
-The theorems above are proved by the so called “Kloostermania” techniques, i.e., we reduce the problems to that of estimation of sums of Kloosterman sums which is then carried out by employing the spectral theory of automorphic forms.
-Let us first consider the proof of Theorem 1.3 which serves as a warm-up for the proof of Theorem 1.1.
+Estimating the error term for this sum is harder than in the case of the classical additive divisor problem.
+Writing $n=bc$ and $n+r=ad$, here all the four variables $a,b,c,d$ are individually restricted by sharp edges, whereas in the sum $\sum_{n\leq X}\tau(n)\tau(n+h)$, the four variables are subject to only one such restriction ($bc\leq X$) and it is possible to treat two of the three remaining variables as varying smoothly.
+1.2. Discussions on the proof
+Unlike in the case of counting integer points in an Euclidean ball.
+the choice of the norm here does not allow us to launch Harmonic Analysis directly into the counting problem.
+So we “forget” the group structure of $SL(2,{\mathbb{Z}})$ temporarily and treat the equation $det(\gamma)=r$ as the equation defining a surface: xyzw=rxy-zw=r.
+Of course, the group structure remains present throughout, only hidden, and after several transformations, it resurfaces as we are led to studying the automomorphic spectrum of $SL(2,{\mathbb{Z}})$ as explained below.
+Let us first see what happens if we count with smooth weights, i.e., the case of (1.1).
 The first step is to eliminate one of the four variables, say $d$, and to interpret the determinant equation as a congruence modulo one of the other variables, say, $a$.
-This is a natural and well-known idea.
-Then we apply Poisson summation (see Lemma 2.1) to one of the remaining two variables, say $b$.
-As usual, the 00-th frequency in the dual sum gives the main term and after one more application of Poisson summation in the variable $c$ to the remaining sum yields a sum of Ramanujan sums (coming from the $zero$-th term in new dual variable) that can be easily estimated, plus the following double sum:
+Then we apply Poisson summation to one of the remaining two variables, say $b$.
+As usual, the zero-th frequency in the dual sum gives the main term and after one more application of Poisson summation to the remaining sum we obtain a sum of Ramanujan sums (coming from the 0-th term in new dual variable) that can be easily estimated and the following double sum:
 $\sum_{\begin{subarray}{c}n\in{\mathbb{Z}}\\ n\neq 0\end{subarray}}\sum_{\begin{subarray}{c}m\in{\mathbb{Z}}\\ m\neq 0\end{subarray}}\iint V\left(\frac{x}{X}\right)V\left(\frac{y}{X}\right)\sum_{a}\frac{1}{a^{2}}V\left(\frac{a}{X}\right)V\left(\frac{r+yx}{aX}\right)e\left(\frac{-nx}{a}\right)e\left(\frac{-my}{a}\right)S(nr,-m,a)\textnormal{d}x\,\textnormal{d}y.$
-At this stage, the Weil bound already gives an asymptotic formula but we get a better error term by applying spectral methods of automorphic forms via the Kuznetsov formula.
-Since the variable $a$ is weighted by a “flat” function $V(x/X)$ (i.e., $y^{j}V^{(j)}(y)\ll 1$), it is not difficult to obtain strong enough bounds for the Bessel transforms after application of the Kuznetsov formula so as to estimate the spectral side efficiently.
-For example, a direct application of [IK04, Theorem 16.9]) yields Theorem 1.3.
-The extra saving one gets by averaging over the variable $a$ is reminiscent of what Heath-Brown calls “double Kloosterman refinement” (see [HB96, §1]).
-In contrast to the proof of Theorem 1.3, several new features emerge when we employ the above strategy towards proving Theorem 1.1.
-After the second application of Poisson summation, if we apply Kuznetsov formula directly then estimating the Bessel transforms that come up becomes quite difficult since we no longer have the advantage of a flat weight function.
-Instead we study the Fourier transform $\widehat{G}(n/a_{1},m/a_{1})$ (see (5.1)) carefully and extract stationary phases of the type $e\left(-2\sqrt{\frac{mnX}{a_{1}}}\right)$ (see Prop.
-3.3).
-However, a straight-forward application of the stationary phase lemmas to the entire interval does not give good error terms and therefore, we need to carry out a delicate analysis involving subdivision of both the range of integration and the range of the stationary point which depends on the three variables of summation $a_{1}$, $m,$ and $n$ (see §5.2 to §7).
-When the stationary point is sufficiently localized, the variables are also under significant restrictions and exploiting these restrictions is crucial in estimating the contributions of the error terms arising out of the stationary phase analysis.
-After we apply the Kuznetsov formula, the phase of the $K$-Bessel function combines with the already existing phase and in certain ranges of the variables we are able to extract a stationary phase that “looks like” $(mn)^{i\kappa_{j}}$ (see(LABEL:lt)).
-Now, a large sieve inequality (see Lemma 2.16) for the twisted coefficients $\rho_{j}(n)n^{i\kappa_{j}}$, that arose in a work of Deshouillers and Iwaniec [DI86] in connection with the Phillips-Sarnak theory on destruction of cusp forms for generic groups (see [PS85]), gives us an extra saving that is crucial.
-In complimentary ranges, the Weyl law and the local (spectral) large sieve of Jutila (Lemma 2.11) are essential.
+At this stage, the Weil bound already gives an asymptotic formula but we get a better error term by applying the Kuznetsov formula.
+Since the variable $a$ is weighted by a “flat” function $V(x/X)$ (i.e., $y^{j}V^{(j)}(y)\ll 1$), it is not very difficult to obtain strong bounds on the Bessel transforms.
+See [GG25] for details.
+In contrast to the smooth case, several new features emerge when we employ the above strategy to the sharp-cuts sum considered in Theorem 1.1.
+In this situation, by fairly elementary means, it is not difficult to obtain an asymptotic formula.
+In fact, an application of the Weil bound allows us to to get an error term of size $O(X^{7/4+\varepsilon})$ rather easily; see (5.7).
+A more elaborate and technically challenging treatment but still within the realm of classical techniques yields the bound $O(X^{5/3+\varepsilon})$; see (3.2).
+However, our main goal is to obtain a stronger error term and to this end we employ the so called “Kloostermania” techniques (see [Li10]), i.e., we use the Kuznetsov formula (see Lemma 2.2) to transform the problem of estimating a sum of Kloosterman sums to that of estimating a sum of Fourier coefficients of Maass forms ranging over the full $GL(2)$ spectrum.
+We begin in a way similar to the smooth case discussed above, after approximating the indicator function of the interval $[1,X]$ by a smooth function, but if we apply the Kuznetsov formula directly after the second application of Poisson summation, then it seems extremely difficult to obtain good estimates for the resulting Bessel transforms since we no longer have the a “flat” weight function.
+To overcome this problem our idea is to analyze the Fourier transform $\widehat{G}(n/a_{1},m/a_{1})$ (see (5.2)) carefully and extract stationary phases of the type $e\left(-2\sqrt{\frac{mnX}{a_{1}}}\right)$ (see Prop.
+3.2).
+However, a straight-forward application of the stationary phase lemma applied to the entire interval produces error terms that are large.
+Instead, we imagine the stationary point, which depends on three variables of summation $a_{1}$, $m,$ and $n$ (see §5.1 to §7), to be moving in an interval and we break up the range of the stationary point into smaller subintervals and we also subdivide the range of integration suitably so that the stationary point is highly localized.
+See §7.2for details.
+When the stationary point is sufficiently localized, the variables are under severe restrictions and exploiting these restrictions leads to strong estimates for the contributions of the error terms arising from stationary phase analysis.
+After we apply the Kuznetsov formula, the phase of the $K$-Bessel function combines with the already existing phase and in certain ranges of the variables we are able to extract a stationary phase that “looks like” $(mn)^{i\kappa_{j}}$ (see(LABEL:lt)).At this stage, a large sieve inequality (see Lemma 2.6) for the twisted coefficients $\rho_{j}(n)n^{i\kappa_{j}}$ that arose in a work of Deshouillers and Iwaniec [HST86] in connection with the Phillips-Sarnak theory on destruction of cusp forms for generic groups (see [PS85]) gives us an extra saving in certain ranges of the variables.
+In the complimentary ranges, the Weyl law and the local (spectral) large sieve of Jutila (Lemma 2.1) suffice.
 In the end we are able to exploit cancellations in each of the three independent variables $a,b$ and $c$.
-Apart from the results on automorphic forms and the Weil bound for Kloosterman sums, the stationary phase method in its different formulations plays a crucial role in the proof.
+The extra saving one gets by exploiting the average over the variable $a$ can be compared to what Heath-Brown calls “double Kloosterman refinement” in HB), though the method of averaging over the moduli in [HB96] is quite different.
+In the proof, we have been brief about some standard details.
+The interested reader can find full details in the PhD thesis of the second author available at https://sites.google.com/view/rachita-guria/research?authuser=0.
 1.3. Notations and conventions
 The letter $\varepsilon$ denotes a positive real number which can be taken to be as small as required and in different occurrences it may assume different values.
-The notation “$f(y)=O(g(y))$” or “$f(y)\ll g(y)$”, where $g$ is a positive function, means that there is a constant $c>0$ such that $|f(y)|\leq cg(y)$ for any $y$ in the concerned domain.
+The notation “f(y)=O(g(y))f(y)=O(g(y))” or $f(y)\ll g(y)$”, where $g$ is a positive function, means that there is a constant $c>0$ such that $|f(y)|\leq cg(y)$ for any $y$ in the concerned domain.
 The dependence of this implied constant on some parameter(s) may sometimes be displayed by a suffix (or suffixes) and may sometimes be suppressed.
 Also, $f(y)\asymp g(y)$ denotes $f(y)\ll g(y)$ as well as $g(y)\ll f(y)$.
-The symbol :=assign:=:= signifies that the left hand side is being defined.
+The symbol :=:= signifies that the left hand side is being defined.
 For a positive integer $n$, $\tau(n),\sigma(n),\phi(n)$ denotes, respectively the total number of positive divisors of $n$, the sum of the divisors of $n$, and the number of residue classes modulo $n$ that are prime to $n$.
@@
 Acknowledgements
-The authors thank Tim Browning, Gergely Harcos, Vinay Kumaraswamy, Gopal Maiti, Ritabrata Munshi, D.
-Surya Ramana, Prahlad Sharma and Igor Shparlinski for helpful comments.
+The authors thank Valentin Blomer, Tim Browning, Étienne Fouvry, Gergely Harcos, Vinay Kumaraswamy, Gopal Maiti, Ritabrata Munshi, D.
+Surya Ramana, Peter Sarnak, Prahlad Sharma and Igor Shparlinski for helpful comments.
 This work is part of the second author’s PhD thesis.
@@
 2. Preliminaries
-In this section we recall some standard results from the literature which will be required in the proofs of our results formula and some applications of the formula.
-Lemma 2.1.
-(Application of Poisson summation)
-[…]
-$\sum_{\begin{subarray}{c}n\equiv a(\textnormal{mod }q)\\ n\in{\mathbb{Z}}\end{subarray}}f\left(n\right)=\frac{1}{q}\sum_{n\in{\mathbb{Z}}}e\left(\frac{an}{q}\right)\hat{f}\left(\frac{n}{q}\right).$
-Proof.
-See, e.g. ,[IK04, Theorem 4.4 and Exercise 4]. ∎
-Lemma 2.2.
-Under the same assumptions on $a,q$ and $f$ as in Lemma 2.1, we have,
-$\sum_{(n,q)=1}g(n)e\left(\frac{a\bar{n}}{q}\right)=\frac{1}{q}\sum_{n}\hat{g}(n/q)S(n,a,q),$
-where
-$S(m,n,q)=\sideset{}{{}^{\star}}{\sum}_{\beta(\textnormal{mod }q)}e\left(\frac{m\beta+n\bar{\beta}}{q}\right)$
-is the Kloosterman sum (see [IK04]) and $\bar{\beta}$ denotes the multiplicative inverse of $\beta$ modulo $q$ (if $q=1$, $\bar{\beta}=S(m,a;q)=1$ by convention).
-Proof.
-This follows from subdividing the sum over $n$ into residue classes coprime to $q$ and then applying Lemma 2.1. ∎
-Lemma 2.3.
-(The Weil bound for Kloosterman sums)
-Let $a,b,c$ be integers, $c$ positive. We have
-$\left|S(a,b;c)\right|\leq\tau(c)(a,b,c)^{\frac{1}{2}}c^{\frac{1}{2}}.$
-Proof.
-See [IK04, Corollary (11.12)]. ∎
-Lemma 2.4.
-(The Perron formula)
-$\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{x^{s}}{s}\textnormal{d}s=h(x)+O\left(x^{c}\min\left(1,\frac{1}{T|\log x|}\right)\right),$
-where
-$h(x)=\begin{cases}1,&\text{ if }x>1\\ \frac{1}{2},&\text{ if }x=1\\ 0,&\text{ if }x<1,\end{cases}$
-for any $x>0$, $x\neq 1$, and any $T,c>0$, with an absolute implied constant. If $x=1$, the same equality holds with the error term replaced by $O(c/T)$.
-Proof.
-See [Dav80, Chap. 17]. ∎
-2.1. Results on exponential integrals
-Lemma 2.5.
-(The $k$-th derivative bound) Suppose $\psi\in C^{1}[a,b]$. Let $\phi$ be a real valued and smooth in $(a,b)$ such that $\left|\phi^{(k)}\right|\geq\lambda_{k}$ for all $x\in(a,b)$. Then
-$\left|\int\limits_{a}^{b}e^{i\phi(x)}\psi(x)\textnormal{d}x\right|\leq c_{k}\lambda_{k}^{-1/k}\left[\|\psi\|_{\infty}+\int\limits_{a}^{b}\left|\psi^{\prime}(x)\right|\textnormal{d}x\right]$
-holds when $k\geq 2$, or $k=1$ and $\phi^{\prime}(x)$ is monotonic. Here $c_{k}>0$ is a constant that depends only on $k$.
-Proof.
-See [Ste93, Eq. (6) & Prop. 2, Chap. 8]. ∎
-Lemma 2.6.
-(The stationary phase lemma) Let $g(x)$ be real-valued function on $[a,b]$ with $g^{\prime}(x_{0})=0$ for some $x_{0}\in[a,b]$. Suppose
-$0<\lambda_{2}\leq g^{\prime\prime}(x)$ (2.1)
-holds throughout the interval, and in addition $|g^{(3)}(x)|\leq\lambda_{3}$ and that $|g^{(4)}(x)|\leq\lambda_{4}$ for all $x\in[a,b]$. Then
-$\int\limits_{a}^{b}e(g(x))\textnormal{d}x=\frac{e(g(x_{0})+1/8)}{\sqrt{g^{\prime\prime}(x_{0})}}+O(R_{1})+O(R_{2}),$
-where
-$R_{1}=\min\left(\frac{1}{\lambda_{2}(x_{0}-a)},\frac{1}{\sqrt{\lambda_{2}}}\right)+\min\left(\frac{1}{\lambda_{2}(b-x_{0})},\frac{1}{\sqrt{\lambda_{2}}}\right)\textnormal{ and }R_{2}=\frac{(b-a)\lambda_{4}}{\lambda_{2}^{2}}+\frac{(b-a)\lambda_{3}^{2}}{\lambda_{2}^{3}}.$
-If instead of (2.1) we have $g^{\prime\prime}(x)\leq-\lambda_{2}<0$ throughout the interval, then the main term becomes
-$\frac{e(g(x_{0})-1/8)}{\sqrt{|g^{\prime\prime}(x_{0})|}}.$
-Proof.
-See [Mon94, Theorem 9]. ∎
-Lemma 2.7.
-(The weighted stationary phase lemma) Let $f(x)$ be a real function, four times continuously differentiable for $\alpha\leq x\leq\beta$, and let $g(x)$ be a real function, three times continuously differentiable for $\alpha\leq x\leq\beta$.
-Suppose that there are positive parameters $T,P,V,Q,$ with
-$Q\geq\beta-\alpha,\quad V\geq\frac{Q}{\sqrt{P}},$
-and positive constants $C_{r}$ such that, for $\alpha\leq x\leq\beta$,
-$\left|f^{(r)}(x)\right|\leq\frac{C_{r}P}{Q^{r}},\quad\left|g^{(s)}(x)\right|\leq\frac{C_{s}T}{V^{s}}\quad\textnormal{and}\quad f^{\prime\prime}(x)\geq\frac{P}{C_{2}Q^{2}},$
-for $r=2,3,4$, and $s=0,12,3$.
-Suppose also that $f^{\prime}(x)$ changes sign from negative to positive at a point $x=x_{0}$ with $\alpha<x_{0}<\beta$.
-If $P$ is sufficiently large in terms of $C_{2},C_{3}$, and $C_{4}$, then we have
-$\int\limits_{\alpha}^{\beta}g(x)e(f(x))\textnormal{d}x=\frac{g(x_{0})e\left(f(x_{0})+\frac{1}{8}\right)}{\sqrt{f^{\prime\prime}(x_{0})}}+\frac{g(\beta)e(f(\beta))}{2\pi if^{\prime}(\beta)}-\frac{g(\alpha)e(f(\alpha))}{2\pi if^{\prime}(\alpha)}$
-$+O\left(\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha)^{3}}+\frac{1}{(\beta-x_{0})^{3}}\right)\right)+O\left(\frac{QT}{P^{3/2}}\left(1+\frac{Q}{V}\right)^{2}\right).$
-Proof.
-See [Hux96, Lemma 5.5.6]. ∎
-Lemma 2.8.
-(The stationary phase lemma with smooth weights) Let $Y\geq 1,T,Q,U,R>0$ and suppose $w$ is a smooth function with support in $[\alpha,\beta]$ satisfying
-$w^{(j)}(t)\ll_{j}TU^{-j}.$
-Suppose that $h$ is a smooth function on $[\alpha,\beta]$ such that
-$\left|h^{\prime}(t)\right|\geq R,\quad\textnormal{ and }\quad h^{(j)}(t)\ll_{j}YQ^{-j},\textnormal{ for }j=2,3\dots,$
-for some $R>0$. Then the integral $I$ defined by
-[…]
-satisfies
-$I\ll_{A}(\beta-\alpha)T\left[\left(\frac{QR}{\sqrt{Y}}\right)^{-A}+(RU)^{-A}\right].$
-Proof.
-See [BKY13, Lemma 8.1]. ∎
-Lemma 2.9.
-(Asymptotic expansion of exponential integrals with smooth weights) Let $0<\delta<\frac{1}{10},T,Y,V,V_{1},Q>0,$ $Z=Q+X+Y+V_{1}+1$ and assume that
-$Y\geq Z^{3\delta},\quad\quad V_{1}\geq V\geq\frac{QZ^{\frac{\delta}{2}}}{\sqrt{Y}}.$ (2.3)
-Suppose $w$ is a smooth function on ${\mathbb{R}}$ with support on an interval $J$ of length $V_{1}$ satisfying
-$w^{(j)}\ll_{j}TV^{-j},$
-for all $j\in{\mathbb{N}}_{0}$, and $h$ is a smooth function on $J$ such that there exists a unique point $t_{0}\in J$ such that $h^{\prime}(t_{0})=0$, and furthermore that
-$h^{\prime\prime}(t)\gg YQ^{-2},\quad h^{(j)}(t)\ll_{j}YQ^{-j},\text{ for }j=1,2,3,\dots,t\in J.$
-Suppose also that $h^{\prime}(t)$ changes sign from negative to positive at a point $t=t_{0}$. Then the integral $I$ defined in (2.2) has an asymptotic expansion of the form
-$I=\frac{e^{ih(t_{0})}}{\sqrt{h^{\prime\prime}(t_{0})}}\sum\limits_{n\leq 3\delta^{-1}A}p_{n}(t_{0})+O_{A,\delta}\left(Z^{-A}\right),$ (2.4)
-where
-$p_{n}(t_{0})=\frac{\sqrt{2\pi}e^{\frac{\pi i}{4}}}{n!}\left(\frac{i}{2h^{\prime\prime}(t_{0})}\right)^{n}G^{(2n)}(t_{0}),\text{ }G(t)=w(t)e^{iH(t)},\textnormal{ and }H(t)=h(t)-h(t_{0})-\frac{h^{\prime\prime}(t_{0})}{2}(t-t_{0})^{2},$ (2.5)
-for some arbitrary $A>0$. Furthermore, each $p_{n}$ is a rational function in $h^{\prime\prime},h^{\prime\prime\prime},\dots,$ satisfying
-$\frac{d^{j}}{dt_{0}^{j}}p_{n}(t_{0})\ll_{j,n}T\left(V^{-j}+Q^{-j}\right)\left(\left(\frac{V^{2}Y}{Q^{2}}\right)^{-n}+Y^{-\frac{n}{3}}\right).$ (2.6)
-Proof.
-See [BKY13, Prop. 8.2]. ∎
+We recall some standard results that will be required.
 2.1. Results from the theory of automorphic forms
-We refer to [DI82] for the definitions and properties of Maass cusp forms and related topics.
-Let us fix an orthonormal basis of Maass cusp forms $\{u_{j}(z)\}_{1}^{\infty}$ for the full modular group $\text{SL}(2,{\mathbb{Z}})$, consisting of common eigenfunctions of all the Hecke operators $T_{n},n\geq 1$.
+Let us fix an orthonormal basis of Maass cusp forms $\{u_{j}(z)\}_{1}^{\infty}$ for the full modular group $\text{SL}(2,{\mathbb{Z}})$, consisting of common eigenfunctions of all the Hecke operators $T_{n},n\geq 1$ (see [DI82] for definitions).
 Thus every $u_{j}(z)$ is an eigenfunction of the Laplace operator with the associated eigenvalue $s_{j}(1-s_{j})$, $s_{j}\in{\mathbb{C}}$.
@@
 Proof.
-See [DI86, Theorem 6]. ∎
+See [HST86, Theorem 6]. ∎
+2.2. Results from Analysis
+Lemma 2.7.
+(The Perron formula)
+[…]
+for any $x>0$, $x\neq 1$, and any $T,c>0$, with an absolute implied constant. If $x=1$, the same equality holds with the error term replaced by $O(c/T)$.
+Proof.
+See [Dav80, Chap. 17]. ∎
+Lemma 2.8.
+(The $k$-th derivative bound) Suppose $\psi\in C^{1}[a,b]$. Let $\phi$ be a real valued and smooth in $(a,b)$ such that $\left|\phi^{(k)}\right|\geq\lambda_{k}$ for all $x\in(a,b)$. Then
+$\left|\int\limits_{a}^{b}e^{i\phi(x)}\psi(x)\textnormal{d}x\right|\leq c_{k}\lambda_{k}^{-1/k}\left[\|\psi\|_{\infty}+\int\limits_{a}^{b}\left|\psi^{\prime}(x)\right|\textnormal{d}x\right]$
+holds when $k\geq 2$, or $k=1$ and $\phi^{\prime}(x)$ is monotonic. Here $c_{k}>0$ is a constant that depends only on $k$.
+Proof.
+See [Ste93, Eq. (6) & Prop. 2, Chap. 8]. ∎
+Lemma 2.9.
+(The stationary phase lemma) Let $g(x)$ be real-valued function on $[a,b]$ with $g^{\prime}(x_{0})=0$ for some $x_{0}\in[a,b]$. Suppose
+$0<\lambda_{2}\leq g^{\prime\prime}(x)$ (2.4)
+holds throughout the interval, and in addition $|g^{(3)}(x)|\leq\lambda_{3}$ and that $|g^{(4)}(x)|\leq\lambda_{4}$ for all $x\in[a,b]$. Then
+$\int\limits_{a}^{b}e(g(x))\textnormal{d}x=\frac{e(g(x_{0})+1/8)}{\sqrt{g^{\prime\prime}(x_{0})}}+O(R_{1})+O(R_{2}),$
+where
+$R_{1}=\min\left(\frac{1}{\lambda_{2}(x_{0}-a)},\frac{1}{\sqrt{\lambda_{2}}}\right)+\min\left(\frac{1}{\lambda_{2}(b-x_{0})},\frac{1}{\sqrt{\lambda_{2}}}\right)\textnormal{ and }R_{2}=\frac{(b-a)\lambda_{4}}{\lambda_{2}^{2}}+\frac{(b-a)\lambda_{3}^{2}}{\lambda_{2}^{3}}.$
+If instead of (2.4) we have $g^{\prime\prime}(x)\leq-\lambda_{2}<0$ throughout the interval, then the main term becomes
+$\frac{e(g(x_{0})-1/8)}{\sqrt{|g^{\prime\prime}(x_{0})|}}.$
+Proof.
+See [Mon94, Theorem 9]. ∎
+Lemma 2.10.
+(The weighted stationary phase lemma) Let $f(x)$ be a real function, four times continuously differentiable for $\alpha\leq x\leq\beta$, and let $g(x)$ be a real function, three times continuously differentiable for $\alpha\leq x\leq\beta$.
+Suppose that there are positive parameters $T,P,V,Q,$ with
+$Q\geq\beta-\alpha,\quad V\geq\frac{Q}{\sqrt{P}},$
+and positive constants $C_{r}$ such that, for $\alpha\leq x\leq\beta$,
+$\left|f^{(r)}(x)\right|\leq\frac{C_{r}P}{Q^{r}},\quad\left|g^{(s)}(x)\right|\leq\frac{C_{s}T}{V^{s}}\quad\textnormal{and}\quad f^{\prime\prime}(x)\geq\frac{P}{C_{2}Q^{2}},$
+for $r=2,3,4$, and $s=0,12,3$.
+Suppose also that $f^{\prime}(x)$ changes sign from negative to positive at a point $x=x_{0}$ with $\alpha<x_{0}<\beta$.
+If $P$ is sufficiently large in terms of $C_{2},C_{3}$, and $C_{4}$, then we have
+$\int\limits_{\alpha}^{\beta}g(x)e(f(x))\textnormal{d}x=\frac{g(x_{0})e\left(f(x_{0})+\frac{1}{8}\right)}{\sqrt{f^{\prime\prime}(x_{0})}}+\frac{g(\beta)e(f(\beta))}{2\pi if^{\prime}(\beta)}-\frac{g(\alpha)e(f(\alpha))}{2\pi if^{\prime}(\alpha)}$
+$+O\left(\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha)^{3}}+\frac{1}{(\beta-x_{0})^{3}}\right)\right)+O\left(\frac{QT}{P^{3/2}}\left(1+\frac{Q}{V}\right)^{2}\right).$
+Proof.
+See [Hux96, Lemma 5.5.6]. ∎
+Lemma 2.11.
+(The stationary phase lemma with smooth weights) Let $Y\geq 1,T,Q,U,R>0$ and suppose $w$ is a smooth function with support in $\alpha,\beta]$ satisfying
+$w^{(j)}(t)\ll_{j}TU^{-j}.$
+Suppose that $h$ is a smooth function on $\alpha,\beta]$ such that
+$\left|h^{\prime}(t)\right|\geq R,\quad\textnormal{ and }\quad h^{(j)}(t)\ll_{j}YQ^{-j},\textnormal{ for }j=2,3\dots,$
+for some $R>0$. Then the integral $I$ defined by
+$I=\int\limits_{{\mathbb{R}}}w(t)e^{ih(t)}dt\quad\textnormal{ satisfies }\quad I\ll_{A}(\beta-\alpha)T\left[\left(\frac{QR}{\sqrt{Y}}\right)^{-A}+(RU)^{-A}\right].$ (2.5)
+Proof.
+See [BKY13, Lemma 8.1]. ∎
+Lemma 2.12.
+(Asymptotic expansion of exponential integrals with smooth weights) Let $0<\delta<\frac{1}{10},T,Y,V,V_{1},Q>0,$ $Z=Q+X+Y+V_{1}+1$ and assume that
+$Y\geq Z^{3\delta},\quad\quad V_{1}\geq V\geq\frac{QZ^{\frac{\delta}{2}}}{\sqrt{Y}}.$ (2.6)
+Suppose $w$ is a smooth function on $\mathbb{R}$} with support on an interval $J$ of length $V_{1}$ satisfying $w^{(j)}\ll_{j}TV^{-j}$, for all $j\in{\mathbb{N}}_{0}$, and $h$ is a smooth function on $J$ such that there exists a unique point $t_{0}\in J$ such that $h^{\prime}(t_{0})=0$, and furthermore that $h^{\prime\prime}(t)\gg YQ^{-2},h^{(j)}(t)\ll_{j}YQ^{-j},\text{ for }j=1,2,3,\dots,t\in J.$ Suppose also that $h^{\prime}(t)$ changes sign from negative to positive at a point $t=t_{0}$.
+Then the integral $I$ defined in (2.5) has an asymptotic expansion of the form
+$I=\frac{e^{ih(t_{0})}}{\sqrt{h^{\prime\prime}(t_{0})}}\sum\limits_{n\leq 3\delta^{-1}A}p_{n}(t_{0})+O_{A,\delta}\left(Z^{-A}\right),$ (2.7)
+where
+$p_{n}(t_{0})=\frac{\sqrt{2\pi}e^{\frac{\pi i}{4}}}{n!}\left(\frac{i}{2h^{\prime\prime}(t_{0})}\right)^{n}G^{(2n)}(t_{0}),\text{ }G(t)=w(t)e^{iH(t)},\textnormal{ and }H(t)=h(t)-h(t_{0})-\frac{h^{\prime\prime}(t_{0})}{2}(t-t_{0})^{2},$ (2.8)
+for some arbitrary $A>0$. Furthermore, each $p_{n}$ is a rational function in $h^{\prime\prime},h^{\prime\prime\prime},\dots,$ satisfying
+$\frac{d^{j}}{dt_{0}^{j}}p_{n}(t_{0})\ll_{j,n}T\left(V^{-j}+Q^{-j}\right)\left(\left(\frac{V^{2}Y}{Q^{2}}\right)^{-n}+Y^{-\frac{n}{3}}\right).$ (2.9)
+Proof.
+See [BKY13, Prop. 8.2]. ∎
 3. Initial reduction and the main steps of the proof
@@
 $S_{r}(X)=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c,d\leq X\\ ad-bc=r\end{subarray}}1=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O(r^{\theta}X^{3/2+\varepsilon}),$ (3.1)
-for $0<|r|\ll X^{1/3}$, since $\mathpzc{S}_{r}(X)=8S_{r}(X)$ by symmetry.
-We first eliminate $d$ and interpret the equality as a congruence modulo $a$, and then express the resulting sum in a suitable form:
+for $0<|r|\ll X^{1/3}$, since $\mathcal{S}_{r}(X)=8S_{r}(X)$ by symmetry. Interpreting the equation as a congruence modulo $a$, we write
 $S_{r}(X)$ $=\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leq a,b,c\leq X\\ bc\equiv-r(\textnormal{mod }a)\\ a-r\leq bc\leq aX-r\end{subarray}}1$
@@
 $S_{r}(X)=S_{r}^{(0)}(X)+O\left((r+X)X^{\varepsilon}\right).$ (3.2)
-Now we attach a weight function to $S_{r}^{(0)}(X)$. Let $X>0$ be a large number and let $H$ be a parameter to be fixed later subject to the condition
+Now we approximate $S_{r}^{(0)}(X)$ by a weighted version. Let $X>0$ be a large number and let $H$ be a parameter to be fixed later subject to the condition
 $\sqrt{X}\leq H\leq X.$ (3.3)
@@
 $(2)\ w(x)=1\textnormal{ if }x\in[1,X],$ (3.5)
-(3) $(3)\ 0<w(x)<1\textnormal{ if }x\in(1-H,1)\cup(X,X+H),$ (3.6)
+$(3)\ 0<w(x)<1\textnormal{ if }x\in(1-H,1)\cup(X,X+H),$ (3.6)
 $(4)\ w^{(j)}(x)\ll_{j}H^{-j}\textnormal{ for every }j\geq 1.$ (3.7)
@@
 $S_{r}^{(0)}(X)=S_{w}(X,r)+O\left(HX^{1+\varepsilon}\right).$ (3.9)
-Our task now is to evaluate $S_{w}(X,r)$ asymptotically. Let us first define $l=(a,c),r_{1}=r/l,c_{1}=c/l,a_{1}=a/l.$ Applying Lemma 2.1) to the $b$-sum and separating the zero frequency, we obtain
+Our task now is to evaluate $S_{w}(X,r)$ asymptotically. Let us define
+$l=(a,c),r_{1}=r/l,c_{1}=c/l,a_{1}=a/l$
+and
+$F_{a,c}(x):=w(x)w\left(\frac{cx}{a}\right),$ (3.10)
+for any integer $a\neq 0$ and any integer $c$. Applying Poisson summation to the $b$-sum and separating the zero frequency, we obtain
 $S_{w}(X,r)$ $=\sum_{l|r}\mathop{\sum\sum}_{\begin{subarray}{c}(a_{1},c_{1})=1\\ c_{1}\neq 0\\ a_{1}\geq 1\end{subarray}}w(la_{1})w(lc_{1})\sum_{\begin{subarray}{c}b\in{\mathbb{Z}}\\ b\equiv-r_{1}\bar{c_{1}}(\textnormal{mod }a_{1})\end{subarray}}w(b)w\left(\frac{bc_{1}}{a_{1}}\right)$ (3.11)
@@
 $=A_{w}(X,r)+B_{w}(X,r),$
-say, where we define for any integer $a\neq 0$ and any integer $c$,
-$F_{a,c}(x):=w(x)w\left(\frac{cx}{a}\right).$ (3.11)
+say.
 3.1. The main term
-The following proposition shows that we can isolate a main term of size $X^{2}$ from $A_{w}(X,r)$. This will be proved in the next section.
+We have,
 Proposition 3.1.
 $A_{w}(X,r)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right).$ (3.12)
+This will be proved in the next section.
 3.2. The error term
-The following proposition expresses $B_{w}(X,r)$ as a sums of Kloosterman sums twisted by an oscillatory phase function where the sum over the modulus is smooth and thus it can be further analyzed by applying the Kuznetsov formula.
+The following intermediate proposition expresses $B_{w}(X,r)$ as a sums of Kloosterman sums twisted by an oscillatory phase function where the sum over the modulus is weighted by a smooth function.
 Proposition 3.2.
-(The first main proposition)
-We have,
+(The first main proposition) We have,
 $B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{\sqrt{2}\pi}\sum_{l|r}\mathop{\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}}_{nl\leq m}\sum_{a_{1}>0}\frac{S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\sin\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)w\left(\frac{ma_{1}}{n}\right)$
@@
 $+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right).$
-The next proposition gives a strong bound for $B_{w}(X,r)$ obtained by applying the spectral methods of automorphic forms.
+Remark 3.2.
+A direct application of the Weil bound for Kloosterman sum at this stage yields the bound
+$B_{w}(X,r)\ll X^{\varepsilon}\left(\frac{X^{2}}{\sqrt{H}}+X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right).$
+Now, from (3.2), (3.9), (3.11), and Prop. 3.1, putting the optimal choice for $H$; i.e., $H=X^{2/3}$, we obtain
+${S_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(\max\left(X^{5/3},|r|\right)X^{\varepsilon}\right).$
+The next step is to exploit the cancellations in the sum of Kloostermna sums using tools from the spectral theory of automorphic forms and the outcome is the next proposition.
 Proposition 3.3.
@@
 3.3. Proof of Theorem 1.1
-The theorem follows from Prop. 3.2 and Prop. 3.4, taking into account (3.10), (3.9) and (3.2) after making the choice $H=\sqrt{X}$.
+The theorem follows from Prop. 3.1 and Prop. 3.3, taking into account (3.11), (3.9) and (3.2) after making the choice $H=\sqrt{X}$.
 4. The Main term and The Proof of Proposition 3.1
@@
 We compare this with the range of the $x$-integral and note that
-$\left[\frac{a_{1}}{c_{1}},\frac{a_{1}X}{c_{1}}\right]\cap[1,X]=\begin{cases}[1,\frac{a_{1}X}{c_{1}}],&\text{ if }a_{1}\leq c_{1}\\ [\frac{a_{1}}{c_{1}},X],&\text{ if }a_{1}>c_{1}\end{cases}.$
+[…]
 In view of the above observations, it is natural to decompose $A_{w}(X,r)$ as
 $A_{w}(X,r)$ $=\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}a_{1}\leq c_{1}\leq\frac{X}{l}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}X}{c_{1}}}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}a_{1}\leq c_{1}\leq\frac{X}{l}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}X}{c_{1}}}^{X}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x$
-$+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}}{c_{1}}}^{X}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}}{c_{1}}}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right),$
-where the contribution of the terms with $a_{1}>X/l$ is absorbed in the error term. We write the above equation more simply as
-$A_{w}(X,r)=A_{1,w}(X)+A_{2,w}(X)+A_{3,w}(X)+A_{4,w}(X)+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right).$ (4.1)
-Executing the integral and removing the coprimality condition by Möbius inversion, we easily obtain
+$+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}}{c_{1}}}^{X}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}}{c_{1}}}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right)$
+$=A_{1,w}(X)+A_{2,w}(X)+A_{3,w}(X)+A_{4,w}(X)+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right),$
+say, where the contribution of the terms with $a_{1}>X/l$ is absorbed in the error term. Now, we easily obtain
 $A_{1,w}(X)$ $=\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}a_{1}\leq c_{1}\leq\frac{X}{l}\\ (a_{1},c_{1})=1\end{subarray}}\left(\frac{a_{1}X}{c_{1}}-1\right)=\frac{\sigma(r)}{r}\frac{X^{2}}{\zeta(2)}+O\left(\frac{\sigma(r)}{r}X(\log X)^{2}\right),$
-and $A_{3,w}(X)$ satisfies the same asymptotic formula.
-Now we consider $A_{2,w}(X)$ and $A_{4,w}(X)$.
-Making the substitution $u=\frac{c_{1}x}{a_{1}}$ in $A_{2,w}(X)$, we note that the length of the $u$-integral is at most $O(H)$ due to the support of $w$ (see (3.4)) and hence $A_{2,w}(X)\ll\frac{\sigma(r)}{r}HX^{1+\varepsilon}$, and the same bound also holds for $A_{4,w}(X)$.
+by executing the integral and removing the coprimality condition by Möbius inversion, and $A_{3,w}(X)$ satisfies the same asymptotic formula.
+Also, using the support of $w$ (see (3.4)), it follows that $A_{2,w}(X),A_{4,w}(X)\ll\frac{\sigma(r)}{r}HX^{1+\varepsilon}$.
 This proves Prop.
-3.2.
+3.1.
 5. The error term: Initial Transformations
-Now we begin analyzing the sum $B_{w}(X,r)$. Interchanging the order of summation and applying Lemma 2.2 to the sum over $c_{1}$,
+Our goal now is to obtain the following intermediate proposition, where we have used the notation
+$\mathop{\sum_{l,m,n}}:=\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}.$ (5.1)
+Proposition 5.1.
+We have,
+$B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)$
+$-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}$
+$+\frac{1}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)$
+$+e\left(2\sqrt{\frac{mnu}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right)$
+$=T_{1,w}(X)-T_{2,w}(X)+T_{3,w}(X)+T_{4,w}(X)+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right),$
+say.
+We proceed to prove the proposition and the proof ends in §7.8.
+Interchanging the order of summation and applying Poisson summation to the sum over $c_{1}$ after subdividing into residue classes modulo $a_{1}$, we have
 $B_{w}(X,r)$ $=\sum_{l|r}\sum_{a_{1}>1}\frac{w(la_{1})}{a_{1}^{2}}\sum_{\begin{subarray}{c}n\in{\mathbb{Z}}\\ n\neq 0\end{subarray}}\sum_{m\in{\mathbb{Z}}}S(nr_{1},-m,a_{1})\widehat{G}(n/a_{1},m/a_{1}),$ (5.2)
@@
 $G(x,y)=w(x)w(ly)w\left(\frac{xy}{a_{1}}\right).$
-The lemma below shows that $\widehat{F_{a_{1},c_{1}}}\left(\frac{n}{a_{1}}\right)$ is negligibly small if $n\gg\frac{X^{1+\varepsilon}}{lH}$ and $\frac{1}{l}\widehat{F_{la_{1},x}}\left(\frac{m}{la_{1}}\right)$ is negligibly small if $m\gg\frac{X^{1+\varepsilon}}{H}$.
+The decay of $\widehat{G}$ limits the effective ranges of the variables $m$ and $n$ to $\frac{X^{1+\varepsilon}}{H}$ and $\frac{X^{1+\varepsilon}}{lH}$, respectively, as shown in the lemma below.
+First recall the definition in (3.10).
 Lemma 5.2.
@@
 $\widehat{F_{a,c}}(y)\ll_{k}H^{1-k}y^{-k}\left(1+\left|\frac{c}{a}\right|^{k-1}\right),$ (5.3)
-for every integer $k\geq 1$, where $F_{a,c}(x)$ is defined in (3.11). In particular, by taking $k=1$, we have the bound
+for every integer $k\geq 1$. In particular, by taking $k=1$, we have the bound
 $\widehat{F_{a,c}}(y)\ll_{k}y^{-1}.$ (5.4)
@@
 Now we consider the other summands.
-Since $w^{(j)}(x)$ is non-zero only in the intervals $(1-H,1)$ and $(X,X+H)$ for every $j\geq 1$, all these terms together contribute only $O(y^{-k}H^{1-k})$, again by (3.7).
+Since $w^{(j)}(x)$ is non-zero only in the intervals (1H,1)(1-H,1) and $(X,X+H)$ for every $j\geq 1$, all these terms together contribute only $O(y^{-k}H^{1-k})$, again by (3.7).
 Finally, by applying the bound (5.4) in one of the integrals and by trivially bounding the other integral using the support of the function $w$, we obtain the bound (5.5).
 ∎
-Now we isolate the $m=0$ term in $B_{w}(X,r)$ and write
-$B_{w}(X,r)=B_{0,w}(X)+B_{1,w}(X),$ (5.5)
+Now we isolate the $m=0$ term in $B_{w}(X,r)$. As the Kloosterman sum becomes a Ramanujan sum which satisfies the bound $|r_{q}(n)|\leq(q,n)$ (see [IK04, Eq. (3.5)]), we easily obtain the following:
+$B_{w}(X,r)=B_{1,w}(X)+O(X^{1+\varepsilon}),$ (5.6)
 where $B_{1,w}(X)$ denotes the sum over all $m\neq 0$.
-5.1. Estimation of $B_{0,w}(X)$
-Note that the Kloosterman sum appearing in $B_{0,w}(X)$ is nothing but the Ramanujan sum $r_{a_{1}}(nr_{1})$, and by the bound $|r_{q}(n)|\leq(q,n)$ (see [IK04, Eq.
-(3.5)]), we easily obtain the following:
-Proposition 5.2.
-We have,
-$B_{0,w}(X)\ll X^{1+\varepsilon}.$ (5.6)
 5.1. Analysis of $B_{1,w}(X)$: initial steps
@@
 Our goal is to improve upon the size of the error term in terms of $X$. Note that the contribution of the terms $a_{1}$ so that $X<la_{1}\leq X+H$ is quite small. We consider a more general sum
-$\mathcal{E}(Y)$ $:=\sum_{l|r}\sum_{0\neq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{0\neq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{\begin{subarray}{c}a_{1}>0\\ la_{1}\in(Y,Y+H]\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}^{2}}\widehat{G}(n/a_{1},m/a_{1}).$
+$\mathcal{E}(Y)$ $=\sum_{l|r}\sum_{0\neq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{0\neq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{\begin{subarray}{c}a_{1}>0\\ la_{1}\in(Y,Y+H]\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}^{2}}\widehat{G}(n/a_{1},m/a_{1}).$
 Proposition 5.3.
@@
 Proof.
-This follows easily from (5.4) and the Weil bound (Lemma 2.3). ∎
+This follows easily from (5.5) and the Weil bound for Kloosterman sums. ∎
 Remark 5.1.
-In view of the above proposition, whenever we have a sum over $a_{1}$, we can disregard the portion $\sum_{la_{1}>X}$ by adding the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$.
+In view of the above proposition, wherever we have a sum over $a_{1}$, we can disregard the part $\sum_{la_{1}>X}$ by adding the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$.
 Thus, from now onwards, whenever we encounter a sum of the above form, we may assume that the variable $a_{1}$ lies in the interval $[1,X]$ but we may not display the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ on every occasion since this term is already present in the error term in the final result, i.e., Prop.
-3.3.
+3.2.
 Now we state a result that will be applied many times. We omit the proof since it follows easily from the Weil bound.
@@
 Let $l\geq 1$, and $\alpha,\beta,\gamma\in{\mathbb{R}}$, such that $\alpha\leq\frac{3}{2}$ and $\beta,\gamma\leq 1$. Then for $0<U\leq X$ we have
-$\Delta(\alpha,\beta,\gamma,l,U)$ $:=\sum_{\begin{subarray}{c}0\neq n\ll\frac{X^{1+\varepsilon}}{lH}\end{subarray}}\frac{1}{n^{\gamma}}\sum_{\begin{subarray}{c}0\neq m\ll\frac{X^{1+\varepsilon}}{H}\end{subarray}}\frac{1}{m^{\beta}}\sum_{\begin{subarray}{c}1\leq la_{1}\leq U\end{subarray}}\frac{w(la_{1})\left|S(nr_{1},-m,a_{1})\right|}{a_{1}^{\alpha}}$
+$\Delta(\alpha,\beta,\gamma,l,U)$ $=\sum_{\begin{subarray}{c}0\neq n\ll\frac{X^{1+\varepsilon}}{lH}\end{subarray}}\frac{1}{n^{\gamma}}\sum_{\begin{subarray}{c}0\neq m\ll\frac{X^{1+\varepsilon}}{H}\end{subarray}}\frac{1}{m^{\beta}}\sum_{\begin{subarray}{c}1\leq la_{1}\leq U\end{subarray}}\frac{w(la_{1})\left|S(nr_{1},-m,a_{1})\right|}{a_{1}^{\alpha}}$
 $\ll U^{-\alpha+\frac{3}{2}+\varepsilon}\left(\frac{X}{H}\right)^{-\gamma-\beta+2}.$
@@
 The other terms can be handled by the Weil bound and they contribute only to the error term.
-Thus we obtain the following intermediate proposition, where we have used the notation
-$\mathop{\sum_{l}\sum_{m}\sum_{n}}:=\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}.$ (5.9)
-Proposition 5.6.
-We have,
-$B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)$
-$-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}$
-$+\frac{1}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)$
-$+e\left(2\sqrt{\frac{mnu}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right).$
-$=T_{1,w}(X)-T_{2,w}(X)+T_{3,w}(X)+T_{4,w}(X)+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right),$
-say.
-We proceed to prove the proposition below and the proof ends in §7.8.
 6. Contributions of $I_{1},I_{2},I_{3},$ and $I_{6}$
@@
 Proof.
-It is enough to estimate the integrals $I_{1},I_{2},I_{3},$ and $I_{6}$ since their contributions can be estimated easily by the Weil bound or using Prop. 5.5.
+It is enough to bound the integrals as the outside sums are estimated trivially after applying the Weil bound (see Prop.
+5.4).
 We first prove (iii) and consider only $I_{2}$ since $I_{6}$ can be handled in a similar manner.
-To bound $I_{2}$, we first interchange the $x$ and $y$ integrals and then use integration by parts in the $x$-integral, obtaining
-$I_{2}$ $=-\frac{1}{2\pi i}\frac{a_{1}}{n}\int\limits_{\frac{X}{l}}^{\frac{X+H}{l}}w(ly)e\left(\frac{-my}{a_{1}}\right)\int\limits_{x\leq la_{1}}\left(w^{\prime}(x)w\left(\frac{xy}{a_{1}}\right)+\frac{y}{a_{1}}w(x)w^{\prime}\left(\frac{xy}{a_{1}}\right)\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y+O\left(\frac{a_{1}H}{nl}\right)$
-$\ll\frac{a_{1}H}{nl},$
+Interchanging the order of the two integrals appearing in $I_{2}$ and using integration by parts in the $x$-integral, we get
+$I_{2}\ll\frac{a_{1}H}{nl},$
 where we have used the bound (3.7) and the fact that $w^{\prime}(x)\neq 0$ only on two intervals of length $H$ (see (3.4), (3.5), and (3.6)).
@@
 In the first one, let us denote it by $I_{3}$$^{\ast}$, we interchange the $x$ and the $y$ integrals and the change of variable $y\rightarrow-y$ yields
-${I_{3}}^{\ast}$ $:=\int\limits_{0}^{a_{1}(H-1)}w(-ly)e\left(\frac{my}{a_{1}}\right)\int\limits_{1}^{\min\{la_{1},a_{1}(H-1)/y)\}}w\left(\frac{-xy}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y$
+${I_{3}}^{\ast}$ $=\int\limits_{0}^{a_{1}(H-1)}w(-ly)e\left(\frac{my}{a_{1}}\right)\int\limits_{1}^{\min\{la_{1},a_{1}(H-1)/y)\}}w\left(\frac{-xy}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y$
 $=\int\limits_{0}^{(H-1)/l}\int\limits_{1}^{la_{1}}+\int\limits_{(H-1)/l}^{a_{1}(H-1)}\int\limits_{1}^{a_{1}(H-1)/y}={I_{3,1}}^{\ast}+{I_{3,2}}^{\ast},$
@@
 ${I_{3,1}}^{\ast}\ll\frac{a_{1}H}{nl}.$
-For ${I_{3,2}}^{\ast}$, we note that from the support of the function $w$ (see (3.4)), $w(-ly)=0$ if $y>(H-1)/l$. Hence,
+For $I_{3,2}$$^{\ast}$, we note that from the support of the function $w$ (see (3.4)), w(ly)=0w(-ly)=0 if y>(H1)/ly>(H-1)/l. Hence,
 ${I_{3,2}}^{\ast}=0.$
@@
 $I_{1}=\frac{1}{2\pi i}\frac{a_{1}}{m}\int\limits_{1}^{la_{1}}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-m}{x}\right)\textnormal{d}x+O\left(\frac{a_{1}H}{m}\right)+O\left(\frac{a_{1}^{2}}{mn}\right).$ (6.1)
-The contribution of the $O$-terms is $O(X^{3/2+\varepsilon})$ by Prop. 5.5. We now analyze the exponential integral
+The contribution of the $O$-terms is $O(X^{3/2+\varepsilon})$ by Prop. 5.4. We now analyze the exponential integral
 $I^{\prime}_{m,n,a_{1}}=\int\limits_{1}^{la_{1}}e(h_{1}(x))\textnormal{d}x,\quad\text{where}\quad h_{1}(x)=-\frac{nx}{a_{1}}-\frac{m}{x}.$
@@
 $x_{0}=\sqrt{\frac{a_{1}m}{n}}.$
-Let us assume further $m,n>0$ since the analysis in the other case is similar and we get the same bounds. We consider two cases.
-Case I: $x_{0}\in[2,2la_{1}/3]$. Here we split the interval $[1,la_{1}]$ in three parts as follows:
+Let us assume further $m,n>0$ since the analysis in the other case is similar and we get the same bound. We consider two cases.
+Case I: $x_{0}\in[2,2la_{1}/3]$. Here we split the interval $1,la_{1}]$ in three parts as follows:
 $I^{\prime}_{m,n,a_{1}}=\left(\int_{1}^{x_{0}-\delta x_{0}}+\int_{x_{0}-\delta x_{0}}^{x_{0}+\delta x_{0}}+\int_{x_{0}+\delta x_{0}}^{la_{1}}\right)e(h_{1}(x))\textnormal{d}x=J_{1}+J_{2}+J_{3},$
@@
 $I_{4}=\frac{1}{2\pi i}\frac{a_{1}}{m}\int_{la_{1}}^{X}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mX}{x}\right)\textnormal{d}x+O\left(\frac{a_{1}H}{m}\right)+O\left(\frac{a_{1}^{2}}{mn}\right).$
-By Prop. 5.5, the contributions of the $O$-terms are absorbed into the error term. Define
-$I_{m,n,a_{1}}:=\int_{la_{1}}^{X}e(h(x))\,\textnormal{d}x,\quad\textnormal{ where }\quad h(x):=-\frac{nx}{a_{1}}-\frac{mX}{x}.$
+By Prop. 5.4, the contributions of the $O$-terms are absorbed into the error term. Define
+$I_{m,n,a_{1}}=\int_{la_{1}}^{X}e(h(x))\,\textnormal{d}x,\quad\textnormal{ where }\quad h(x)=-\frac{nx}{a_{1}}-\frac{mX}{x}.$
 Differentiating $h(x)$, we get
 $h^{\prime}(x)=-\frac{n}{a_{1}}+\frac{mX}{x^{2}},\quad\text{and}\quad h^{\prime\prime}(x)=-\frac{2mX}{x^{3}}.$ (7.1)
-If $mn<0$, then there is no stationary point and $|h^{\prime}(x)|\geq\frac{n}{a_{1}}$ throughout the interval and in this case we use the first derivative bound directly to the integral $I_{m,n,a_{1}}$.
-This gives $I_{m,n,a_{1}}\ll\frac{a_{1}}{n}$ and the contribution of this is sufficiently small by Prop.
-5.5.
-The case left to consider is $mn>0$.
+If $mn<0$, then there is no stationary point we easily get the bound $I_{m,n,a_{1}}\ll\frac{a_{1}}{n}$ and the contribution of this term is sufficiently small by Prop.
+5.4.
+The case left to consider is $mn>0$ and we shall assume that both $m$ and $n$ are positive since the complimentary case will be similar.
 In this case the stationary point is given by
 $x_{0}=x_{0}(m,n,a_{1}):=\sqrt{\frac{ma_{1}X}{n}}.$ (7.2)
-We will assume that $m,n>0$ in our analysis since the other case is identical.
 Now, if we give an upper bound for exponential integrals and then sum over the three variables $m,n,$ and $a_{1}$ as in the case of $I_{1}$ then the error term we get is not good enough.
-We, therefore, refine our analysis by localizing the stationary point $x_{0}$ and extracting the stationary phases wherever they occur and estimating the ensuing error terms carefully.
+We, therefore, refine our analysis by localizing the stationary point $x_{0}$ and extracting the stationary phases and estimating the ensuing error terms carefully.
 This leads to the following:
@@
 $-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\Biggr{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+\frac{X^{2}}{H}\right)\right).$
-Before we begin the proof, let us first observe that from (7.2), we record the following crude range for $x_{0}$: $x_{0}\ll X\sqrt{X/H}\ll X^{5/4}$.
+Before we proceed with the proof, let us first observe the following crude range for $x_{0}$: $x_{0}\ll X\sqrt{X/H}\ll X^{5/4}$. This is immediate from (7.2).
 7.2. Localizing the stationary point and decomposition of $B_{1,w}(X,I_{4})$
-Let us define, for a fixed $a_{1}$,
-$k=k(a_{1})$ $:=\max\left\{j\geq 0:\frac{X}{2^{j}}\geq la_{1}+1\right\}=\left[\frac{\log(X/(la_{1}+1))}{\log 2}\right].$ (7.3)
+For a fixed $a_{1}$, define
+$k=k(a_{1})$ $=\max\left\{j\geq 0:\frac{X}{2^{j}}\geq la_{1}+1\right\}=\left[\frac{\log(X/(la_{1}+1))}{\log 2}\right].$ (7.3)
 Now, (7.3) gives the following upper bound on the variable $j$:
-for $\text{for }0\leq j\leq k=k(a_{1})\leq J(X),\text{ we have }X/2^{j}\geq la_{1},\text{ where }J(X):=\left[\frac{\log(X/2)}{\log 2}\right].$ (7.4)
-We make a dyadic subdivision of the range of $x_{0}$, i.e., the interval $[1,X^{5/4}]$ in the following manner:
-$\mathcal{I}_{0}:=\left[X+\delta,X^{\frac{5}{4}}\right],\mathcal{I}_{j}:=\left[\frac{X}{2^{j}}+\delta,\frac{X}{2^{j-1}}-\delta\right]\text{ for }1\leq j\leq k,\text{ and }\mathcal{I}_{k+1}:=[la_{1}+\delta,\frac{X}{2^{k}}-\delta],\mathcal{I}_{k+2}:=[1,la_{1}-\delta].$ (7.5)
+$\text{for }0\leq j\leq k=k(a_{1})\leq J(X),\text{ we have }X/2^{j}\geq la_{1},\text{ where }J(X):=\left[\frac{\log(X/2)}{\log 2}\right].$ (7.4)
+We make a dyadic subdivision of the range of $x_{0}$, i.e., the interval $1,X^{5/4}]$ in the following manner:
+$\mathcal{I}_{0}=\left[X+\delta,X^{\frac{5}{4}}\right],\mathcal{I}_{j}=\left[\frac{X}{2^{j}}+\delta,\frac{X}{2^{j-1}}-\delta\right]\text{ for }1\leq j\leq k,\text{ and }\mathcal{I}_{k+1}=[la_{1}+\delta,\frac{X}{2^{k}}-\delta],\mathcal{I}_{k+2}=[1,la_{1}-\delta].$ (7.5)
 Let us define
@@
 Here $\delta>0$ is to be chosen later (we will see that $\delta=1/100$ suffices).
-Clearly, we can write the interval $[1,X^{5/4}]$ as a disjoint union of the $\mathcal{I}_{j}$’s and $\mathcal{I}^{\prime}_{j}$s, and thus we express $B_{1,w}(X,I_{4})$ as
-$B_{1,w}(X,I_{4})$ $=\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{0\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$ (7.7)
-$+\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{1\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}^{\prime}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$
+Clearly, we can write the interval $1,X^{5/4}]$ as a disjoint union of the $\mathcal{I}_{j}$’s and $\mathcal{I}^{\prime}_{j}$s, and thus we decompose the sum as follows:
+$B_{1,w}(X,I_{4})$ $=\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{0\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$ (7.7)
 $=\mathcal{S}+\mathcal{T},$
-say.
-7.3. Subdividing the range of integration and decomposition of the sum $\mathcal{S}$
-We first subdivide the range of the integral $[la_{1},X]$. Let
-$\mathcal{F}_{j}:=\left[\frac{X}{2^{j}},\frac{X}{2^{j-1}}\right],\text{ for }1\leq j\leq k\text{ and }\mathcal{F}_{k+1}:;=\left[la_{1},\frac{X}{2^{k}}\right].$ (7.8)
-Clearly, $[la_{1},X]=\bigcup\limits_{1\leq j\leq k+1}\mathcal{F}_{j},$ where an interval $\mathcal{F}_{j}$ can intersect its neighbour only at the boundary point. Let us define
+where in the sum $\mathcal{S}$, the stationary point $x_{0}\in\cup_{j}\mathcal{I}_{j}$ and in the sum $\mathcal{T}$, $x_{0}\in\cup_{j}\mathcal{I}^{\prime}_{j}$.
+7.3. Subdivision of the range of integration and decomposition of the sum $\mathcal{S}$
+Let
+$\mathcal{F}_{j}=\left[\frac{X}{2^{j}},\frac{X}{2^{j-1}}\right],\text{ for }1\leq j\leq k\text{ and }\mathcal{F}_{k+1}:;=\left[la_{1},\frac{X}{2^{k}}\right].$ (7.8)
+Clearly, $la_{1},X]=\bigcup\limits_{1\leq j\leq k+1}\mathcal{F}_{j},$ where an interval $\mathcal{F}_{j}$ can intersect its neighbour only at the boundary point. Let us define
 $y_{j}:=\frac{X}{2^{j}},\text{ for }0\leq j\leq k+1,\text{ and }y_{k+2}:=\frac{la_{1}}{2}.$ (7.9)
@@
 $y_{k+1}\leq la_{1}\leq y_{k}.$ (7.10)
-Note that for $11\leq j\leq k+$ 1, the interval $\mathcal{I}_{j}$ is contained in $\mathcal{F}_{j}$ and the edges of $\mathcal{I}_{j}$ are away from the edges of $\mathcal{F}_{j}$.
+Note that for $1\leq j\leq k+1$, the interval $\mathcal{I}_{j}$ is contained in $\mathcal{F}_{j}$ and the edges of $\mathcal{I}_{j}$ are away from the edges of $\mathcal{F}_{j}$.
 For isolating the stationary phase when the stationary point $x_{0}\in\mathcal{I}_{j}$ we express the sum $\mathcal{S}$ as
@@
 $I^{+}_{m,n,a_{1}}(y_{j}):=\int\limits_{la_{1}}^{y_{j}}e(h(x))\textnormal{d}x.$
-Here the main idea is to subdivide the ranges of $m$, $n$ and $a_{1}$ according to the position of $x_{0}$.
-First we interchange the sum over $j$ and the sum over $a_{1}$ using (7.4) and (7.9).
-Then we bring the sum over $j$ outside the sums over the variables $m$ and $n$.
-Thus we have,
+Here the main idea is to subdivide the ranges of $m$, $n$ and $a_{1}$ according to the position of $x_{0}$. Interchanging the order of summations and using (7.4) and (7.9), we have
 $Q_{1}(X)\ll\sum_{l|r}\sum_{1\leq j\leq J(X)}\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{\begin{subarray}{c}0<la_{1}\leq y_{j}\\ \sqrt{\frac{ma_{1}X}{n}}>y_{j}+\delta\end{subarray}}\frac{1}{m}\frac{|S(nr_{1},-m,a_{1})|w(la_{1})}{a_{1}}|I_{m,n,a_{1}}^{+}(y_{j})|.$ (7.13)
-Note that the sum over $a_{1}$ is vacuous if $j$ is so large that $X/2^{j}<la_{1}$. We now make a dyadic subdivision of the range of $x_{0}$ and write
-$[y_{j}+\delta,X^{3/2}]=\bigcup_{\begin{subarray}{c}Y=2^{j}\delta\\ 0\leq j\leq L(X)\end{subarray}}[y_{j}+Y,y_{j}+2Y],$
+Note that the sum over $a_{1}$ is vacuous if $j$ is so large that $X/2^{j}<la_{1}$. Now write
+$y_{j}+\delta,X^{3/2}]=\bigcup_{\begin{subarray}{c}Y=2^{j}\delta\\ 0\leq j\leq L(X)\end{subarray}}[y_{j}+Y,y_{j}+2Y],$
 where $L(X)=O(\log X)$. We consider two cases.
 Case I: $\delta\leq Y\leq\frac{y_{j}}{2}$. Suppose $y_{j}+Y\leq x_{0}\leq y_{j}+2Y,$ where $\delta\leq Y\leq\frac{y_{j}}{2}$. This implies
-$\frac{n}{m}\leq\frac{a_{1}X}{(y_{j}+Y)^{2}}\ll\frac{1}{l}\left(\frac{X}{y_{j}}\right)^{2},$ (7.14)
-and
-$\frac{n}{mX}(y_{j}+Y)^{2}\leq a_{1}\leq\frac{n}{mX}(y_{j}+2Y)^{2},$ (7.15)
+$\frac{n}{m}\leq\frac{a_{1}X}{(y_{j}+Y)^{2}}\ll\frac{1}{l}\left(\frac{X}{y_{j}}\right)^{2},\quad\textnormal{ and }\quad\frac{n}{mX}(y_{j}+Y)^{2}\leq a_{1}\leq\frac{n}{mX}(y_{j}+2Y)^{2},$ (7.14)
 and hence, the size of $a_{1}$ is given by $a_{1}\asymp\frac{n{y_{j}}^{2}}{mX}$ and the number of such $a_{1}$ is $O\left(\frac{ny_{j}Y}{mX}+1\right)$. Moreover, from (7.14)
@@
 By (7.1), the first inequality of (7.15) and Prop.
-5.5, this case contributes $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ in $Q_{1}(X)$ (see (7.12)).
+5.4, this case contributes $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ in $Q_{1}(X)$ (see (7.12)).
 Combining this with the bound in Case I, we prove Prop.
-7.3.
+7.2.
 ∎
@@
 $=\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3},$
-say. By an argument similar to the one used in proving Prop. 7.3, we can show that
+say. By an argument similar to the one used in proving Prop. 7.2, it follows that
 $\mathcal{T}_{1},\mathcal{T}_{3}\ll X^{3/2+\varepsilon}.$
@@
 7.6. Analysis of $\mathcal{S}_{3}$, $\mathcal{S}_{5}$, $\mathcal{S}_{7}$ and extraction of the stationary phase
-Since $x_{0}\in\mathcal{I}_{j}$, and $\mathcal{I}_{j}\subset\mathcal{F}_{j}$, we extract the stationary phases by applying Lemma 2.6 and thus obtain that for $11\leq j\leq k+$ 1,
+Since $x_{0}\in\mathcal{I}_{j}$, and $\mathcal{I}_{j}\subset\mathcal{F}_{j}$, we extract the stationary phases by applying Lemma 2.9 and thus obtain that for $1\leq j\leq k+1$,
 $\int\limits_{\mathcal{F}_{j}}e(h(x))\textnormal{d}x=\frac{m^{\frac{1}{4}}X^{\frac{1}{4}}a_{1}^{\frac{3}{4}}}{\sqrt{2}n^{\frac{3}{4}}}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)+O(R_{1})+O(R_{2}),$ (7.18)
@@
 As in the proof of Prop.
-7.3, by introducing dyadic subdivision of the range of the stationary point and analyzing the ensuing restrictions on the variables $m,n$ and $a_{1}$, we show that the contributions of the error terms in $\mathcal{S}_{3}$, $\mathcal{S}_{5}$ and $\mathcal{S}_{7}$ is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the Weil bound.
+7.2, by introducing dyadic subdivision of the range of the stationary point and analyzing the ensuing restrictions on the variables $m,n$ and $a_{1}$, we show that the contributions of the error terms in $\mathcal{S}_{3}$, $\mathcal{S}_{5}$ and $\mathcal{S}_{7}$ is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the Weil bound.
 Furthermore, from (7.18), we observe that the stationary phase analysis of the integrals in $\mathcal{S}_{3},\mathcal{S}_{5},\mathcal{S}_{7}$ give the same main term.
 Combining all these estimates in one place, we conclude that
-$Q_{3}(X):=\mathcal{S}_{3}+\mathcal{S}_{5}+\mathcal{S}_{7}$
+$Q_{3}(X)=\mathcal{S}_{3}+\mathcal{S}_{5}+\mathcal{S}_{7}$
 $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}\sum_{a_{1}>0}}_{x_{0}\in\cup_{1\leq j\leq k+1}\mathcal{I}_{j}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)+O\left(X^{\frac{3}{2}+\varepsilon}\right).$
-Now we extend the range of $x_{0}$ to the whole interval $[la_{1},X]$ and subtract the contribution of the extra terms that have added.
+Now we extend the range of $x_{0}$ to the whole interval $la_{1},X]$ and subtract the contribution of the extra terms that have been added.
 Note that the total contribution of these extra terms is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the same kind of argument used in the proof of Prop.
-7.4.
+7.3.
 Thus, we have
 $Q_{3}(X)$ $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}\sum_{a_{1}>0}}_{\sqrt{\frac{ma_{1}X}{n}}\in[la_{1},X]}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)+O\left(X^{\frac{3}{2}+\varepsilon}\right).$ (7.19)
-Combining Prop. 7.3, Prop. 7.4, (7.20) and also considering the case $m,n<0$, we prove Prop. 7.1.
+Combining Prop. 7.2, Prop. 7.3, (7.19) and also considering the case $m,n<0$, we prove Prop. 7.1.
 7.7. Contribution of $I_{5}$
@@
 $I_{5}=a_{1}\int\limits_{X}^{X+H}w(u)\int\limits_{la_{1}}^{X}\frac{1}{x}e\left(\frac{-nx}{a_{1}}\right)e\left(\frac{-mu}{x}\right)\textnormal{d}x\textnormal{d}u+O\left(\frac{a_{1}H^{2}}{X}\right).$ (7.20)
-By Prop. 5.5, the contribution of the $O$-term in $B_{1,w}(X)$ is $O\left(X^{3/2+\varepsilon}\right)$.
+By Prop. 5.4, the contribution of the $O$-term in $B_{1,w}(X)$ is $O\left(X^{3/2+\varepsilon}\right)$.
 Let us now compare the above expression for $I_{5}$ with that of $I_{4}$.
@@
 7.1 and Prop.
-7.5, we arrive at Prop.
-5.6, except for the conditions on $a_{1}$.
+7.4, we arrive at Prop.
+5.1, except for the conditions on $a_{1}$.
 Now, the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ appearing in the sum over $a_{1}$ in Prop.
@@
 A similar consideration shows that the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ is equivalent to saying $0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}$.
-Remark 7.6.
-At this stage, a direct application of the Weil bound for Kloosterman sum (see Lemma 2.3), yields the bound
-$B_{w}(X,r)\ll X^{\varepsilon}\left(\frac{X^{2}}{\sqrt{H}}+X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right).$
-Now, from (3.2), (3.9), (3.10), and Prop. 3.2, the optimal choice for $H$ is given by $H=X^{2/3}$ and this gives us
-${S_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(X^{5/3+\varepsilon}\right).$
-8. Removal of the sharp-cut condition and the proof of Proposition 3.3
-To simplify notation, let us, henceforth, denote the variable $a_{1}$ by $c$.
+7.9. Removal of the sharp-cut condition and the proof of Proposition 3.2
+To simplify notation, let us, henceforth, denote the variable $a_{1}$ by $c$ with the understanding that this has no connection the original variable $c$.
 In Prop.
-5.6, we expressed $\widehat{G}(n/c,m/c)$ in a form that enabled us to see its oscillating behaviour and its order of magnitude.
-The next step is to apply the Kuznetsov formula to the sums over the variable $c$, but first we need to make the variable smooth.
-We show how to do it for $T_{1,w}(X)$.
+5.1, we expressed $\widehat{G}(n/c,m/c)$ in a form that enabled us to see its oscillating behaviour and its order of magnitude.
+We now need to make the variable smooth and we show how to do it for $T_{1,w}(X)$.
 The condition $c\leq\min\left\{\frac{mX}{nl^{2}},\frac{nX}{m}\right\}$ can be restated as
@@
 Proposition 7.5.
-$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{0<c\leq\frac{nX}{m}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
-$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{mc}{n}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
+$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{nl\leq m}\sum_{0<c\leq\frac{nX}{m}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
+$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{mc}{n}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
 An analogous statement holds in the complimentary situation, i.e.,
-$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m\leq nl}\sum_{0<c\leq\frac{mX}{nl^{2}}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
-$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m\leq nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{cnl^{2}}{m}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
+$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{m\leq nl}\sum_{0<c\leq\frac{mX}{nl^{2}}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
+$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{m\leq nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{cnl^{2}}{m}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
 Proof.
@@
 where
-$R_{1,w}(X)$ $:=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right),$
+$R_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right),$
 and
-${R^{\prime}}_{1,w}(X)$ $:=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m<nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{nl^{2}c}{m}\right).$
+${R^{\prime}}_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{m<nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{nl^{2}c}{m}\right).$
 We can decompose $T_{3,w}(X)$ in a similar manner as
@@
 where
-$R_{3,w}(X)$ $:=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right),$ (8.4)
+$R_{3,w}(X)$ $=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right),$ (7.24)
 and
-${R^{\prime}}_{3,w}(X)$ $:=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{nl^{2}c}{m}\frac{X}{u}\right).$
+${R^{\prime}}_{3,w}(X)$ $=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{nl^{2}c}{m}\frac{X}{u}\right).$
 The sum $T_{2,w}(X)$ (resp.
@@
 Combining these expressions for $T_{i,w}(X),i=1,2,3,4$, we obtain Prop.
-3.3.
+3.2.
 8. Proof of Proposition 3.3: initial steps
-We only discuss the cases of $T_{1,w}(X)$ and $T_{3,w}(X)$ below as the other two sums can be handled similarly.
-Moreover, it is enough to consider $R_{1,w}(X)$ and $R_{3,w}(X)$ as ${R^{\prime}}_{1,w}(X)$ and ${R^{\prime}}_{3,w}(X)$ are similar to $R_{1,w}(X)$ and $R_{3,w}(X)$, respectively.
+It is enough to consider the sums $T_{1,w}(X)$ and $T_{3,w}(X)$ as the other two sums can be handled similarly.
+Moreover, it is enough to consider $R_{1,w}(X)$ and $R_{3,w}(X)$ as ${R^{\prime}}_{1,w}(X)$ and ${R^{\prime}}_{3,w}(X)$ are similar to the other two sums.
 8.1. Analysis of $R_{1,w}(X)$
@@
 for every $x>0$. Thus we have,
-$R_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{C\text{ dyadic}}}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (9.1)
-×$\times\sum_{c}V\left(\frac{c}{C}\right)\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right).$
-When $C$ is small enough, the Weil bound already gives a good bound (see Prop. 5.5) and we collect all such terms in the error term and write
-$R_{1,w}(X)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{\begin{subarray}{c}C\text{ dyadic}\\ C\geq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\end{subarray}}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots+O\left(X^{\frac{5}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{1}{2}}\right).$ (9.2)
-Now we fix a divisor $l$ of $r$ and consider the sum
-$R_{1}(M,N,C,X):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right)V\left(\frac{c}{C}\right)g(m,n,c),$
+$R_{1,w}(X)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{\begin{subarray}{c}C\text{ dyadic}\\ C\geq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\end{subarray}}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots+O\left(X^{\frac{5}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{1}{2}}\right),$ (8.1)
+where we have collected the contributions of the smaller $c$’s in the error term after applying the Weil bound (see Prop. 5.4). It is, therefore, enough to estimate the sum
+$R_{1}(M,N,C,X)=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}(m/M)b_{2}(n/N)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right)V\left(\frac{c}{C}\right)g(m,n,c),$
 where
-$g(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{3}{4}}.$ (9.3)
-We define
+$g(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{3}{4}}$ (8.2)
+and $l$ is a fixed divisor of $r$. Let us define
 $C_{0}:=\max\left\{2^{j}:2^{j}\leq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\right\},$ (8.3)
 so that the contribution of the terms corresponding to $C\leq C_{0}$ is absorbed in the error term after we make the choice $H=\sqrt{X}$ at the end.
-Note that in the sum $R_{1}(M,N,C,X)$, the dyadic variables $M$ and $N$ satisfy $1\leq M\ll\frac{X^{1+\varepsilon}}{H}$, $1\leq N\ll\frac{X^{1+\varepsilon}}{lH}$ and from the support of the function $w$ (see (3.6)), we have $C\leq\frac{8NX}{M}$ as $N\leq n\leq 4N$ and $M\leq m\leq 4M$.
-For estimating $R_{1,w}(X)$, it is enough to estimate $R_{1}(M,N,C,X)$ and here is the main result we need.
+Recall that $1\leq M\ll\frac{X^{1+\varepsilon}}{H}$, $1\leq N\ll\frac{X^{1+\varepsilon}}{lH}$ and from (3.6), we have $C\leq\frac{8NX}{M}$ since $N\leq n\leq 4N$ and $M\leq m\leq 4M$.
 Proposition 8.1.
@@
 $C_{0}^{\star}:=\max\left\{2^{j}:2^{j}\leq\frac{X^{\frac{2}{3}+\varepsilon}}{l}\right\},$ (8.5)
-for a suitable $\varepsilon>0$. By Prop. 5.5, we have,
-$\frac{e\left(\frac{1}{8}\right)}{\sqrt{2}}$ $\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{0<c\leq C_{0}^{\star}}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)\ll X^{\frac{3}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{3}{2}}.$ (9.7)
+for a suitable $\varepsilon>0$. By Prop. 5.4, we have,
+$\frac{e\left(\frac{1}{8}\right)}{\sqrt{2}}$ $\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}_{nl\leq m\frac{u}{X}}\sum_{0<c\leq C_{0}^{\star}}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)\ll X^{\frac{3}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{3}{2}}.$ (8.6)
 The choice of $C_{0}^{\star}$ is dictated by the above bound, since the optimal choice for $H$ is $\sqrt{X}$.
 For larger values of $c$ we follow the same approach as above and we get the following sum after dyadic subdivision which is analogous to $R_{1}(M,N,C,X)$:
-$R_{3}(M,N,C,X,u)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)V\left(\frac{c}{C}\right)g^{\star}(m,n,c),$
+$R_{3}(M,N,C,X,u)$ $=\mathop{\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right)V\left(\frac{c}{C}\right)g^{\star}(m,n,c),$
 where
 $g^{\star}(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{1}{4}}.$ (8.7)
-We then have the following proposition which can be compared with Prop. 9.1.
+We then have the following proposition which can be compared with Prop. 8.1.
 Proposition 8.2.
@@
 $R_{3}(M,N,C,X,u)\ll r^{\theta}X^{\varepsilon}\left(X^{\frac{1}{4}}\left(\frac{X}{H}\right)^{2}+X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right).$
-We give the details of the proof of Prop. 9.1 below and the proof of Prop. 9.2 is almost identical and we only give a very brief sketch of it at the end.
+We give the details of the proof of Prop. 8.1 below and the proof of Prop. 8.2 is almost identical and we only give a very brief sketch of it at the end.
 9. Application of The Kuznetsov formula
@@
 $=\Sigma_{\text{disc.}}+\Sigma_{\text{cont.}},$
-say, where
-$\check{f}_{m,n,C}(\eta)=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$
-From the support of $V$, the argument $t$ of $f_{m,n,C}(t)$ (see (10.1)) is very small. Indeed, it lies in the the interval
+say, where $\check{f}_{m,n,C}(\eta)$ is defined in (2.2). From the support of $V$, the argument $t$ of $f_{m,n,C}(t)$ (see (9.1)) is very small. Indeed, it lies in the the interval
 $\frac{\pi\sqrt{mnr_{1}}}{C}<t\leq\frac{4\pi\sqrt{mnr_{1}}}{C}\leq\frac{16\pi\sqrt{MNr_{1}}}{C}\ll 1,$ (9.2)
-since $C>C_{0}$ (see (9.4)), $H\geq\sqrt{X}$ (see (3.3)) and $r=O(X^{1/3})$ (see Remark 16.1). We now expand $K_{2i\eta}(t)$ as a power series:
+since $C>C_{0}$ (see (8.3)), $H\geq\sqrt{X}$ (see (3.3)) and $r=O(X^{1/3})$ (see Remark 15.1). We now expand $K_{2i\eta}(t)$ as a power series Reference?:
 $K_{2i\eta}(t)$ $=\frac{\pi}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\left(\frac{\left(\frac{t}{2}\right)^{-2i\eta+2\nu}}{\Gamma(1-2i\eta)\nu!(1-2i\eta)\dots(\nu-2i\eta)}-\frac{\left(\frac{t}{2}\right)^{2i\eta+2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}\right)$
@@
 where
-$\mathcal{D}_{\pm}(M,N,C):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,\pm}(\kappa_{j}),$ (10.3)
-$\mathcal{C}_{\pm}(M,N,C)$ $:=\frac{1}{\pi}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{-\infty}^{\infty}(nmr_{1})^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}_{m,n,C,\pm}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta,$ (10.4)
+$\mathcal{D}_{\pm}(M,N,C)=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,\pm}(\kappa_{j}),$ (9.3)
+$\mathcal{C}_{\pm}(M,N,C)$ $=\frac{1}{\pi}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{-\infty}^{\infty}(nmr_{1})^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{\cosh(\pi\eta)\check{f}_{m,n,C,\pm}(\eta)}{\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta,$ (9.4)
 and
-$\check{f}_{m,n,C,\pm}(\eta)$ $:=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta,\pm}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$ (10.5)
+$\check{f}_{m,n,C,\pm}(\eta)$ $=\frac{4}{\pi}\int\limits_{0}^{\infty}K_{2i\eta,\pm}(t)f_{m,n,C}(t)\frac{\textnormal{d}t}{t}.$ (9.5)
 Thus we decompose $R_{1}(M,N,C,X)$ as
@@
 where
-$I_{\pm}(m,n,C,\nu,\eta,r_{1})$ $:=\int t^{\mp 2i\eta+2\nu}w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(\frac{-X^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right)\frac{\textnormal{d}t}{t}.$
+$I_{\pm}(m,n,C,\nu,\eta,r_{1})$ $=\int t^{\mp 2i\eta+2\nu}w\left(\frac{4\pi m\sqrt{mnr_{1}}}{nt}\right)V\left(\frac{4\pi\sqrt{mnr_{1}}}{tC}\right)g\left(m,n,\frac{4\pi\sqrt{mnr_{1}}}{t}\right)e\left(\frac{-X^{\frac{1}{2}}t^{\frac{1}{2}}(mn)^{\frac{1}{4}}}{\sqrt{\pi}r_{1}^{1/4}}\right)\frac{\textnormal{d}t}{t}.$
 We collect various estimates for the above integrals. By the change of variables $t=u^{2}\sqrt{r_{1}}$,
@@
 10. Estimation of $\mathcal{D}_{+}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$
-Here, the phase function $-h_{+}$ has no stationary point and the first derivative bound together with the Weyl law suffices.
+Here, the phase function $h_{+}$ has no stationary point and the first derivative bound together with the Weyl law suffices.
 Proposition 10.1.
 We have, for any $\varepsilon>0$
-$\check{f}_{m,n,C,+}(\eta)\ll\begin{cases}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta^{\frac{3}{2}}}e^{-\pi|\eta|}&\text{if }\eta>0,\\ X^{-100},&\text{if }\eta\gg\frac{MCX^{\varepsilon}}{NH}.\end{cases}$
+[…]
 Proof.
@@
 Let $1\ll K_{1},K_{2}\ll X^{2}$, and $\lambda\geq 0$. For any fixed $c_{1},c_{2}>0$ and any positive integer $q$, we have,
-$\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $:=\sum_{\begin{subarray}{c}c_{1}K_{1}\leq\kappa_{j}\leq c_{2}K_{2}\end{subarray}}\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)|\rho_{j}(nq)\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\lambda}}$
-$\ll\begin{cases}q^{\theta+\varepsilon}(MN)^{1+\varepsilon}K_{2}^{2-\lambda+\varepsilon}&\text{if}\quad 0\leq\lambda\leq 2\\ q^{\theta+\varepsilon}(MN)^{1+\varepsilon}\text{if}\quad\lambda>2.\end{cases}$
+$\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $=\sum_{\begin{subarray}{c}c_{1}K_{1}\leq\kappa_{j}\leq c_{2}K_{2}\end{subarray}}\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)|\rho_{j}(nq)\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\lambda}}$
+[…]
 Proof.
-We subdivide the $j$-sum into dyadic pieces with $\kappa_{j}\asymp K$ such that $c_{1}K_{1}\leq K\leq c_{2}K_{2}$, and apply the Cauchy-Schwarz inequality to the $m$-sum and the $n$-sum, getting
+We subdivide the $j$-sum into dyadic pieces with $\kappa_{j}\asymp K$ such that $c_{1}K_{1}\leq K\leq c_{2}K_{2}$, and apply the Cauchy-Schwarz inequality to both the sums over $m$ and $n$, getting
 $\Theta_{c_{1},c_{2}}(M,N,K_{1},K_{2},\lambda)$ $\ll\sum_{c_{1}K_{1}\leq K\leq c_{2}K_{2}}\frac{1}{K^{\lambda}}\sum_{\kappa_{j}\asymp K}\frac{\cosh(\pi\kappa_{j})}{e^{\pi\kappa_{j}}}\sqrt{MN}\left(\sum_{n\sim N}\frac{|\rho_{j}(nq)|^{2}}{\cosh(\pi\kappa_{j})}\right)^{\frac{1}{2}}\left(\sum_{m\sim M}\frac{|\rho_{j}(m)|^{2}}{\cosh(\pi\kappa_{j})}\right)^{\frac{1}{2}}$
@@
 Using (9.3), Prop.
-11.1, Prop.
-11.3, and (9.5), we easily get
+10.1, Prop.
+10.3, and (8.4), we easily get
 $\mathcal{D}_{+}(M,N,C)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\mathop{\sum_{m\sim M}\sum_{n\sim N}}_{nl\leq m}\sum_{1\ll\kappa_{j}\ll\frac{MCX^{\varepsilon}}{NH}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\kappa_{j}^{\frac{3}{2}}}\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}X^{\varepsilon}\left(r_{1}^{\theta}MN\left(\frac{MC}{NH}\right)^{\frac{1}{2}}\right).$
@@
 For $1<\lambda\leq 2$, we have
-$\Theta(M,N,K,\lambda)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{0}^{\infty}(nr_{1}m)^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta$
+$\Theta(M,N,K,\lambda)$ $=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{0}^{\infty}(nr_{1}m)^{-i\eta}\sigma_{2i\eta}(nr_{1})\sigma_{2i\eta}(m)\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta$
 $\ll MNX^{\varepsilon}.$
@@
 $\Theta(M,N,K,\lambda)$ $\ll\sum_{m\asymp M}\sum_{n\asymp N}(mnr_{1})^{\varepsilon}\left[\int_{0}^{\varepsilon}+\int_{\varepsilon}^{\infty}\right]\frac{e^{-\pi|\eta|}\cosh(\pi\eta)}{\eta^{\lambda}\left|\zeta(1+2i\eta)\right|^{2}}\textnormal{d}\eta.$
-We know from [Tit86, (2.1.16)] that for $0<\eta<\varepsilon$,
-$\eta^{\lambda}\zeta(1+2i\eta)^{2}=\eta^{\lambda}\left(\frac{1}{2i\eta}+\gamma+O\left(\eta\right)\right)^{2}\gg\eta^{\lambda-2}.$
-Also, from [Tit86, (3.6.5)], for $\eta\geq\varepsilon$,
-$\frac{1}{\left|\zeta(1+2i\eta)\right|}\ll_{\varepsilon}|\log\eta|^{7}.$
+We know from [Tit86, (2.1.16)] that for $0<\eta<\varepsilon$ and also from [Tit86, (3.6.5)], for $\eta\geq\varepsilon$ respectively,
+$\eta^{\lambda}\zeta(1+2i\eta)^{2}=\eta^{\lambda}\left(\frac{1}{2i\eta}+\gamma+O\left(\eta\right)\right)^{2}\gg\eta^{\lambda-2},\textnormal{ and }\frac{1}{\left|\zeta(1+2i\eta)\right|}\ll_{\varepsilon}|\log\eta|^{7}.$
 The proposition follows since both the integrals are bounded by constants that depend only on $\varepsilon$. ∎
@@
 Proof.
-This is clear from Prop. 11.1, (10.4) and Prop. 11.5. ∎
+This is clear from Prop. 10.1, (9.4) and Prop. 10.5. ∎
 11. Estimation of $\mathcal{D}_{-}(M,N,C)$ for $C_{0}\leq C\leq C_{1}$
 Our next task is to estimate $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$ (see (9.3), (9.4) and (9.6)).
-The sum $\mathcal{D}_{-}(M,N,C)$ (see (10.3)) is considerably more difficult to estimate than $\mathcal{D}_{+}(M,N,C)$ because the phase function $-h_{-}(u)$ (see (10.10)) can have a stationary point inside the range of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ (see (10.8)).
+The sum $\mathcal{D}_{-}(M,N,C)$ (see (9.3)) is considerably more difficult to estimate than $\mathcal{D}_{+}(M,N,C)$ because the phase function $h_{-}(u)$ (see (9.10)) can have a stationary point inside the range of the integral $I_{-}(m,n,C,\nu,\eta,r_{1})$ (see (9.8)).
 Differentiating $h_{-}(u)$ (see (9.10)),
@@
 $C_{1}:=\max\left\{2^{j}:2^{j}\leq\frac{XN}{100M}\right\}.$ (11.3)
-From (3.5) and the fact that $V$ is supported in $[1,4]$, it follows that $w(\frac{4\pi m\sqrt{mn}}{nu^{2}})=1$ if $C_{0}<C\leq C_{1}$. In this range of $C$, the integral we need to analyze becomes
-$I_{-}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu+i\eta}\int_{{\mathbb{R}}}A(u)e(-h_{-}(u))\textnormal{d}u,$
-where
-$A(u)=\frac{u^{4\nu}}{u}V\left(\frac{4\pi\sqrt{mn}}{u^{2}C}\right)g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right).$
+From (3.5) and the fact that $V$ is supported in $[1,4]$, it follows that $w\left(\frac{4\pi m\sqrt{mn}}{nu^{2}}\right)=1$ if $C_{0}<C\leq C_{1}$.
+In this range of $C$, the integral we need to analyze becomes
+$I_{-}(m,n,C,\nu,\eta,r_{1})=2r_{1}^{\nu+i\eta}\int_{{\mathbb{R}}}A(u)e(-h_{-}(u))\textnormal{d}u,\textnormal{ where }A(u)=\frac{u^{4\nu}}{u}V\left(\frac{4\pi\sqrt{mn}}{u^{2}C}\right)g\left(m,n,\frac{4\pi\sqrt{mn}}{u^{2}}\right).$
 When $C$ is in this range we are able to apply Lemma 2.12 since the condition (2.6), which translates to the condition $\eta\gg(mC/nH)X^{\varepsilon}$, is satisfied.
@@
 Dropping the condition $nl\leq m$ by positivity as we had done earlier will be too costly here.
-We remove this condition by approximating the indicator function $\mathbf{1}_{[1,\infty)}(m/nl)$ by the Perron formula (Lemma 2.4).
+We remove this condition by approximating the indicator function $\mathbf{1}_{[1,\infty)}(m/nl)$ by the Perron formula.
 Furthermore, the variables $m$ and $n$ occurring in the argument of the function $V$ are separated by Mellin inversion.
@@
 $p_{k}(u_{0})=\frac{1}{k!}\left(\frac{i}{4\pi h^{\prime\prime}_{-}(u_{0})}\right)^{k}G^{(2k)}(u_{0}),$ (11.5)
-where $A$ and $\delta$ are arbitrary subject to the conditions $A>0$ and $100<\delta<1/$ 10. We define
+where $A$ and $\delta$ are arbitrary subject to the conditions $A>0$ and $0<\delta<1/10$. We define
 $G(u)=A(u)e(-H(u)),\quad\text{where}\quad H(u)=h_{-}(u)-h_{-}(u_{0})-\frac{1}{2}h^{\prime\prime}_{-}(u_{0})(u-u_{0})^{2}.$ (11.6)
@@
 where
-$\mathcal{D}_{-,k}(M,N,C)$ $:=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\frac{1}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}$ (12.8)
-×$\times\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\left[\frac{2r_{1}^{\nu+i\eta}}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)p_{k}(u_{0})\right],$
+$\mathcal{D}_{-,k}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\rho_{j}(nr_{1})\rho_{j}(m)\frac{1}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}$ (11.8)
+$\times\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\left[\frac{2r_{1}^{\nu+i\eta}}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)p_{k}(u_{0})\right],$
 and
-$\mathcal{D}_{-,Z}(M,N,C)$ $:=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\left|\frac{\rho_{j}(nr_{1})\rho_{j}(m)\kappa_{j}^{-A}}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}\right|$ $\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}r_{1}^{\nu}}{\nu!\left|(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})\right|}.$ (12.9)
+$\mathcal{D}_{-,Z}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{j=1}^{\infty}\left|\frac{\rho_{j}(nr_{1})\rho_{j}(m)\kappa_{j}^{-A}}{\sinh(2\pi\kappa_{j})\Gamma(1+2i\kappa_{j})}\right|$ $\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}r_{1}^{\nu}}{\nu!\left|(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})\right|}.$ (11.9)
 We first consider the case $k=0$ in (11.8). Let us first define
-$\mathcal{L}(m,n,\kappa_{j},\nu,C):=2\frac{A(u_{0})}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)$ (12.10)
+$\mathcal{L}(m,n,\kappa_{j},\nu,C)=2\frac{A(u_{0})}{\sqrt{h^{\prime\prime}_{-}(u_{0})}}e\left(-h_{-}(u_{0})-\frac{1}{8}\right)$ (11.10)
 $=\sqrt{\frac{2\pi}{\kappa_{j}}}e\left(-\frac{2\kappa_{j}}{\sqrt{\pi}}-\frac{1}{8}\right)u_{0}^{4i\eta}{u_{0}}^{4\nu}V\left(\frac{\pi^{2}Xmn}{\kappa_{j}^{2}C}\right)\left(\frac{\pi^{2}X}{\kappa_{j}^{2}}\right)^{3/4}$
@@
 by (11.2) and (8.2). From the support of $V$, the expression for $u_{0}$ (see (11.2)), and taking into account the condition $nl\leq m$, we see that the effective range of $\kappa_{j}$ is given by
-$\frac{\pi}{2}\left(\frac{mnX}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq\pi\left(\frac{mnX}{C}\right)^{\frac{1}{2}};$ (12.11)
-in particular, we have
-$\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}.$ (12.12)
+$\frac{\pi}{2}\left(\frac{mnX}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq\pi\left(\frac{mnX}{C}\right)^{\frac{1}{2}};\quad\textnormal{in particular}\quad\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}.$ (11.11)
 Even though the presence of the function $V$ restricts the range of $\kappa_{j}$, we, nevertheless, add the condition (11.11) to the sum over $\kappa_{j}$.
@@
 Hence, we write $\mathcal{D}_{-,0}(M,N,C)$ in the following way.
-$\mathcal{D}_{-,0}(M,N,C):=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)T(m,n,C),$ (12.13)
+$\mathcal{D}_{-,0}(M,N,C)=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)T(m,n,C),$ (11.12)
 where
-$T(m,n,C):=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C),$
+$T(m,n,C)=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C),$
 where
-$\mathcal{V}(\kappa_{j},\nu,r_{1})$ $:=\frac{1}{\sinh(2\pi\kappa_{j})}\frac{2^{-2i\kappa_{j}-2\nu}r_{1}^{\nu+i\kappa_{j}}}{\Gamma(1+2i\kappa_{j})\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\ll\frac{e^{-\pi\kappa_{j}}r_{1}^{\nu}}{\nu!\sqrt{\kappa_{j}}},$ (12.14)
+$\mathcal{V}(\kappa_{j},\nu,r_{1})$ $=\frac{1}{\sinh(2\pi\kappa_{j})}\frac{2^{-2i\kappa_{j}-2\nu}r_{1}^{\nu+i\kappa_{j}}}{\Gamma(1+2i\kappa_{j})\nu!(1+2i\kappa_{j})\dots(\nu+2i\kappa_{j})}\ll\frac{e^{-\pi\kappa_{j}}r_{1}^{\nu}}{\nu!\sqrt{\kappa_{j}}},$ (11.13)
 which follows from the Stirling asymptotic (see [IK04, Eq. (5.113)]). Let us also define
-$\tilde{T}(m,n,C):=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C)|.$ (12.15)
+$\tilde{T}(m,n,C)=\sum_{\nu=0}^{\infty}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)\mathcal{V}(\kappa_{j},\nu,r_{1})\mathcal{L}(m,n,\kappa_{j},\nu,C)|.$ (11.14)
 11.3. Removal of the condition $nl\leq m$
-Our next step is to capture the condition $nl\leq m$ by Lemma 2.4. So, we write
+Our next step is to capture the condition $nl\leq m$ by Lemma the Perron formula (see [Dav80, Chap. 17]). So, we write
 $\mathbf{1}_{nl\leq m}=\frac{1}{2\pi i}\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\left(\frac{m}{nl}\right)^{s_{1}}\frac{\textnormal{d}s_{1}}{s_{1}}-\frac{1}{2}\mathbf{1}_{nl=m}+O\left(\left(\frac{m}{nl}\right)^{\sigma_{1}}\min\left(1,\frac{1}{T|\log(m/nl)|}\right)+\frac{1}{T}\right),$
@@
 where
-$E_{1}:=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m\neq nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{\sigma_{1}}\frac{1}{|\log(m/nl)|}\tilde{T}(m,n,C),$
-$E_{2}:=\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\tilde{T}(m,n,C)\frac{\textnormal{d}|s_{1}|}{|s_{1}|},$
+$E_{1}=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m\neq nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{\sigma_{1}}\frac{1}{|\log(m/nl)|}\tilde{T}(m,n,C),$
+$E_{2}=\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\int\limits_{\sigma_{1}-iT}^{\sigma_{1}+iT}\tilde{T}(m,n,C)\frac{\textnormal{d}|s_{1}|}{|s_{1}|},$
 and
-$E_{3}:=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\tilde{T}(m,n,C).$
+$E_{3}=\frac{1}{T}\mathop{\sum_{m}\sum_{n}}_{m=nl}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\tilde{T}(m,n,C).$
 Proposition 11.2.
 We have,
-$E_{1}$ $\ll r_{1}^{\theta}X^{3/4}\frac{X^{1+\varepsilon}}{TH}MN,$ (12.17)
-$E_{2}$ $\ll{r_{1}}^{\theta}X^{3/4+\varepsilon}\sqrt{MN},$
-$E_{3}$ $\ll r_{1}^{\theta}\frac{X^{3/4+\varepsilon}}{T}\sqrt{MN}.$
+$E_{1}\ll r_{1}^{\theta}X^{3/4}\frac{X^{1+\varepsilon}}{TH}MN,\quad E_{2}\ll{r_{1}}^{\theta}X^{3/4+\varepsilon}\sqrt{MN},\quad E_{3}\ll r_{1}^{\theta}\frac{X^{3/4+\varepsilon}}{T}\sqrt{MN}.$ (11.16)
 Proof.
@@
 $\tilde{T}(m,n,C)$ $\ll X^{3/4}\sum_{\frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq\kappa_{j}\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi\kappa_{j}}}{\kappa_{j}^{5/2}},$ (11.17)
-which yields the bound for $E_{1}$ once we recall Prop. 11.3.
+which yields the bound for $E_{1}$ once we recall Prop. 10.3.
 For the sum $E_{2}$, using (11.17) and estimating the integral trivially, we have,
@@
 where
-$U(M,N,C):=\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\sum_{\kappa_{j}\sim K}\frac{\kappa_{j}^{2\sigma_{2}}}{\kappa_{j}^{2}}\Bigg{|}\sum_{\nu=0}^{\infty}\mathcal{V}(\kappa_{j},\nu,r_{1})\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (12.21)
-×$\times\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\right)^{4\nu}\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\Bigg{|}$
+$U(M,N,C)=\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\sum_{\kappa_{j}\sim K}\frac{\kappa_{j}^{2\sigma_{2}}}{\kappa_{j}^{2}}\Bigg{|}\sum_{\nu=0}^{\infty}\mathcal{V}(\kappa_{j},\nu,r_{1})\mathop{\sum_{m}\sum_{n}}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (11.20)
+$\times\left(\frac{2\kappa_{j}}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\right)^{4\nu}\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\Bigg{|}$
 $\ll\sum_{\begin{subarray}{c}K\text{ dyadic}\\ \frac{\pi}{2}N\left(\frac{Xl}{C}\right)^{\frac{1}{2}}\leq K\leq 4\pi\left(\frac{MNX}{C}\right)^{\frac{1}{2}}\end{subarray}}\frac{K^{2\sigma_{2}}}{K^{\frac{5}{2}}}\sum_{\nu=0}^{\infty}\left(\frac{2K}{\sqrt{\pi}X^{\frac{1}{2}}}\right)^{4\nu}\frac{r_{1}^{\nu}}{\nu!}$
-×$\times\sum_{\kappa_{j}\sim K}e^{-\pi\kappa_{j}}\left|\sum_{m}\sum_{n}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}-\nu}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\right|$
+$\times\sum_{\kappa_{j}\sim K}e^{-\pi\kappa_{j}}\left|\sum_{m}\sum_{n}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\frac{m}{nl}\right)^{s_{1}}(mn)^{-s_{2}-\nu}\rho_{j}(nr_{1})\rho_{j}(m)(mn)^{-i\kappa_{j}}\right|$
 by (11.13) and integration by parts. Observing that $e^{-\pi\kappa_{j}}\asymp 1/\cosh\pi\kappa_{j}$, the above bound yields
@@
 A routine calculation shows that for $k=1,2$, we have
-$||p_{k}||_{\infty}\ll\frac{1}{\eta}||p_{0}||_{\infty},\quad\text{and}\quad||p_{0}||_{\infty}\ll||A||_{\infty}\ll\frac{\left(\frac{\sqrt{MN}}{C}\right)^{2\nu}}{\left(\frac{\sqrt{MN}}{C}\right)^{1/2}}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}},$ (12.24)
+[…]
 where we use (9.11), (11.5) and (11.6).
@@
 $\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{1}{\eta^{3/2}}.$ (11.24)
-Combining (12.25), (12.8), (11.2), (10.2), ignoring the condition $nl\leq m$ by positivity and by Prop. 11.3, we have for any $k\geq 1$,
+Combining (11.24), (11.8), (10.2), (9.2), ignoring the condition $nl\leq m$ by positivity and by Prop. 10.3, we have for any $k\geq 1$,
 $\mathcal{D}_{-,k}(M,N,C)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\left(r_{1}^{\theta+\varepsilon}MN\right).$ (11.25)
@@
 $\mathcal{D}_{-,Z}(M,N,C)$ $\ll r^{\theta}\left(\frac{X}{H}\right)^{2+\varepsilon}.$ (11.26)
-Bringing together (12.7), (12.23), (12.26), (12.27), and taking $\delta=1/20$ and $A=2$, we complete the proof of Prop. 12.1.
+Bringing together (11.7), (11.22), (11.25), (11.26), and taking $\delta=1/20$ and $A=2$, we complete the proof of Prop. 11.1.
 12. Estimation of $\mathcal{D}_{-}(M,N,C)$ for $C>C_{1}$
@@
 where we define (see (9.7))
-$\check{f}_{m,n,C,-,\mu}(\eta):=\frac{4}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}I_{-,\mu}(m,n,C,\nu,\eta,r_{1}),$ (13.7)
+$\check{f}_{m,n,C,-,\mu}(\eta)=\frac{4}{\sinh(2\pi\eta)}\sum_{\nu=0}^{\infty}\frac{2^{-2\nu}}{\Gamma(1+2i\eta)\nu!(1+2i\eta)\dots(\nu+2i\eta)}I_{-,\mu}(m,n,C,\nu,\eta,r_{1}),$ (12.7)
 and by (12.6),
-$\mathfrak{D}_{-,\mu}(M,N,C)$ $:=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,\mu}(\kappa_{j}).$ (13.8)
+$\mathfrak{D}_{-,\mu}(M,N,C)$ $=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\kappa_{j}\ll\frac{X^{1+\varepsilon}}{H}}\rho_{j}(nr_{1})\rho_{j}(m)\check{f}_{m,n,C,-,\mu}(\kappa_{j}).$ (12.8)
 Now our aim is to estimate the above two sums separately.
@@
 7.2 and we only give a sketch.
-First note that since $\eta\ll X^{1+\varepsilon}/H$, $u_{0}=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\ll\frac{X^{\frac{1}{2}+\varepsilon}}{H}.$ We write the range $\left[\beta_{1}+\delta,\frac{X^{\frac{1}{2}+\varepsilon}}{H}\right]$ of $u_{0}$ as a union of dyadic subintervals of the form $[\beta_{1}+Y,\beta_{1}+2Y]$, where the dyadic variable $Y$ can take at most $O(\log X)$ values.
+First note that since $\eta\ll X^{1+\varepsilon}/H$, $u_{0}=\frac{2\eta}{\sqrt{\pi}X^{\frac{1}{2}}(mn)^{\frac{1}{4}}}\ll\frac{X^{\frac{1}{2}+\varepsilon}}{H}.$ We write the range $\left[\beta_{1}+\delta,\frac{X^{\frac{1}{2}+\varepsilon}}{H}\right]$ of $u_{0}$ as a union of dyadic subintervals of the form $\beta_{1}+Y,\beta_{1}+2Y]$, where the dyadic variable $Y$ can take at most $O(\log X)$ values.
 Let us consider two cases separately: $\delta\leq Y\leq\frac{\beta_{1}}{2}$ and $Y>\frac{\beta_{1}}{2}$.
@@
 $\ll\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\sum_{\begin{subarray}{c}Y\textnormal{ dyadic}\\ \delta\leq Y\leq\frac{\beta_{1}}{2}\end{subarray}}\frac{1}{YX^{\frac{1}{2}}(MN)^{\frac{1}{4}}}\sum_{c_{m,n}(\beta_{1}+Y)\leq\kappa_{j}\leq c_{m,n}(\beta_{1}+2Y)}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi|\kappa_{j}|}}{\sqrt{\kappa_{j}}}$
-by (14.5). Applying Prop. 13.3, we now get
+by (13.5). Applying Prop. 12.3, we now get
 $\mathfrak{D}_{-,1,2}(M,N,C,I)\ll r_{1}^{\theta}X^{\varepsilon}\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\left(\sqrt{M}+\frac{\sqrt{N}}{(MN)^{\frac{1}{4}}}\right),$ (13.6)
@@
 We have,
-$Y>\frac{\beta_{1}}{2}\implies Y\geq\beta_{0},\quad\text{where}\quad\beta_{0}:=\frac{1}{1000}\frac{\sqrt{M}(MN)^{\frac{1}{4}}}{\sqrt{XN}}.$ (14.9)
+$Y>\frac{\beta_{1}}{2}\implies Y\geq\beta_{0},\quad\text{where}\quad\beta_{0}=\frac{1}{1000}\frac{\sqrt{M}(MN)^{\frac{1}{4}}}{\sqrt{XN}}.$ (13.9)
 Thus, using (13.9) and (13.7), we have
@@
 The proof is finished in view of (13.6) and (13.10).
-Now we consider part (ii). Here, $\eta$ is highly localized and Lemma 2.5 with $k=2$ suffices and we apply Prop. 13.3 to handle the sum over $\kappa_{j}$. We omit the details. ∎
+Now we consider part (ii). Here, $\eta$ is highly localized and Lemma 2.8 with $k=2$ suffices and we apply Prop. 12.3 to handle the sum over $\kappa_{j}$. We omit the details. ∎
 13.3. Estimation of $\mathfrak{D}_{-,1,5}(M,N,C)$ and the weighted stationary phase analysis
@@
 where $\alpha_{1}$ and $\beta_{1}$ are defined in (13.1) and
-$R_{1}:=\frac{|A(\beta_{1})|}{|h_{-}^{\prime}(\beta_{1})|}+\frac{|A(\alpha_{1})|}{|h_{-}^{\prime}(\alpha_{1})|},\text{ }R_{2}:=\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha_{1})^{3}}+\frac{1}{(\beta_{1}-x_{0})^{3}}\right),\text{ and }R_{3}:=\frac{QT}{P^{\frac{3}{2}}}\left(1+\frac{Q}{V}\right)^{2},$ (14.11)
+$R_{1}=\frac{|A(\beta_{1})|}{|h_{-}^{\prime}(\beta_{1})|}+\frac{|A(\alpha_{1})|}{|h_{-}^{\prime}(\alpha_{1})|},\text{ }R_{2}=\frac{Q^{4}T}{P^{2}}\left(1+\frac{Q}{V}\right)^{2}\left(\frac{1}{(x_{0}-\alpha_{1})^{3}}+\frac{1}{(\beta_{1}-x_{0})^{3}}\right),\text{ and }R_{3}=\frac{QT}{P^{\frac{3}{2}}}\left(1+\frac{Q}{V}\right)^{2},$ (13.11)
 where $Q,T,P$ and $V$ are defined in Lemma 2.10. It is clear that the leading term here is $\mathcal{L}(m,n,\kappa_{j},\nu,C)$ (see (LABEL:lt)). Therefore,
 $\mathfrak{D}_{-,1,5}(M,N,C)$ $=4\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\nu=0}^{\infty}\sum_{\begin{subarray}{c}\kappa_{j}\\ \alpha_{1}+\delta<u_{0}<\beta_{1}-\delta\end{subarray}}\rho_{j}(nr_{1})\rho_{j}(m)$
-×$\times\mathcal{V}(\kappa_{j},\nu,r_{1})\Big{[}\mathcal{L}(m,n,\kappa_{j},\nu,C)+O(R_{1}+R_{2}+R_{3})\Big{]}.$
+$\times\mathcal{V}(\kappa_{j},\nu,r_{1})\Big{[}\mathcal{L}(m,n,\kappa_{j},\nu,C)+O(R_{1}+R_{2}+R_{3})\Big{]}.$
 Since we can handle the contribution of the portion of the sum where the stationary point $u_{0}$ varies over a small interval easily, we rewrite the above sum as
@@
 So we have only one extra condition $\kappa_{j}>\pi m$ from which we need to separate the variables $m$ and $\kappa_{j}$.
-By Lemma 2.4,
+By the Perron formula,
 $G(M,N,C)$ $=\sum_{\kappa_{j}>\pi m}(\cdots)=\sum_{\kappa_{j}}(\cdots)\frac{1}{2\pi i}\int\limits_{\sigma-iT}^{\sigma+iT}\left(\frac{\kappa_{j}}{\pi m}\right)^{s}\frac{\textnormal{d}s}{s}-\frac{1}{2}\sum_{\kappa_{j}}(\cdots)\mathbf{1}_{\kappa_{j}=\pi m}$ (13.14)
@@
 By (LABEL:lt), (11.13) and (9.2), the contribution of the term $\frac{1}{2}\mathbf{1}_{\kappa_{j}=\pi m}$ is bounded by
-$X^{3/4}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)\sum_{\begin{subarray}{c}\kappa_{j}\\ \kappa_{j}=\pi m\end{subarray}}|\rho_{j}(nr_{1})\rho_{j}(m)|\frac{e^{-\pi\kappa_{j}}}{\kappa_{j}^{5/2}}$ (14.16)
 […]
@@
 Combining (13.3), Prop.
-14.2, and Prop.
-14.3, Prop.
+13.2, and Prop.
+13.3, Prop.
 13.1 follows.
@@
 $\check{f}_{m,n,C,-}(\eta)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{e^{-\pi|\eta|}}{\eta}\sum_{\nu=0}^{\infty}\left(\frac{\sqrt{MNr_{1}}}{C}\right)^{2\nu}\frac{1}{\nu!}\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}\frac{e^{-\pi|\eta|}}{\eta}.$
-Thus, proceeding as in the proof of Prop. 11.5 and by (10.4), we have,
+Thus, proceeding as in the proof of Prop. 10.5 and by (9.4), we have,
 $\mathcal{C}_{-}(M,N,C)$ $\ll\left(\frac{C}{N^{2}l}\right)^{\frac{3}{4}}MNX^{\varepsilon}.$
@@
 Bringing together (9.6) and the estimates for $\mathcal{D}_{+}(M,N,C)$, $\mathcal{C}_{+}(M,N,C)$, $\mathcal{D}_{-}(M,N,C)$ and $\mathcal{C}_{-}(M,N,C)$, i.e., Prop.
-11.4, Prop.
-11.6, Prop.
-12.1, Prop.
+10.4, Prop.
+10.6, Prop.
+11.1, Prop.
 12.1 and Prop.
-15.1, we have proved Prop.
-9.1.
+14.1, we have proved Prop.
+8.1.
 15.2. A sketch of the proof of Proposition 8.2
@@
 The proofs of the analogues of Prop.
-11.4, Prop.
-11.6, Prop.
-12.1, Prop.
-13.1, and Prop.
+10.4, Prop.
+10.6, Prop.
+11.1, Prop.
+12.1, and Prop.
 14.1 are along the same lines but with some minor changes.
@@
 11.1 comes from the term $E_{2}$ (see Prop.
-12.2), $\mathcal{D}_{-,0}(M,N,C)$ (see (12.23)), and $\mathcal{D}_{-,k}(M,N,C)$ (see (12.26)).
+11.2), $\mathcal{D}_{-,0}(M,N,C)$ (see (11.22)), and $\mathcal{D}_{-,k}(M,N,C)$ (see (11.25)).
 The term analogous to $E_{2}$ now contributes $O\left(r^{\theta}X^{\frac{1}{3}}\left(X/H\right)^{\frac{3}{2}}\right)$ and the same contribution also comes from the analogue of $\mathcal{D}_{-,0}(M,N,C)$.
@@
 For the analogue of $\mathcal{D}_{+}(M,N,C)$ in Prop.
-11.4, we now have the bound $O\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$.
-Together, the analogous of $\mathcal{C}_{-}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$ in Prop.s 15.1 and 11.6 now contribute $O\left(X^{{\frac{1}{4}}+\varepsilon}{\left(X/H\right)}^{\frac{3}{2}}\right)$.
+10.4, we now have the bound $O\left(r^{\theta}X^{\frac{1}{4}+\varepsilon}\left(X/H\right)^{2}\right)$.
+Together, the analogous of $\mathcal{C}_{-}(M,N,C)$ and $\mathcal{C}_{+}(M,N,C)$ in Prop.s 14.1 and 10.6 now contribute $O\left(X^{{\frac{1}{4}}+\varepsilon}{\left(X/H\right)}^{\frac{3}{2}}\right)$.
 15.3. Proof of Proposition 3.3
 Now, from Prop.
-9.1, (9.1), and (9.2), we obtain the bound $O\left(r^{\theta}\frac{X^{2+\varepsilon}}{H}\right)$ for $R_{1,w}(X)$, and as discussed earlier, the bound for ${R^{\prime}}_{1,w}(X)$ is the same.
+8.1 and (8.1), we obtain the bound $O\left(r^{\theta}\frac{X^{2+\varepsilon}}{H}\right)$ for $R_{1,w}(X)$, and as discussed earlier, the bound for ${R^{\prime}}_{1,w}(X)$ is the same.
 Hence we obtain the bound $O\left(r^{\theta}X^{\varepsilon}\left(\frac{X^{2}}{H}+X\sqrt{H}\right)\right)$ for $T_{1,w}(X)$ (see (7.22)) and the same bound holds for $T_{2,w}(X)$.
@@
 3.3 follows from the above discussion and Prop.
-5.6.
+5.1.
 Remark 15.1.